### v2

# A Comprehensive Survey on Data Augmentation

Abstract.
Data augmentation is a series of techniques that generate high-quality artificial data by manipulating existing data samples. By leveraging data augmentation techniques, AI models can achieve significantly improved applicability in tasks involving scarce or imbalanced datasets, thereby substantially enhancing AI models’ generalization capabilities. Existing literature surveys only focus on a certain type of specific modality data, and categorize these methods from modality-specific and operation-centric perspectives, which lacks a consistent summary of data augmentation methods across multiple modalities and limits the comprehension of how existing data samples serve the data augmentation process. To bridge this gap, we propose a more enlightening taxonomy that encompasses data augmentation techniques for different common data modalities. Specifically, from a data-centric perspective, this survey proposes a modality-independent taxonomy by investigating how to take advantage of the intrinsic relationship between data samples, including single-wise, pair-wise, and population-wise sample data augmentation methods. Additionally, we categorize data augmentation methods across five data modalities through a unified inductive approach.
Data augmentation, Data-centric taxonomy, Multi-modality
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXX
1. Introduction
With the rapid development of Artificial Intelligence (AI) in the past decades, AI methods have shown their superiority over human beings and other traditional approaches across most tasks, from universal tasks in our daily lives such as image recognition (He et al., 2016) and text translation (Bahdanau et al., 2015) to complicated scientific research tasks such as protein structure prediction (Jumper et al., 2021) and weather forecasting (Bi et al., 2023). Recently, powered by Latent Diffusion Models (Rombach et al., 2022) and Transformer (Vaswani et al., 2017), Stable Diffusion (Rombach et al., 2022) and ChatGPT (OpenAI, 2023) have greatly changed human beings’ working, living, and entertainment. The success of AI products is usually attributed to AI models’ in-depth understanding of the accumulated data, thereby intrinsically uncovering data patterns and learning data-task correlations. The performances of AI models will be affected by the quantity and quality of training data, e.g. models trained with limited data suffer from over-fitting with significantly degenerative performance on testing datasets, and models trained with unbalanced samples will lead to poor generalization ability. In most situations, researchers have to overcome difficulties caused by limited data sample amounts and imbalanced distribution of data samples. An intuitive solution to these problems is to acquire more data, but data are scarce or difficult to collect in many cases, and labeling the data is yet another labor-intensive work.
To solve these problems, data augmentation has been extensively applied and proven to be effective and efficient (Zha et al., 2023). The core idea of data augmentation is to artificially enlarge the training dataset by creating modified copies of existing data. It also introduces more diversity and fills the gap between training datasets and real-world applications. As technology continues to evolve, a variety of techniques are applied to augment data samples. Some simple approaches only randomly mask part of the data (DeVries and Taylor, 2017b; Zhong et al., 2020; Wei and Zou, 2019), while more sophisticated augmentation schemes involve generative adversarial networks (Bowles et al., 2018) or reinforcement learning agents (Cubuk et al., 2019).
An ideal survey on data augmentation should be modality-independent, because only in this way can it focus on the intrinsic mechanisms of data augmentation, regardless of the data type, and thus provide insight into the nature of data augmentation. Existing literature surveys summarize data augmentation methods from different perspectives (Mikołajczyk and Grochowski, 2018; Shorten and Khoshgoftaar, 2019; Khalifa et al., 2022; Mumuni and Mumuni, 2022; Feng et al., 2021; Shorten et al., 2021; Bayer et al., 2022; Ding et al., 2022; Zhao et al., 2022; Zhou et al., 2022; Borisov et al., 2022; Sauber-Cole and Khoshgoftaar, 2022; Fonseca and Bacao, 2023; Wen et al., 2021; Iglesias et al., 2023; Iwana and Uchida, 2021). As summarized in Table. 1, each of them concentrates on a certain data modality. Most of them categorize data augmentation methods with modality-specific (e.g. image modality and text modality), or operation-centric (e.g. feature scaling operation, data perturbation operation) taxonomies. These surveys reflect the development of data augmentation and its application in different learning scenarios. However, they fail to cover data augmentation methods across data types, uncover their common patterns despite different modalities, and thus limit readers’ understanding of the essence of data augmentation. To fill this gap, we investigate data augmentation methods comprehensively, across five most popular data modalities. We analyze how data augmentation methods utilize different sample numbers and how they leverage different components of the data information. Then, we propose a unified taxonomy that focuses on the way data augmentation derives new data samples. Finally, we investigate up-to-date literature and categorize them by our proposed taxonomy.
The main contributions of this survey can be summarized as follows:
  • We propose a novel modality-independent taxonomy from a data-centric perspective that accommodates data augmentation techniques for all modalities consistently and inductively;
  • To the best of our knowledge, this is the first survey that covers data augmentation techniques across five data modalities, i.e. image, text, graph, tabular, and time-series modalities;
  • We investigate how information is consistently contained in each of these data modalities and can be utilized for data augmentation.
  • We include and categorize up-to-date literature in data augmentation.
Table 1. A comparison of data modalities and taxonomies discussed by existing surveys and ours.
Data Modality Taxonomies
Image Text Graph Tabular Time-series Modality Technical Procedural Sample amount Information type
IIPhDW 2018 (Mikołajczyk and Grochowski, 2018)
J Big Data 2019 (Shorten and Khoshgoftaar, 2019)
Artif Intell Rev 2022 (Khalifa et al., 2022)
Array 2022 (Mumuni and Mumuni, 2022)
IJCNLP 2021 (Feng et al., 2021)
J Big Data 2021 (Shorten et al., 2021)
Comput Surv 2022 (Bayer et al., 2022)
KDD 2022 (Ding et al., 2022)
arXiv 2022 (Zhao et al., 2022)
arXiv 2022 (Zhou et al., 2022)
TNNLS 2022 (Borisov et al., 2022)
J Big Data 2022 (Sauber-Cole and Khoshgoftaar, 2022)
J Big Data 2023 (Fonseca and Bacao, 2023)
IJCAI 2021 (Wen et al., 2021)
Neural Comput Appl 2023 (Iglesias et al., 2023)
Plos one 2021 (Iwana and Uchida, 2021)
Ours
2. Background
2.1. Early Evolution
Despite the fact that data augmentation only became popular in no more than a decade, some of its embryonic forms were proposed quite early. The concept of data augmentation was conceived at the very beginning of Deep Learning. For example, the use of random distortion can be found in LeNet (LeCun et al., 1998), where ninefold the distorted images are added to the dataset to verify that increasing the size of the training set can effectively reduce test error. Since then, data augmentation has gradually been recognized as a best practice for training Convolutional Neural Networks (CNNs) (Simard et al., 2003; Cireşan et al., 2011; Ciregan et al., 2012). AlexNet (Krizhevsky et al., 2012) explicitly employs several data augmentation techniques to reduce overfitting. It augments the datasets by extracting patches from images and altering the color intensity. It also applies dropout, which randomly sets the output of some neurons to zero to prevent their co-adaption. Although dropout is often regarded as an approach to regularization, it has a similar idea to data augmentation and indeed inspires a set of augmentation methods (DeVries and Taylor, 2017b). These applications of data augmentation mainly focus on increasing the size of the training dataset and introducing diversity to it. SMOTE (Chawla et al., 2002) is another early and representative work on data augmentation but from a different perspective. Focusing on addressing the class imbalance problem, it suggests that oversampling the minority class can achieve better classification performance when categories are not equally represented in the dataset.
2.2. Existing Surveys
Given the rapid development of CNNs and their wide application in image-related tasks, most early data augmentation methods are proposed for image data. Also, it is natural to augment image data because many data augmentation methods suit the traits of CNNs very well, such as translation invariance. Published in 2018, (Mikołajczyk and Grochowski, 2018) is one of the first surveys on data augmentation. It gives a brief discussion of the effect of traditional and neural-based augmentation methods. In 2019, the use of data augmentation on image data is thoroughly studied in (Shorten and Khoshgoftaar, 2019), which is up to now the most extensive work on this track. It categorizes image data augmentation methods into basic image manipulation and deep learning approaches, with a wide coverage. It also introduces how meta-learning is used to wisely choose and combine various basic operations for data augmentation. (Khalifa et al., 2022) is a recent survey on image data augmentation published in 2022. It formalizes commonly used operations for image data augmentation with equations and illustrates their effects with examples. In the same year, (Mumuni and Mumuni, 2022) also elaborates on image data augmentation with plenty of examples. It adopts taxonomies similar to those of (Shorten and Khoshgoftaar, 2019) but includes more modern approaches and evaluates the improvements in model performance when these augmentation methods are used.
Data augmentation on text data is not thoroughly researched as early as image data, possibly due to the discrete and correlated nature of text components. Still, several augmentation methods are proposed for or transplanted to text data. (Feng et al., 2021) is among the first surveys on text data augmentation. It adopts a simple taxonomy but gives more emphasis on the applications. Built on previous work on image data augmentation, (Shorten et al., 2021) continues to investigate how data augmentation is used to enrich text datasets, and methods are classified as symbolic augmentation and neural augmentation. (Bayer et al., 2022) categorizes text data augmentation methods by identifying whether they are used on the data space or feature space. It also presents the performance improvements of different methods.
Following the success of data augmentation methods for image and text data, there is a growing trend in augmenting graph data, too. Since the graph represents data in a relatively complex form compared with image and text, there are more opportunities for data augmentation. This also leads to more possible ways to categorize graph data augmentation methods. For example, (Ding et al., 2022) views graph data augmentation from the employed techniques and application scenarios, while (Zhao et al., 2022) discusses graph data augmentation in terms of its data modality, task level, and whether the augmentation operation is rule-based or learned. (Zhou et al., 2022) tries to elaborate on all feasible taxonomies for graph data augmentation and uses detailed schematic illustrations to explain the overview framework as well as the idea of some typical methods.
Tabular data is common in reality, but there is not as much room for data augmentation as other data types because tabular data lacks some features for augmentation to utilize. Some surveys only discuss augmentation techniques in the context of deep learning on tabular data (Borisov et al., 2022), and others mainly focus on how to generate tabular data (Sauber-Cole and Khoshgoftaar, 2022; Fonseca and Bacao, 2023).
With its growing application in scenarios such as speech recognition and IoT, time-series data is attracting more attention. (Wen et al., 2021) discusses basic and advanced augmentation methods. (Iglesias et al., 2023) focuses on generative approaches. (Iwana and Uchida, 2021) gives more discussion on mixing time-series data.
Till now, data augmentation techniques for different data types have been exhaustively discussed; however, there has not been a comprehensive survey that summarizes data augmentation for all types of data. Such a survey can be profitable because data augmentation techniques for different modalities share some common methodologies when leveraging information from existing data, such as altering some features of a sample or mixing values between multiple samples. Summarizing these similarities can help reveal the common pattern of data augmentation. A comprehensive survey does mean simply combining existing surveys because the proposed taxonomies in these surveys view data augmentation from different angles and are incompatible with each other. Some taxonomies contrast basic operations to advanced approaches; some categorize methods by their application scenarios. Besides, these taxonomies provide limited insight into what happens to the information. Indeed, huge gaps lie between operations for data modalities due to their distinctive nature; these operation-based or application-based taxonomies reflect the development of data augmentation techniques. However, from such taxonomies, we cannot grasp the essence of data augmentation and lack perception of the big picture of data augmentation. Hence there is a need for a comprehensive survey that covers data augmentation techniques for all data modalities with an all-embracing and information-centric taxonomy.
3. Taxonomy
Before presenting the taxonomy, we formalize data augmentation as follows. Given a labeled dataset $\mathcal{D}_{L}=\{\mathbf{X},\mathbf{y}\}$, where $\mathbf{X}$ stands for the data, and $\mathbf{y}$ stands for the labels, data augmentation can be represented by a certain function $f_{\theta}$, such that the augmented dataset $\mathcal{\tilde{D}}_{L}=\{\mathbf{\tilde{X}},\mathbf{\tilde{y}}\}$ is derived by:
(1) $f_{\theta}:\mathcal{D}_{L}=\{\mathbf{X},\mathbf{y}\}\rightarrow\mathcal{\tilde{D}}_{L}=\{\mathbf{\tilde{X}},\mathbf{\tilde{y}}\}$
This representation also applies to an unlabeled dataset $\mathcal{D}_{U}=\{\mathbf{X}\}$ or a partially labeled dataset $\mathcal{D}=\{\mathbf{X}_{L}\cup\mathbf{X}_{U},\mathbf{y}_{L}\}$, but for simplicity, in this section, we only use a labeled dataset to explain the taxonomy. Some data augmentation approaches are used in the input data space, while others are used in the latent feature space. This section uses $\mathbf{X}$ as a space-agnostic representation.
In this comprehensive survey, we try to propose a two-tier taxonomy from an information-centric perspective that can be applied to all data modalities. The main consideration of our taxonomy is whence the information in the augmented data comes. It differentiates data augmentation techniques by asking two research questions: RQ1: How many samples are used to generate each new sample? RQ2: Which part of the information is used to generate new data? The answers to these questions constitute the taxonomy hierarchy which we present below.
By answering RQ1, we can divide data augmentation approaches into Individual, Multiple, and Populational Augmentation, which build the first tier of the taxonomy in our survey.
Individual Augmentation. This type of data augmentation method transforms one data sample without referring to others to create new data. It introduces a certain perturbation to the data, such as masking and noise addition. The information of new data comes from exactly one sample in the original dataset, and each new sample often lies around its original sample in the feature space. We formalize Individual Augmentation as follows, where $\epsilon(\mathbf{x}_{i})$ stands for a certain perturbation that is related to the original data $\mathbf{x}_{i}$:
(2) $\mathbf{\tilde{x}}=\mathbf{x}_{i}+\epsilon(\mathbf{x}_{i}),\qquad\tilde{y}=y_{i}$
We further answer RQ2 to build the second tier of the taxonomy. To answer this question, we must consider how the information is contained in the data. Typically, data is made up of a certain type of element that carries unit information with a value, such as the coloring of a pixel or the wording in a sentence. For most data, there is also some structural information that connects these elements in a certain way, such as the positional relationship of pixels or the syntax of a sentence. The key to designing data augmentation methods is to think of how to properly leverage the two types of information. So, we can categorize Individual Augmentation according to which part of the information is used for data augmentation:
  • Value-based Transformation, which perturbs the value that an element carries.
  • Structure-based Transformation, which perturbs elements’ structural relationships.
Multiple Augmentation. This type of data augmentation method makes use of multiple data to acquire new ones. Information from different data samples is combined by operations such as interpolation and concatenation. New samples derived from these methods often lie between the source data in the feature space. We formalize Multiple Augmentation as follows, where $i$ and $j$ are two data samples from the original dataset, and $\lambda$ is a scalar factor that often lies between $[0,1]$:
(3) $\mathbf{\tilde{x}}=\lambda\cdot\mathbf{x}_{i}+(1-\lambda)\cdot\mathbf{x}_{j},\qquad\tilde{y}=\lambda\cdot y_{i}+(1-\lambda)\cdot y_{j}$
Now we answer RQ2 for Multiple Augmentation. Similarly to Individual Augmentation, this type of data augmentation is also conducted utilizing either value of structure information, and the detailed classification is as follows:
  • Value-based Mixture, which arithmetically mixes the values of multiple data.
  • Structure-based Combination, which pieces multiple data (or parts of them) together.
Populational Augmentation. This type of data augmentation method does not explicitly make use of one or multiple data but rather uses the population of the data. It comprehensively learns the features and abstractions of the entire dataset and samples completely new data from a learned distribution of the data. The generated data should also fall into this distribution. We formalize Populational Augmentation as follows, where $P(\cdot)$ stands for a probability distribution learned from the original dataset $\mathcal{D}_{L}=\{\mathbf{X},\mathbf{y}\}$:
(4) $\vspace{-1.5mm}\mathbf{\tilde{x}}\sim P(\mathbf{X}),\qquad\tilde{y}\sim P(\mathbf{y}|\mathbf{\tilde{x}})$
The answer to RQ2 for Populational Augmentation is different from the two categories mentioned above. When generating new data based on population distributions, most methods simultaneously consider value and structural information. Some methods also introduce external knowledge when learning from the data population. So, we have the following classification:
  • Vanilla Generation, where the generation process only resorts to the existing dataset.
  • Exogenous Generation, where the generation process refers to external resources such as other datasets or expert knowledge.
We have presented a general taxonomy for data augmentation, whereas the category is at a high level, and each data modality has its specific data augmentation techniques depending on the nature of that modality. In addition, the equations that we use as a general formalization are only abstract and conceptual. They do not literally represent all specific data augmentation techniques in these categories. From paragraph 4 to paragraph 8, we will elaborate on how this taxonomy applies to each data modality and review some typical data augmentation methods.
Figure 1. A taxonomy for data augmentation (left) and how information in each data modality is utilized (right).
4. Data Augmentation for Image Data
The elemental components of an image are pixels. A pixel carries unit information that uses a numerical value to represent the color. Each pixel also has spatial relationships with other pixels, so, images are well-structured. Data augmentation for image data revolves around the manipulation of the colors and spatial relationships of image pixels.
4.1. Individual Image Augmentation
4.1.1. Value-based Image Transformation
Two kinds of methods are used to perturb the value a pixel carries, and thus change the color it displays. One is called Pixel Erasing, which completely masks out the color information in an area of the data; another is called Photometric Transformation, which only imposes a soft adjustment to change the appearance of the image, but the essential information remains visible.
Pixel Erasing. Pixel Erasing is one of the simplest yet effective types of method in data augmentation. The idea is to mask some information from the original input space or the feature space. Typical work like Cutout (DeVries and Taylor, 2017b) and random erasing (Zhong et al., 2020) first selects a rectangle region and erases all pixels in that region. The erased region is then filled with a certain color or Gaussian noises.
Table 2. Region selection strategy for Pixel Erasing methods
Name Ref Augmentation Strategy
Random Location-based
Cutout (DeVries and Taylor, 2017b)
Image-aware Random Erasing (Zhong et al., 2020)
Object-aware Random Erasing (Zhong et al., 2020)
Region-aware Random Erasing (Yang et al., 2019)
Hide-and-seek (Kumar Singh and Jae Lee, 2017)
Grid-Mask (Chen et al., 2020a)
Random Erasing in the frequency domain (Hwang et al., 2023)
If erasing is performed by randomly selecting an area in the image, it is called Image-aware Random Erasing (IRE) (Zhong et al., 2020). In some tasks, the location of the target in the image is known, so random erasing can be restricted in the object’s bounding box, known as Object-aware random erasing (ORE) (Zhong et al., 2020). Further argued in (Yang et al., 2019), the erased area within the bounding box should be limited to a manageable size and a part of the less informative background should also be erased. (Yang et al., 2019) proposed Region-aware Random Erasing (RRE), which first introduces ORE with upper bounds for width and height in the bounding boxes, and then randomly erases the background.
Random Erasing typically adds one or two big rectangular masks to the image, and some other methods deploy more flexible implementations. Hide-And-Seek (Kumar Singh and Jae Lee, 2017) proposes to randomly remove multiple parts of the image so that the model has to learn from multiple parts of the image other than the most discriminating part. Similarly, the Grid-Mask method (Chen et al., 2020a) generates a set of squared masks under the grid layout. It can prevent excessive deletion and reservation and achieve a reasonable balance between deleted and reserved regional information.
In most research, random erasing is used in the spatial domain, while REF (Hwang et al., 2023) extends the use of random erasing to the frequency domain. It uses the discrete Fourier transform to process the image, applies random erasing on the image’s frequency domain, and uses the inverse discrete Fourier transform to recover the image to the spatial domain.
Photometric Transformation. In contrast to Pixel Erasing, which put masks on an area of the image, Photometric Transformation changes the overall color and style of the image.
Some methods alter the value of each pixel without considering other pixels. The simplest way is to adjust the brightness of the image. Color inversion replaces the pixels with their “opposite” color. Color casting applies different operations in different channels. For example, red color casting increases the image’s red channel while keeping the green channel and blue channel untouched. Noise injection adds Gaussian or other types of noises to the images. It helps the model learn from various distorted features and become robust. In addition, this method can produce a large amount of augmented data with low cost and complexity.
Some other methods decide how to adjust the value of pixels according to other pixels. Histogram equalization (Cheng and Shi, 2004; Abdullah-Al-Wadud et al., 2007), for example, enhances the image by enhancing its contrast, and white balancing simulates alternative light cast on the object. Kernel filters are matrices that slide across and are convoluted with original the image to generate new images. Depending on the value of the matrix, a kernel filter can be used to sharpen or blur the image, or make other changes. A sharpening filter helps extract the features, and a blurring filter can improve the robustness of the model. (Reichenbach et al., 1990), for instance, uses kernel filters to improve edge detection performance.
Greyscaling transforms a colored picture into black-and-white by combining channels for different colors. This change forces the model to focus on the shapes or edges of target objects instead of their colors. Apart from this, greyscaling would be especially useful in compressing data and improving the model’s efficiency. A counter-operation of grey-scaling is Grayscale-to-Color Conversion (GCC) (Gupta et al., 2020). Since most pre-trained models for image-related tasks are trained on colored images, single-channeled greyscale images should be colored to fit the models.
4.1.2. Structure-based Image Transformation
This type of Individual Image Augmentation manipulates the spatial relationship between pixels. Image Cropping breaks the original image into pieces, and Geometric Transformation perturbs the structure within the original data.
Image Cropping. Image Cropping refers to selecting an area from the original image and removing other parts of the image. Some cropping techniques are discussed in the context of reducing the cost of image storage and transmission, improving image quality, and reducing computation complexity and cost when training artificial intelligence models. However, it can also be used as a data augmentation technique, in that image cropping can produce multiple samples by selecting different regions from the original image, and prevent the models from being affected by irrelevant background information in the images. Primitive cropping methods make use of manual selection or random cropping, while some methods employ more advanced techniques to perform this task, including attention-based approaches and aesthetics-based approaches (Chen et al., 2016).
Attention-based approaches use certain attention scores to evaluate a subimage. Regions that contain the main topic or other types of information often have high attention scores. (Ma and Guo, 2004) proposes a general solution, which uses the entropy, area ratio, and center distance of the image to evaluate the performance of the image segmentation and accordingly adjust the cropping parameters. (Ciocca et al., 2007) uses an adaptive cropping method that applies different processes depending classification result of the image. These studies discuss the way to evaluate the importance of each part of the image, while they overlook how to search for the optimal cropping rectangle. Formulations of this problem and algorithms with low computational complexity are given by (Chen et al., 2016). Apart from these automatic methods, there are also some interactive image cropping methods. (Santella et al., 2006), a gaze-based image cropping approach, uses eye tracking techniques to capture human fixations and thus infer the location of important contents on the image. Unlike fully hand-crafted image cropping, the implicit interaction applied in eye tracking is less burdensome and less time-consuming.
Aesthetics-based approaches (Nishiyama et al., 2009; Cheng et al., 2010; Zhang et al., 2012) are similar to quality evaluation. For example, (Nishiyama et al., 2009) detects several subject regions and one background region, extracts features from the regions, estimates and combines posterior probabilities, and determines the quality score according to these probabilities. A2-RL (Li et al., 2018b) integrates image cropping into a sequential decision-making process and an aesthetics-aware reinforcement learning model to solve the best cropping problem.
Beyond attention-based and aesthetic-based cropping, (Yan et al., 2013) proposed a change-based approach. It takes into account the content that is removed and extracts two sets of specially designed features to model what is changed after cropping.
Geometric Transformation. As one of the most classical image data augmentation methods, Geometric Transformation changes the spatial relationship of pixels within an image while keeping the value of each pixel untouched to create more data without extra sampling or labeling. It increases the number of training samples, extends the input space, and better represents the real world by deriving more possible views from the original images. For example, flipping the image of a cat can be understood as taking another picture of the cat from a different orientation.
Affine transformations like flipping, shearing, rotation, and translation manipulate the image while preserving the colinearity and ratios of distances. Image flipping reflects the image around its vertical axis, horizontal axis, or both axes, rotation turns the image by an angle about a fixed point. Image shearing fixs all pixels along a given line, and other pixels are shifted parallel to that line by a distance proportional to their perpendicular distance from the line. Translation means moving an image in a certain direction. It can introduce more variance in the position of target objects in the image.
In non-affine transformation, the spatial relationship between pixels is further changed compared with affine transformations because the colinearity and ratio of distances are broken. One example of non-affine transformation is perspective transformation (Wang et al., 2019), which mimics images taken from other views of the object. Image stretching (Wang et al., 2008) is a kind of content-aware image resizing. It preserves or even enlarges the prominent contents in the image and compresses the background or unimportant foreground information. For example, if there is a person in the middle of an image, the middle part of the image is made wider, and the left and right parts are made narrower. Distortion (Simard et al., 2003; Jaderberg et al., 2015) also stretches the pixels but often focuses on a certain region. It aims at fixing distortion in the original images and is especially useful in digit classification tasks where the handwriting is poor or when the images are distorted by a camera lens (e.g. wide-angle lens) when collecting them. PatchShuffle (Kang et al., 2017) divides an image into block matrices and shuffles the pixels within each block under a certain probability. This transformation introduces more variance in the image and extends the input space or feature space of the training data.
4.1.3. Image Value-structure Transformation
Automatic Data Augmentation methods search for the optimized augmentation policy. The policy is derived from a search space that includes both value-based transformations and structure-based transformations. When implementing the augmentation, both pixels’ coloring and the positions are changed.
Automatic Data Augmentation. AutoAugment (Cubuk et al., 2019) is among the most significant achievements in Automatic Data Augmentation. It represents data augmentation as a decision-making process and uses reinforcement learning to search for the best operations. In AutoAugment, an operation can be any basic value-based transformation like shearing, or structure-based transformation like histogram equalization. A sub-policy consists of two operations along with their probabilities and magnitudes with which the operations are applied. A policy consists of many sub-policies. All sub-policies and their parameters form the search space (action space), and a controller RNN (agent) picks the sub-polices from the search space and gets a policy (action). Images are processed under the policy, which produces several augmented images. The augmented images are fed to downstream networks (environment) and then AutoAugment uses the network’s performance (reward) to update the controller. As a result, the controller is trained to pick the policies and parameters that can yield the best network performance. This optimal policy will be used on all images and is also transferable so that it can be used on other datasets, too. Fast AutoAugment (Lim et al., 2019) proposes a more efficient searching algorithm for augmentation policies, which works significantly faster than AutoAugment and thus allows a larger number of sub-policies, which should benefit the generalization performance. Similarly, RandAugment (Cubuk et al., 2020) removes the need for a separate search phase on a proxy task, and makes the training process less complicated and less expensive. Population Based Augmentation (PBA) (Ho et al., 2019) learns a augmentation schedule rather than a fixed policy as in AutoAugment to improve efficiency. These methods combine the benefits of both value-based transformations and structure-based transformations and further discover the potential of data augmentation with individual image data.
4.2. Multiple Image Augmentation
4.2.1. Value-based Image Mixture
This type of Multiple Image Augmentation computes an arithmetical mixture (e.g. interpolation) of the color from different images.
Image MixUp. Image mixup is a category of data augmentation methods that combine two images pixel-wise. Each pixel of the augmented image is produced by interpolating the pixels from the original images. Mixup (Zhang et al., 2017) mixes two random images on the input layer. Pixels from two original images each contribute a portion to the new image:
(5) $\mathbf{\tilde{x}}=\lambda\mathbf{x}_{i}+(1-\lambda)\mathbf{x}_{j},\qquad\mathbf{\tilde{y}}=\lambda\mathbf{y}_{i}+(1-\lambda)\mathbf{y}_{j}$
where $(\mathbf{x}_{i},\mathbf{y}_{i})$ and $(\mathbf{x}_{j},\mathbf{y}_{j})$ are two labeled images from the dataset, and $\lambda\in[0,1]$. As a variant, manifold mixup (Verma et al., 2019) is operated on the hidden layers of the neural network. Most mixup methods combine two different images, while AugMix (Hendrycks et al., 2019) applies three augmentation operations to one source image, and combines the three augmented images to generate a new sample.
In the above-mentioned mixup approaches, the mixing factor is even across the entire image, so they are region-consistent. In region-variant mixup approaches, the mixing operation varies in strength across regions. For example, SmoothMix (Lee et al., 2020) applies a soft-edged mask when mixing. Two more sophisticated methods are Co-mixup (Kim et al., 2021) and Puzzle mix (Kim et al., 2020), which both use saliency to decide the mixing strength of different regions, potentially leading to better mixing performances.
Most mixup methods create new samples with soft labels, as in Eq.5, while there are some exceptions. Sample Paring (Inoue, 2018), for example, uses the label of one of the original images, and that of the other image is left unused. ReMix (Chou et al., 2020) is designed to address the class imbalance problem, and the label of the augmented data is assigned with the label from the minority class.
Neural Blending. In mixup, we explicitly and arithmetically blend different images. However, this process can benefit from the use of neural networks.
AutoMix (Liu et al., 2022a), for example, integrates a Mix Block in the training process. It builds a bridge between the selection of mixup policy and the optimization of the model. In AutoMix, the augmentation module is optimized along with the classification feature map, and the mixup functions and mixing ratio will updated according to the classification loss. Another image blending technique is proposed as Smart Augmentation (Lemley et al., 2017). At the input end of Smart Augmentation, two random samples are fed into the network through two channels, and a CNN sub-network yields a one-channel output with the same shape as the input, which means one new image is generated based on the information of the two input images. The augmented image is then used for downstream tasks, and the augmentation sub-network adjusts to the task performance accordingly.
4.2.2. Structure-based Image Combination
Methods that combine images based on the position of pixels and image substructure are called Image Patching. It refers to putting together small fragments of many images or putting a part of an image over a second image.
Image Patching. RICAP (Takahashi et al., 2019) randomly crops four images, and patches them to construct a new image. CutMix (Yun et al., 2019) makes use of two images. It is developed based on Cutout (DeVries and Taylor, 2017b), but the removed region is replaced by a patch from another image. Attentive CutMix (Walawalkar et al., 2020) further enhances the CutMix method by adding a feature extraction module that selects an important or representative region, and pastes this attentive patch onto the other image. Image patching is not limited to rectangular shapes. CowMask (French et al., 2020), for example, introduces patching with irregular shape rather than rectangular.
Figure 2. A conceptual demonstration of (a) Value-based Individual Image Transformation, (b) Structure-based Individual Image Transformation, (c) Value-based Multiple Image Mixture, and (d) Structure-based Multiple Image Combination.
4.3. Populational Image Augmentation
4.3.1. Vanilla Image Generation
Generative Adversarial Networks. The rise of Generative Adversarial Networks (GAN) (Goodfellow et al., 2014) makes it possible for humans to generate new images on a large scale using computers. This nature also leads to a new path of data augmentation: creating new images instead of modifying or combining old ones.
GANs have now been widely used as a data augmentation technique (Bowles et al., 2018), especially in the medical field (Frid-Adar et al., 2018; Kaur et al., 2021), where data is to some extent scarce. Apart from simply enlarging the dataset, GAN can be used in other way such as removing background noise and cleaning the dataset (Kupyn et al., 2018; Sharma et al., 2019; Ashraf et al., 2021). It can also generate different views or poses of the object based on the original image (Zhang et al., 2022), and thus add more variation to the dataset.
4.3.2. Exogenous Image Generation
Neural Style Transfer introduces style information from external dataset when augmenting an entire dataset. Computer Graphic Modelling is to leverage expert knowledge on how objects in the existing dataset looks like and the power of computer software to generate new image samples.
Neural Style Transfer. Neural Style Transfer (Gatys et al., 2016) learns the artistic style from a group of image data and uses this artistic style to depict the topic of images from another dataset.
As is (Gatys et al., 2016), Neural Style Transfer can be used to mimic the drawing styles of particular artists. Such an augmentation process can help the model learn a more general representation of the topics. (Li et al., 2018a) proposes a more realistic use of neural style transfer. It transfers the weather or illumination from reference images to the target dataset. This is particularly useful because it is hard to collect sufficient images under certain weather.
Computer Graphics Modelling. Computer Graphics Modelling refers to artificially creating images by computer software, and it can be used to produce datasets for model training (Gaidon et al., 2016; McCormac et al., 2017; Barbosa et al., 2018). These datasets have several advantages over ordinary image datasets. First, it requires little human efforts to collect data. Second, by applying different aspects, poses, and lighting condition, computer graphics modelling can generate images with high diversity. Third, these images are all fully labelled with accurate ground truth.
Table 3. A summary for image data augmentation
Tier 1: Quantity of source data Tier 2: Information type Augmentation method Typical literature
Individual Image Augmentation Value-based Transformation Pixel Erasing Cutout (DeVries and Taylor, 2017b) (arXiv 2017), Random Erasing (Zhong et al., 2020) (AAAI 2020), Region-aware Random Erasing (Yang et al., 2019) (ICCT 2019), Hide-and-seek (Kumar Singh and Jae Lee, 2017) (ICCV 2017), Grid-Mask (Chen et al., 2020a) (arXiv 2020), REF (Hwang et al., 2023) (ICEIC 2023)
Photometric Transformation multi-peak GHE (Cheng and Shi, 2004) (Digital Signal Processing 2024), DHE (Abdullah-Al-Wadud et al., 2007) (TCE 2007), (Reichenbach et al., 1990) (ICPR 1990), GCC (Gupta et al., 2020) (JDI 2020)
Structure-based Transformation Image Cropping (Ma and Guo, 2004) (CCNC 2004), Self-adaptive Image Cropping (Ciocca et al., 2007) (TCE 2007), Automatic Image Cropping (Chen et al., 2016) (CVPR 2016), Gaze-based Cropping (Santella et al., 2006) (CHI 2006), Sensation-based Cropping (Nishiyama et al., 2009) (MM 2009), (Cheng et al., 2010) (MM 2010), (Zhang et al., 2012) (TIP 2012), Change-based Image Cropping (Yan et al., 2013) (CVPR 2013), A2-RL (Li et al., 2018b) (CVPR 2018)
Geometric Transformation Perspective Transformation (Wang et al., 2019) (IEEE Access 2019), Stretching (Wang et al., 2008) (SIGGRAPH 2008), Distortion (Simard et al., 2003) (ICDAR 2003), STN (Jaderberg et al., 2015) (NeurIPS 2015), PatchShuffle (Kang et al., 2017) (arXiv 2017)
Value-structure Transformation Automatic Augmentation AutoAugment (Cubuk et al., 2019) (CVPR 2019), Fast AutoAugment (Lim et al., 2019) (NeurIPS 2019), RandAugment (Cubuk et al., 2020) (CVPR 2020), PBA (Ho et al., 2019) (ICML 2019)
Multiple Image Augmentation Value-based Mixture Image Mixup Mixup (Zhang et al., 2017) (arXiv 2017), Manifold Mixup (Verma et al., 2019) (ICML 2019), AugMix (Hendrycks et al., 2019) (arXiv 2019), SmoothMix (Lee et al., 2020) (CVPR 2020), Co-Mixup (Kim et al., 2021) (arXiv 2021), Puzzle Mix (Kim et al., 2020) (ICML 2020), ReMix (Chou et al., 2020) (ECCV 2020), Sample Paring (Inoue, 2018) (arXiv 2018)
Neural Blending AutoMix (Liu et al., 2022a) (ECCV 2022), Smart Augmentation (Lemley et al., 2017) (IEEE Access 2017)
Structure-based Combination Image Patching RICAP (Takahashi et al., 2019) (TCSVT 2019), CutMix (Yun et al., 2019) (ICCV 2019), Attentive CutMix (Walawalkar et al., 2020) (ICASSP 2020), CowMask (French et al., 2020) (arXiv 2020)
Populational Image Augmentation Vanilla Generation Generative Adversarial Networks GAN Augmentation (Bowles et al., 2018) (arXiv 2018), (Frid-Adar et al., 2018) (Neurocomputing 2018), (Kaur et al., 2021) (ICAIA 2021), DeblurGAN (Kupyn et al., 2018) (CVPR 2018), (Sharma et al., 2019) (ACCV 2019), (Ashraf et al., 2021) (IEEE Access 2021), Re-ID (Zhang et al., 2022) (Neural Computing and Applications 2022)
Exogenous Generation Neural Style Transfer Neural Style Transfer (Gatys et al., 2016) (CVPR 2016), Photorealistic Image Stylization (Li et al., 2018a) (ECCV 2018)
Computer Graphics Modelling Virtual KITTI (Gaidon et al., 2016) (CVPR 2016), SceneNet RGB-D (McCormac et al., 2017) (CVPR 2017), SOMAnet (Barbosa et al., 2018) (CVIU 2018)
5. Data Augmentation for Text Data
The elemental components of text data are words, and sometimes the words are processed in the form of tokens for later convenience. Words are organized into sentences following syntactic rules. The sentence is a sequential representation since one word comes after another, but there is also a tree-structured grammatical dependency behind the sentence. Data augmentation for text data often changes the wording or reshapes the structure of a sentence.
5.1. Individual Text Augmentation
5.1.1. Value-based Text Transformation
Text data augmentation of this type uses word-level manipulation. It applies changes to the choice of words or adds or removes some words.
Invariant Replacement. Invariant Replacement refers to those augmentations that try to use alternative wording in a given sentence but convey the same meaning.
EDA (Wei and Zou, 2019), for example, proposes a set of text augmentations, including synonym replacement. It randomly selects several words from a sentence, substituting them with their synonyms, i.e. words that mean exactly or nearly the same.
There are several ways to implement Synonym Replacement. (Marivate and Sefara, 2020) uses WordNet (Miller, 1995), a lexical database to find synonyms for verbs and nouns in sentences and thus make replacements. Apart from WordNet, (Marivate and Sefara, 2020) also derives word embeddings by Word2Vec (Mikolov et al., 2013) and replaces a word with others who have high cosine similarity. Another possibility is to use a knowledge graph. A knowledge graph describes the relationships of entities, and by utilizing the “is equivalent” relation, we can easily find a series of synonyms for a given word.
A special case of synonym replacement is discussed in (Liu et al., 2011). Non-standard tokens are frequently used in people’s daily lives, especially for text messages and social network websites, posing challenges to language processing tasks. (Liu et al., 2011) models the process of generating non-standard tokens from standard dictionary words by letter removal or substitution and proposes a letter transformation approach to handle non-standard tokens. Although used as a normalization method, this technique could also be applied in data augmentation tasks to transform non-standard tokens to standard form, or vice versa.
Token Addition. Token Addition refers to adding extra tokens to the original text, either introducing junk tokens as noise or enriching the sentences and making them more informative.
The random insertion, proposed in EDA (Wei and Zou, 2019), inserts a random synonym of a random word at a random position in the sentence. The words added in the EDA play the role of data noising and may carry either little or harmful information. As a variant of EDA, the technique of AEDA (Karimi et al., 2021) only inserts punctuation marks. It only adds limited noise and avoids the potentially negative effect of random insertion. INSNET (Lu et al., 2022) is an insertion-based text generation model that can sequentially insert new tokens into an incomplete sentence, and with consecutive insertion steps, the sentence is enriched and contains more information. LeCA (Chen et al., 2021) introduces lexical constraints to improve the performance of neural machine translation. In LeCA, a set of words is supposed to appear in the translation result. These words are called constraints whose embeddings are appended to those of the original tokens. The embeddings are then fed to the transformers’ encoding layer, and these constraint words are very likely to be included in the output text. A very similar research is (Hokamp and Liu, 2017), which also uses lexical constraints. The difference is that in LeCA, the constraints are given a soft suggestion, rather than being forced to appear as in (Hokamp and Liu, 2017).
Token Deletion. Random deletion (Wei and Zou, 2019) randomly removes some words from a sentence. In (Xie et al., 2017), noising techniques are used to avoid overfitting, and the randomly selected word is replaced by a blank placeholder token “_”, rather than simply being removed. In (Duan et al., 2023), apart from word dropout which is similar to random deletion, it also uses word blanking to replace a word with a special “¡BLANK¿” placeholder.
On Word Selection. Most methods randomly select the word to be replaced, added, and deleted. But some method considers the role a word plays in the sentence when deciding whether to impose perturbation. For example, TEXTFOOLER (Jin et al., 2020a) ranks words by their importance and replaces the important words with their synonyms.
Syntax-aware data augmentation (Duan et al., 2023) argues that most previous text data augmentation approaches are only computation-motivated and cannot assure the good quality of the generated text. Instead of simply modifying words with the same probability, Syntax-aware data augmentation calculates the probability of being modified for each word based on its syntactic role in the sentence. Specifically, it utilizes a dependency tree to calculate the importance of each word, i.e. the depth in the tree. A word far from the root is considered less important syntactically and is more likely to be removed (by means of dropout and blanking) or replaced during the augmentation process.
Table 4. Word selection strategy
Name Ref Augmentation Strategy
Random Role-based Learned
EDA (Wei and Zou, 2019)
- (Marivate and Sefara, 2020)
TEXTFOOLER (Jin et al., 2020a)
LetterTran (Liu et al., 2011)
AEDA (Karimi et al., 2021)
INSNET (Lu et al., 2022)
LeCA (Chen et al., 2021)
GBS (Hokamp and Liu, 2017)
- (Xie et al., 2017)
STA (Guo et al., 2022)
Instead of using a syntax tree to assess word importance, Selective Text Augmentation (STA) (Guo et al., 2022) focuses more on the statistical and semantic features of words to describe the role of words. It proposes two perspectives to view words in text classification tasks, namely, statistical correlation, which means how frequently a word co-occurs with a certain category, and semantic similarity, which measures the semantic overlap of the given words and the category label. Using the metrics as axes, STA divides the roles of words into four quadrants: gold, venture, bonus, and trivial. Words with different roles are treated differently when imposing augmentation operations on the sentence. For example, gold words will be kept in selective deletion and synonyms of gold words are added in selective insertion.
5.1.2. Structure-based Text Transformation
We categorize transformation that perturbs sentence structure into two types: Sentence Cropping that splits a sentence into many pieces, and Sentence Morphing that swaps the location of words or grammatical components.
Sentence Cropping. Borrowing the idea of image cropping, which selects segments from the images, Sentence Cropping (Şahin and Steedman, 2018) divides sentences into several small pieces. It often uses dependency trees to extract the relationships between subjects and objects of the sentence and produces multiple smaller sentences that focus on different parts of the original sentence. This operation breaks the syntactic solidity and semantic comprehensiveness of the entire sentence but preserves local syntactic tags and shallow semantic labels.
Sentence Morphing. These methods manipulate sentence structure by altering the position of words or terms and thus changing the syntax of the sentence.
Random swap (Artetxe et al., 2018; Wei and Zou, 2019) randomly chooses two words in a sentence and swaps their position. This perturbs the structure and adds noise to the text data across paragraphs. (Min et al., 2020) studies two types of syntactic augmentations: inversion and passivization. Inversion swaps the subject and object in a sentence, and the semantics might be different in most cases. Passivization converts a sentence to its passive form, and the meaning of the new sentence is typically consistent with the original one. Sentence rotation (Şahin and Steedman, 2018) borrows the idea of image rotation, where the image is rotated around its center. In sentence rotation, the root of the sentence is chosen as the center of the sentence, and other sentence fragments are connected to their center via the dependency tree. For some languages, such an operation is semantic-preserving, while for other languages the sentence’s meaning is subject to change. The DRNNs model (Xu et al., 2016) also utilizes the dependency tree and subject-object knowledge for data augmentation. It samples the shortest dependency path from a sentence’s dependency parse tree and then extracts two sub-paths, i.e. the subject-predicate path and the object-predicate path. Augmentation is done by changing the order of the two paths.
Table 5. Demonstration of some elementary operations for text data augmentation
Operation Sentence
Original I am an art student and I paint a lot of pictures. Many people pretend that they understand modern art.
Synonym Replacement (Wei and Zou, 2019) I am an art student and I draw a lot of pictures. Many people pretend that they understand contemporary art.
Similarity Replacement (Marivate and Sefara, 2020) I am an art student and I paint a lot of photos. Many people pretend that they understand modern literature.
Random Insertion (Wei and Zou, 2019) I am an art student and picture I paint a lot of pictures. Many appreciate people pretend that they understand modern art.
Punctuation Insertion (Karimi et al., 2021) I am an, art student and: I paint a lot of pictures. Many people? pretend that they. understand modern art.
Random Swap (Wei and Zou, 2019) I am an of student and I paint a lot art pictures. Many people they that pretend understand modern art.
Random Deletion (Wei and Zou, 2019) I am an art student and I paint a lot of pictures. Many people pretend that they understand modern art.
Placeholder Replacement (Xie et al., 2017) I am an art _ and I paint a lot of pictures. Many _ pretend that they understand modern art.
Inversion (Min et al., 2020) I am an art student and a lot of pictures paint me. Many people pretend that they understand modern art.
Passivization (Min et al., 2020) I am an art student and I paint a lot of pictures. Many people pretend that modern art are understood by them.
5.1.3. Text Value-structure Manipulation
This type of method perturbs both value and structure information of a sentence. And two typical methods are Back-translation Augmentation and Hierarchical Data Augmentation.
Back-translation Augmentation. Back-translation (Sennrich et al., 2016) means translating text data from one language to another, or a sequence of other languages, and finally translating back to the original language. The semantic information is preserved during the translation process, but the process can introduce some variance in syntax and wording. Such transformation can help the model learn more flexible expressions that convey the same meaning.
Hierarchical Data Augmentation. Hierarchical Data Augmentation (HDA) (Yu et al., 2019) augments text data at both word and sentence levels. It evaluates the roles of words and sentences by training a hierarchical attention network to obtain the attention values of the text data at both levels. The attention value is used to help decide the importance of components and augment the data by cropping and concatenating these components.
5.2. Multiple Text Augmentation
5.2.1. Value-based Text Mixture
Mixup (Zhang et al., 2017) has been a widely used data augmentation approach especially used for image data. However, unlike image data, whose coloring information is expressed by numerical values and can be interpolated at the input layer, it is hard to calculate an interpolation for the words. So, text data are usually interpolated at embedding layers.
Text Mixup. The adaptation of Mixup in sentence classification is studied in (Guo et al., 2019), which proposes wordMixup and senMixup. WordMixup interpolates a pair of word embeddings and then feeds the interpolation to the sentence encoder. SenMixup, on the other hand, interpolates a pair of sentence embeddings. The embedding vector produced by WordMixup or SenMixup is then passed to the softmax layer to produce the target distribution of the classification task. The use of mixup in text data augmentation has also been extensively studied in many research, such as nonlinear Mixup (Guo, 2020) and DoubleMix (Chen et al., 2022). A special example is GPT3Mix (Yoo et al., 2021), which makes use of the latest large language model to mix sentences. By inserting two source sentences in a prompt template that describes the instruction to the model, we can get a mixed sentence of high quality with a soft label.
5.2.2. Structure-based Text Combination
Some methods combine text data by merging text fragments from one sentence into another.
Text Fragment Merging. Substructure Substitution (Shi et al., 2021) and Subtree Swapping (Dehouck and Gómez-Rodríguez, 2020) both use tree structures to analyze the part of speech or role in the sentences and thus make reasonable swaps. Some simpler method like GECA (Andreas, 2019) aims to remove a fragment from a sentence and populate it with another fragment that occurs in some common environment.
5.3. Populational Text Augmentation
5.3.1. Vanilla Text Generation
We have discussed how augmented text data samples can be produced based on existing text materials. While modifying existing data is not the only solution to data augmentation. By leveraging various generative models, we can easily acquire high-quality text data on a large scale, and such abundance in data availability could greatly improve the performance of language models especially in terms of their generalization and robustness. Data augmentation by text generation previously relied on Traditional Models, while Large Language Models have now become the dominant technique.
Traditional Models. The use in data augmentation of some early generative language models, such as seq2seq (Prakash et al., 2016), VAE (Bowman et al., 2015), and CVAE (Zhou and Wang, 2017) has been investigated in (Malandrakis et al., 2019), which shows improvement in the accuracy of an extrinsic low resource classification task.
Large Language Models. Recent advances in large language models (LLM) and generative pre-training (GPT) (Radford et al., 2018) models have demonstrated unprecedented power in text understanding, classification, and generation tasks. Apart from supporting intelligent chat-bots that help people in their daily lives, the capacity of these state-of-the-art language generation models establishes a new paradigm of text data augmentation.
LAMBADA (Anaby-Tavor et al., 2020) is among the first attempts to use a state-of-the-art language generator to synthesize new text data for data augmentation, and its procedure is as follows. LAMBADA first trains a baseline classifier from a small dataset, and then it fine-tunes the GPT-2 (Radford et al., 2019) model. A set of labels is thereafter fed to the fine-tuned language model, which generates a large dataset of labeled sentences. These sentences are evaluated by the classifier trained prior according to its confidence score, and top-ranked sentences are retained to compose the synthesized dataset. Similarly, (Quteineh et al., 2020) integrates GPT-2 in its text data augmentation process and improves the performance of active learning on tiny datasets. It uses the language model to generate a tree of tokens as candidates and construct a sentence from them. The set of new sentences is repeatedly refined by active learning, and the outcome is an augmented text dataset.
Some other research also applies GPT-2 to effectively augment low-resource text datasets, such as (Yang et al., 2020) and (Edwards et al., 2022). Despite their success, both of them point out the need for more sophisticated methods to generate text data. (Yang et al., 2020) for example, argues that human-in-the-loop for relabeling might be rewarding. (Edwards et al., 2022) adopts human-in-the-loop for seed selection and further discusses the potential benefits of injecting expert knowledge for specialized domains. Notably, Data Boost (Liu et al., 2020) proposes a reinforcement learning framework to guide the process of text data augmentation with GPT-2 and succeeds in generating data of high quality and high similarity to the original data.
The rise of GPT-3 (Brown et al., 2020) and later LLMs is now providing more possibilities for text generative augmentation. Their use and improvements are studied in the latest research such as (Balkus and Yan, 2022) and (Cochran et al., 2023).
Table 6. A summary for text data augmentation
Tier 1: Quantity of source data Tier 2: Information type Augmentation method Typical literature
Individual Text Augmentation Value-based Transformation Invariant Replacement EDA (Wei and Zou, 2019) (EMNLP-IJCNLP 2019), (Marivate and Sefara, 2020) (CD-MAKE 2020), TEXTFOOLER (Jin et al., 2020a) (AAAI 2020), LetterTran (Liu et al., 2011) (ACL 2011), (Duan et al., 2023) (TASLP 2023), STA (Guo et al., 2022) (arXiv 2022)
Token Addition EDA (Wei and Zou, 2019) (EMNLP-IJCNLP 2019), AEDA (Karimi et al., 2021) (EMNLP 2021), INSNET (Lu et al., 2022) (NeurIPS 2022), LeCA (Chen et al., 2021) (IJCAI 2021), GBS (Hokamp and Liu, 2017) (ACL 2017), STA (Guo et al., 2022) (arXiv 2022)
Token Deletion EDA (Wei and Zou, 2019) (EMNLP-IJCNLP 2019), (Xie et al., 2017) (ICLR 2017), (Duan et al., 2023) (TASLP 2023), STA (Guo et al., 2022) (arXiv 2022)
Structure-based Transformation Sentence Cropping Sentence cropping (Şahin and Steedman, 2018) (EMNLP 2018), STA (Guo et al., 2022) (arXiv 2022)
Sentence Morphing EDA (Wei and Zou, 2019) (EMNLP-IJCNLP 2019), (Artetxe et al., 2018) (ICLR 2018), Sentence Rotation (Şahin and Steedman, 2018) (EMNLP 2018), Inversion & Passivization (Min et al., 2020) (ACL 2020), DRNNs (Xu et al., 2016) (COLING 2016)
Value-structure Transformation Back-translation Back-translation (Sennrich et al., 2016) (ACL 2016)
Hierarchical HDA (Yu et al., 2019) (IEEE Access 2019)
Multiple Text Augmentation Value-based Mixture Text Mixup wordMixup & senMixup (Guo et al., 2019) (arXiv 2019), Nonlinear Mixup (Guo, 2020) (AAAI 2020), DoubleMix (Chen et al., 2022) (COLING 2022), GPT3Mix (Yoo et al., 2021) (EMNLP 2021)
Structure-based Combination Text Fragment Merging Substructure Substitution (Shi et al., 2021) (IJCNLP 2021), GECA (Andreas, 2019) (arXiv 2019), Subtree Swapping (Dehouck and Gómez-Rodríguez, 2020) (COLING 2020)
Populational Text Augmentation Vanilla Generation Traditional Models (Malandrakis et al., 2019) (EMNLP-IJCNLP 2019)
LLMs LAMBADA (Anaby-Tavor et al., 2020) (AAAI 2020), (Quteineh et al., 2020) (EMNLP 2020), (Yang et al., 2020) (EMNLP 2020), (Edwards et al., 2022) (DaSH 2021), Data Boost (Liu et al., 2020) (EMNLP 2020), (Balkus and Yan, 2022) (Natural Language Engineering 2022), (Cochran et al., 2023) (ICAIE 2023)
6. Data Augmentation for Graph Data
A graph is represented by $\mathbf{G}=\{\mathcal{V},\mathcal{E}\}$, where $\mathcal{V}$ stands for the vertex set, and $\mathcal{E}$ stands for the edge set. The topology structure can be alternatively represented by an adjacency matrix $\mathbf{A}\in\mathbb{R}^{|\mathcal{V}|\times|\mathcal{V}|}$, where $\mathbf{A}_{ij}$ is $1$ if there is an edge connecting vertex $i$ and vertex $j$, and is 00 otherwise. In a deep graph learning scenario, the graph also contains a feature matrix $\mathbf{X}\in\mathbb{R}^{|\mathcal{V}|\times d}$ that represents the attributes of the nodes. Hence another representation for a graph is $\mathbf{G}=\{\mathbf{A},\mathbf{X}\}$. Graph data augmentation is to perturb the values in the feature matrix $\mathbf{X}$ of nodes or the topology structure represented by the adjacency matrix $\mathbf{A}$.
6.1. Individual Graph Augmentation
6.1.1. Value-based Graph Transformation
Graph nodes often contain rich attributes in the feature matrix. This matrix is generally similar to tabular data, so augmenting node attributes does not need to consider the specialty of graph data.
Node Attributes Augmentation. Various typical augmentation approaches on node attributes from random masking (You et al., 2020; Feng et al., 2020; Thakoor et al., 2021) to perturbing attributes with random noises (Yang et al., 2021) can be applied, as we do for common tabular data. These methods are unified in Eq. 6, where $\mathbf{M}\in\{0,1\}^{N\times N}$ is a binary mask, $\odot$ denotes the Hadamard product, and $\mathbf{N_{W}}\sim\mathcal{N}(0,\Sigma)$ is the Gaussian white noise.
(6) $\mathbf{\tilde{X}}=\mathbf{X}\odot\mathbf{M}+\mathbf{N}_{W}$
Some other methods are more sophisticated. For example, (Song et al., 2021) first learns the topology information of the graph by random walk, and uses the topology embedding together with the original features as inputs to a dual graph neural network. Similarly, LA-GNN (Liu et al., 2022d) applies a local augmentation mechanism. It samples features from nodes’ neighborhoods as additional information when augmenting the attributes. A different approach is proposed in DropMessage (Fang et al., 2023). Instead of directly modifying the attributes on the nodes, it applies random dropping in the message-passing phase.
6.1.2. Structure-based Graph Transformation
Graph topology structure has immense information representation capacity. Many augmentation methods designed for graph data focus on modifying the topology structure. Naive Topology Perturbation methods modify a single node or edge component or consider a set of nodes as a whole and modify the substructure. There is also Subgraphing that extracts subgraphs from the complete graph. Graph Diffusion, Graph Rewiring and Graph Structure Learning are more sophisticated techniques that focus on adjusting the edges to obtain a better graph.
Naive Topology Perturbation. Due to its straightforward idea and easy implementation, node dropping and addition are among the first applied augmentation approaches in graph learning tasks. As given in Eq. 7, Node dropping (You et al., 2020) randomly discards some nodes from the graph. Such an operation will also remove the attributes that these nodes carry and the edges on these nodes.
(7) $\tilde{\mathcal{V}}=\text{Drop}(\mathcal{V}),\qquad\tilde{\mathbf{G}}=\{\tilde{\mathbf{A}},\tilde{\mathbf{X}}\}=\{\mathbf{A}[\tilde{\mathcal{V}},\tilde{\mathcal{V}}],\mathbf{X}[\tilde{\mathcal{V}},:]\}$
In some special cases where node dropping is adopted, the edges on the removed nodes are not simply dropped. For example, in MoCL (Sun et al., 2021c), when dropping a general carbon, the two sides of the carbon atom are connected to ensure the completeness of the molecule. So, in MoCL, rather than being removed, the node is more of being “skipped”. As for node addition, an example is MPNN (Gilmer et al., 2017). It introduces a latent master node that is connected to all other nodes, which helps messages pass through the graph network.
An edge describes the relationship between two nodes, and by perturbing the edges, we also introduce diversity to the structural information of a graph. DropEdge (Rong et al., 2019) perturbs the edge set by randomly setting a portion of the adjacency matrix to zero. It first produces a sparse matrix from a random subset of the original edges, then derives the augmented edge set by subtracting the sparse matrix from the original adjacency matrix. Edge masking in (Thakoor et al., 2021) constructs a Bernoulli distribution, and randomly masks edges under that probability. FairDrop (Spinelli et al., 2021) computes a mask according to the network’s homophily, and use this mask to guide the edge dropping process to remove unfair connections. AD-GCL (Suresh et al., 2021) applies a GNN-augmenter to learn how to drop edges so that the mutual information of the augmented graph and that of the original graph is minimized.
Graph substructure perturbation is beneficial for molecular graphs. A molecule graph represents atoms with nodes and chemical bonds with edges. Functional groups, which often decide the chemical characteristics of molecules, can be represented by substructures. So, by wisely sampling substructures, the global semantics of a molecule graph can be well contained. MoCL (Sun et al., 2021c), for example, utilizes substructure information as one of its augmentation scheme. It performs graph augmentation by replacing functional groups, effectively introducing reasonable noises to the graph. GEAR (Liu et al., 2022e) also employs such technique. It augment molecule graphs by so-called environment replacement, which introduces noise from the environment subgraphs.
Subgraphing. Subgraphing is based on the idea that graph semantics are partially represented in local structures (You et al., 2020), and some subgraphing methods, such as those used in GCC (Qiu et al., 2020) and SUGAR (Sun et al., 2021a) use random walk to sample a subgraph from the original graph. GrpahCrop (Wang et al., 2020) applies a node-centric strategy to crop connectivity-preserved subgraphs and expand the dataset with these augmented subgraphs. For each subgraph, the rest of the graph is removed, but compared with edge dropping and node dropping methods, GraphCrop (Wang et al., 2020) better maintains the topology characteristics of the original graph. Similarly, SUBG-CON (Jiao et al., 2020) uses the Personalized PageRank algorithm to select the important nodes. NeuralSparse (Zheng et al., 2020) proposes to obtain the subgraph by learning. It uses a sparsification network to generate sparsified subgraphs.
Graph Diffusion. Traditional message passing in graph neural networks only leverages one-hop connections, while the edges of a real graph are often noisy or arbitrarily defined, and many relationships in the networks are indirectly contained in higher-order neighborhoods. To address this issue, GDC (Gasteiger et al., 2019), as in Eq. $\mathbf{\tilde{A}}=\sum_{k=0}^{\infty}\theta_{k}\mathbf{T}^{k}$ , proposes to calculate a series of transition matrices $\mathbf{T}^{k}$ by means such as random walk and combines these transition matrices with weights $\theta_{k}$ to obtain a new adjacency matrix $\mathbf{\tilde{A}}$ and thus generate a new graph that directly contains some node connection hidden in the original graph. The new graph can also be understood as being processed by a denoising filter because graph diffusion smooths out the neighborhood over the graph (Gasteiger et al., 2019). This technique is widely used by self-supervised approaches to transform the graph and generate different views for contrastive learning, such as MVGRL (Hassani and Khasahmadi, 2020), SelfGNN (Kefato and Girdzijauskas, 2021), and MERIT (Jin et al., 2021).
Graph Rewiring. The topology structure has significant influences on the performance of GNN models. Message passing plays an indispensable role in GNN models. Still, this process suffers from the over-squashing problem when information is spreading over long distances in a graph with the bottleneck (Topping et al., 2021). Also, the over-smoothing problem leads to GNNs’ performance degeneration on heterophilous graphs (Guo et al., 2023). Moreover, the graph’s robustness to perturbations poses the stability problem to GNN models (Kenlay et al., 2021). To address these issues, a series of methods referred to as Graph Rewiring is proposed, which adjust the structure of the graphs to help models learn better graph representations and yield better inference in downstream tasks (Guo et al., 2023).
Graph rewiring, as presented in Eq. 8, is initially proposed as randomly swapping edges (Schneider et al., 2011) to improve a network’s robustness and mitigate malicious attacks. Some later research such as smart rewiring (Louzada et al., 2013) takes the degrees of chosen nodes when swapping their edges. Stochastic Discrete Ricci Flow (Topping et al., 2021) is a curvature-based graph rewiring method, which in each step adds an edge that can improve the most negatively curved edge, and then removes the edge with the most positive curvature. Its results demonstrate the ability of graph rewiring to alleviate the over-squashing problem. DiffWire (Arnaiz-Rodríguez et al., 2022) introduces the CT-LAYER and GAP-LAYER in a GNN model, which respectively learns the commute times and spectral gap of the graph. It rewires the adjacency graph so that the graph’s spectral gap is minimized, and by doing so tries to predict the optimal graph structure for the downstream task. DHGR (Bi et al., 2022) first learns a similarity matrix from the original graph. Then it rewires the graph by adding edges between high-similarity node pairs and removing edges from low-similarity ones. Such an operation can effectively improve the homophily of the graph and thus improve GNN performance. HDHGR (Guo et al., 2023) is a method that follows DHGR (Bi et al., 2022). It computes multiple similarity matrices under different meta-paths and thus generates several rewired meta-path subgraphs.
(8) $\vspace{-2mm}\mathbf{\tilde{A}}=\overbrace{\mathbf{A}\odot(\mathbf{1}-\mathbbm{1}_{r})}^{\text{edge removal}}+\overbrace{(\mathbf{1}-\mathbf{A})\odot\mathbbm{1}_{r}}^{\text{edge addition}}$
Graph Structure Learning. Graph Neural Networks are subject to noisy, wrong, or incomplete graph structures. Similar to graph rewiring, graph structure learning (GSL) also aims at simultaneously optimizing graph structure and its corresponding representations (Zhu et al., 2021) to deal with the above-mentioned problem, however, GSL is often implemented in a more complicated manner and discussed in a broader context than graph rewiring. GSL applies various techniques to learn and model graph structures, and it is often depicted with a pipeline that includes iterative learning and updating (Zhu et al., 2021). Also, unlike graph rewiring, which only rewires a graph, GSL can be used to learn and construct a graph structure from non-graph datasets (Zhu et al., 2021), so that many models specific to graph datasets, e.g. GNNs can be trained on these data. Based on how the graph structure is modeled, GSL can be categorized into three classes: metric-based, neural, and direct modeling (Zhu et al., 2021).
Metric-based modeling evaluates the similarity between node pairs through some kernel functions, and in the new graph, the weights of the edges are decided by the similarities of the node pairs. Many early GSL methods fall into this group. For example, IDGL (Chen et al., 2020b) and HGSL (Zhao et al., 2021) both use cosine function for similarity learning to evaluate the optimal adjacency of nodes.
Neural approaches use neural networks for both graph learning and graph convolution tasks, as presented in the work of GLCN (Jiang et al., 2019), which introduces a simple single-layer neural network to learn a new graph from the original node attributes before the graph data is learned by the graph convolution module. PTDNet (Luo et al., 2021) focuses on dropping noisy, task-irrelevant edges from the graph structure. VIB-GSL (Sun et al., 2022) learns graph structure from an information theory perspective. SUBLIME (Liu et al., 2022f) proposes an unsupervised learning paradigm for GSL that benefits more types of downstream task.
Direct GSL approaches do not involve estimating edge weights from the vertex or generating a graph by training neural networks. Pro-GNN (Jin et al., 2020b) for example, directly updates the learned adjacency matrix itself during the training process. To avoid sub-optimal results, it jointly learns the graph structure and downstream GNN model. The learned graph and model parameters are updated sequentially in each training epoch. LDS-GNN (Franceschi et al., 2019) treats graph learning as a probabilistic distribution problem. When learning from the objectives, it updates the parameter of a Bernoulli distribution, and samples the learned graph from this distribution.
6.2. Multiple Graph Augmentation
6.2.1. Value-based Graph Mixture
Feature or label information from different nodes can be mixed by Graph Propagation, which is often used for tasks such as node classification.
Graph Propagation. Propagation refers to spreading information from one node to other nodes along the graph. This type of method can be divided into feature propagation and label propagation.
Feature propagation shares node features with other nodes. After feature vectors of some nodes are masked out, GRAND (Feng et al., 2020) further augments the node features by randomly propagating them through edges. MV-GCN (Yuan et al., 2021) first creates three different views from the original graph and propagates the features of the three views respectively before their representations are combined.
Label propagation algorithm (LPA) (Zhu and Ghahramani, 2002) propagates labels from labeled nodes to all unlabeled nodes according to their proximity through areas with dense unlabeled nodes. AutoGRL (Sun et al., 2021b) applies LPA in their semi-supervised framework to label more nodes utilizing a few ground truth nodes. (Huang et al., 2020) performs two types of label propagation, correcting the base predictions and smoothing the final prediction, respectively. Though not in a typical graph-based scenario, (Iscen et al., 2019) constructs a nearest neighbor graph on the classifier’s manifolds. The affinity matrix, which is the analogy to the adjacency matrix in this case, is constructed based on the $k$ nearest neighbors of each node’s embeddings, and this matrix decides how labels propagate.
6.2.2. Graph Value-structure Mixutre
Graph Mixup refers to mixing graph nodes or graph instances. This often involves both node attribute values and the topology structure information.
Graph Mixup. The technique of mixup has been extensively applied in various augmentation approaches for image and text tasks. By interpolating feature vectors and labels, mixup expands the feature space of training data and effectively improves the model’s robustness and generalization ability. However, the application of mixup in graph data augmentation is challenged by the special structure of graph datasets. Due to the irregular connectivity and various topological structures of graph data, it is hard to directly combine nodes or graphs, as we do when combining two images.
For node-level mixup, one solution to this problem is temporarily ignoring the edges and mixing node features in latent space. For example, GraphMix (Verma et al., 2021) interpolates the hidden states and labels of input nodes to train a better FCN layer, whose parameters are shared with the GNN layer. A topology-involved node mixup method is proposed in (Wang et al., 2021). It randomly pairs nodes and includes the nodes’ local topology representations when interpolating them. Notwithstanding the topology information’s involvement in (Wang et al., 2021), it dose not mentions the construction of edges for the new node. S-Mixup (Structural Mixup) (Kim et al., 2023) also considers neighboring edges when pairing nodes but works in the input space and discusses how to connect the node with the original graph with edges. It matches nodes based on their classification confidence of a GNN classifier and selects edges with high gradients to connect with the mixed node.
(Wang et al., 2021) also proposes a method for graph-level mixup. It uses GNNs to embed the complex and irregular graph structures and conduct mixup in the embedding space. However, this method cannot mix graphs in the input layer. To directly mix two inputs at the graph level, ifMixup (Guo and Mao, 2021) adds dummy nodes so that two graphs have the same number of nodes. These dummy nodes have all-zero feature vectors and are not connected to any other nodes, or in other words, the weights of edges on dummy nodes are zeros. When mixing the graph, the interpolation of a normal, one-hot labeled edge and a zero-weighted edge yields an edge with a soft label. Similarly, S-Mixup (Graph Mixup with Soft Alignments) (Ling et al., 2023) also aligns graphs first before mixing them. It obtains an assignment matrix by a soft assignment and uses the assignment matrix to transform a graph so that every node corresponds with a node in the other graph. Another approach that mixes graph-structured datasets at the graph level and input space is Graph Transplant (Park et al., 2022). It extracts a salient subgraph from the source graph and a random subgraph from the destination graph. The augmented graph is produced by transplanting the source subgraph to the destination subgraph and adaptively interpolating the labels of the original graphs. Compared with the above-mentioned approaches, G-Mixup (Han et al., 2022) proposes a more sophisticated way to interpolate graph data. It first uses graphs within the same class to estimate a graphon, i.e. graph generator:
(9) $\mathbf{\mathcal{G}}\rightarrow W_{\mathbf{\mathcal{G}}},\mathbf{\mathcal{H}}\rightarrow W_{\mathbf{\mathcal{H}}}$
The graphon $W:[0,1]^{2}\rightarrow[0,1]$ is a matrix, where $W(i,j)$ describes the probability of an edge between nodes $i$ and $j$, and graphs of the same class are considered to be generated from the same graphon. Graphons of two classes of graphs are then interpolated to produce a mixed graphon:
(10) $W_{\mathbf{\mathcal{I}}}=\lambda W_{\mathbf{\mathcal{G}}}+(1-\lambda)W_{\mathbf{\mathcal{H}}}$
By sampling from the graphon mixup $W_{\mathbf{\mathcal{I}}}$, we acquire a series of graphs with the same soft labels:
(11) $\mathbf{y}_{\mathbf{\mathcal{I}}}=\lambda\mathbf{y}_{\mathbf{\mathcal{G}}}+(1-\lambda)\mathbf{y}_{\mathbf{\mathcal{H}}}$
At the graphon estimation phase, G-Mixup aligns the original node features and obtains the graphon node features average pooling, which are used as node features of the generated graphs. DAGAD (Liu et al., 2022b) mixes graph representations by performing random permutations and it shows that data augmentation is very helpful in scenarios such as graph anomaly detection.
6.3. Populational Graph Augmentation
6.3.1. Vanilla Graph Generation
Despite that Variational Autoencoders (VAEs) and Generative Adversarial Networks (GANs) can all be used to generate graph data, they face many problems in practice. VAE models require things like a massive graph-matching process, and GANs on graphs easily fall into mode collapse (Liu et al., 2023a). So, graph data augmentation resort to another recently proposed technique, namely, Generative Diffusion Models.
Generative Diffusion Models
The denoising diffusion model (Ho et al., 2020) is a generative paradigm that has become popular in recent years. Though not yet widely applied as an approach to graph data augmentation, the Generative Diffusion Model boasts enormous potential in graph generation and can possibly be adopted as a promising data augmentation technique. To implement diffusion models on graph data, some properties of the graph data must be properly treated. For example, diffusion models often rely on a continuous Gaussian noise process, but the graph structures are discrete (Zhang et al., 2023). DiGress (Vignac et al., 2023) designs a discrete denoising diffusion model for graph generation and achieves state-of-the-art performance on molecular and non-molecular datasets. The use of Generative Diffusion Models for data augmentation especially benefits studies on molecules, proteins, and materials, whose data are often in the form of graph instances.
Table 7. A summary for graph data augmentation
Tier 1: Quantity of source data Tier 2: Information type Augmentation method Typical literature
Individual Graph Augmentation Value-based Transformation Node Attributes Augmentation Attribute Masking (You et al., 2020) (NeurIPS 2020), GRAND (Feng et al., 2020) (NeurIPS 2020), BGRL (Thakoor et al., 2021) (ICLR 2021), GASSL (Yang et al., 2021) (NeurIPS 2021), Topological Regularization (Song et al., 2021) (arXiv 2021), LA-GNN (Liu et al., 2022d) (ICML 2022), DropMessage (Fang et al., 2023) (AAAI 2023)
Structure-based Transformation Naive Topology Perturbation Node Dropping & Edge perturbation (You et al., 2020) (NeurIPS 2020), MPNN (Gilmer et al., 2017) (ICML 2017), BGRL (Thakoor et al., 2021) (ICLR 2021), DropEdge (Rong et al., 2019) (ICLR 2019), AD-GCL (Suresh et al., 2021) (NeurIPS 2021), FairDrop (Spinelli et al., 2021) (TAI 2021)
Subgraphing Subgraph (You et al., 2020) (NeurIPS 2020), GCC (Qiu et al., 2020) (KDD 2020), SUGAR (Sun et al., 2021a) (WWW 2021), GrpahCrop (Wang et al., 2020) (arXiv 2020), SUBG-CON (Jiao et al., 2020) (ICDM 2020), NeuralSparse (Zheng et al., 2020) (ICML 2020), MoCL (Sun et al., 2021c) (KDD 2021), GEAR (Liu et al., 2022e) (KDD 2021)
Graph Diffusion GDC (Gasteiger et al., 2019) (NeurIPS 2019), MVGRL (Hassani and Khasahmadi, 2020) (ICML 2020), SelfGNN (Kefato and Girdzijauskas, 2021) (arXiv 2021), MERIT (Jin et al., 2021) (arXiv 2021)
Graph Rewiring (Schneider et al., 2011) (PNAS 2011), Smart Rewiring (Louzada et al., 2013) (J. Complex Netw. 2013), SDRF (Topping et al., 2021) (ICLR 2021), DiffWire (Arnaiz-Rodríguez et al., 2022) (LoG 2022), DHGR (Bi et al., 2022) (arXiv 2022), HDHGR (Guo et al., 2023) (WWW 2023)
Graph Structure Learning IDGL (Chen et al., 2020b) (NeurIPS 2020), HGSL (Zhao et al., 2021) (AAAI 2021), GLCN (Jiang et al., 2019) (CVPR 2019), PTDNet (Luo et al., 2021) (WSDM 2021), VIB-GSL (Sun et al., 2022) (AAAI 2022), SUBLIME (Liu et al., 2022f) (WWW 2022), Pro-GNN (Jin et al., 2020b) (KDD 2020), LDS-GNN (Franceschi et al., 2019) (ICML 2019)
Multiple Graph Augmentation Value-based Mixture Propagation Label propagation (Zhu and Ghahramani, 2002) (2002), GRAND (Feng et al., 2020) (NeurIPS 2020), MV-GCN (Yuan et al., 2021) (CIKM 2021), AutoGRL (Sun et al., 2021b) (IJCNN 2021), C&S (Huang et al., 2020) (ICLR 2020), (Iscen et al., 2019) (CVPR 2019)
Value-structure Mixture Graph Mixup GraphMix (Verma et al., 2021) (AAAI 2021), (Wang et al., 2021) (WWW 2021), Structural Mixup (Kim et al., 2023) (CIKM 2023), ifMixup (Guo and Mao, 2021) (arXiv 2021), Graph Mixup with Soft Alignments (Ling et al., 2023) (arXiv 2023), Graph Transplant (Park et al., 2022) (AAAI 2022), G-Mixup (Han et al., 2022) (ICML 2022), DAGAD (Liu et al., 2022b) (ICDM 2022)
Populational Graph Augmentation Vanilla Generation Diffusion Models DiGress (Vignac et al., 2023) (ICLR 2023)
7. Data Augmentation for Tabular Data
A table is made up of cells, which store numerical or categorical values that represent attributes of the described entities. Cells are organized into columns and rows, but this structural information is trivial, and switching the sequence of the columns and rows does not affect the information conveyed by the table. So, most data augmentation methods leverage the value information.
7.1. Individual Tabular Augmentation
7.1.1. Value-based Table Transformation
Tabular data is much simpler compared with other modalities, so it is hard to design complex augmentation methods that make use of spatial or sequential relationships within a data sample as we do for image or text data, and most data augmentation for tabular data perturbs the value stored in the table cells. Table Masking is a simple technique often applied for tabular data augmentation. These corrupted data are often used with the original data under a self-supervised learning framework to learn stronger representations of the dataset. Another approach to tabular data augmentation borrows ideas from Feature Engineering. Some work on Feature Engineering design automatic feature selection and transformation frameworks to generate optimal feature space, and the operations can also be used to augment tabular datasets.
Table Masking. VIME (Yoon et al., 2020) uses a mask generator to corrupt features. In a self-supervised scenario, two pretext tasks respectively reconstruct the original data and estimate the mask vector. The model updates its weights based on the reconstruction loss and the mask vector estimation loss of the two pretext tasks. VIME also considers a semi-supervised scenario. In that case, multiple masks are generated and applied to the feature of the unlabeled samples. Despite not being labeled, the model predictions on these corrupted samples should be similar, and thus a consistency loss is calculated. For labeled samples, a supervised loss is calculated based on the prediction. The two losses jointly contribute to the back-propagation phase. SCARF (Bahri et al., 2021) applies data augmentation in a contrastive learning framework. Its corruption process is similar to that of VIME but it uses InfoNCE loss to evaluate the similarity between the embeddings of the corrupted and original views and then update the model accordingly. MET (Majmundar et al., 2022) works in a pattern similar to VIME and SCARF, which calculates the reconstruction loss of a partially masked sample. MTR (Onishi and Meguro, 2023) randomly replaces part of the tokenized embeddings, instead of working directly in the input space.
Feature Engineering. (Khurana et al., 2018) continuously generates new features and records them in a transformation graph. It uses reinforcement learning to select the feature set with best cross-validation performance. Some methods formulate the feature selection problem into a multi-agent reinforcement learning task (Liu et al., 2019, 2021a; Fan et al., 2020). They assign each feature to an agent, and the agents learn from the reward whether to select or unselect the features. However, the use of multi-agent reinforcement learning increases the computational burden and hardware cost, and some single-agent frameworks are further proposed for this problem (Liu et al., 2021b, 2023b). In these methods, only one agent is employed to traverse the whole feature set using the Monte Carlo algorithm and early stopping. Compared to the above-mentioned automatic feature selection methods, RAFT (Wang et al., 2022) and GRFG (Xiao et al., 2023a) emphasize feature transformation and generation. Unlike latent representation learning methods, they use arithmetic operations to generate new features, providing better traceability and explainability. GAINS (Xiao et al., 2023b) and MOAT (Wang et al., 2023) extend feature selection and transformation to continuous space. In GAINS (Xiao et al., 2023b), the selected feature subset is encoded into a continuous embedding space. It searches for better embeddings in the learned space and uses the decoder to generate an optimal selection of the dataset. MOAT (Wang et al., 2023) adapts this process to a feature transformation scenario, where features go through arithmetic operations transformation before the continuous embedding and searching stage, and the decoded features are also processed by a series of transformations. TFWT (Zhang et al., 2024) computes weights for samples and features, and the weights are combined with the data for later tasks. Although these studies focus on feature engineering, their way of thinking and application of reinforcement learning is very enlightening for data augmentation tasks.
7.1.2. Structure-based Table Transformation
Table Subsetting
Inspired by image cropping (Chen et al., 2016; Yan et al., 2013), SubTab (Ucar et al., 2021) divides a table into multiple subsets by splitting its columns. Each subset contains only a percentage of all features. The model is forced to learn the representation of the samples by these features and is required to reconstruct all the features from the subset.
7.2. Multiple Tabular Augmentation
7.2.1. Value-based Tabular Mixture
Mixup is the technique to interpolate samples to produce new data. Tabular data, due to its will-aligned structure, is suitable for this augmentation approach.
Table Mixup. SMOTE (Chawla et al., 2002) is among the first works to study data interpolation, even before the proposal of Mixup (Zhang et al., 2017). It addresses the problem of class imbalance by oversampling from the minority class. The samples are selected within the $k$ nearest neighbors in the feature space and new synthetic data are produced by a random interpolation operation. Since tabular features are often nominal, two variants, namely SMOTE-NC and SMOTE-N, are also proposed along with the initial version of SMOTE (Chawla et al., 2002). As argued in (Darabi and Elor, 2021), SMOTE is subject to unrealistic samples when interpolating in a sparse feature space. To solve this problem, TAEI (Darabi and Elor, 2021) first encodes the sparse data into a dense latent space, interpolates the new data, and then decodes them to obtain the original data plus the augmented data. Contrastive Mixup (Darabi et al., 2021) applies mixup for tabular data in a self- and semi-supervised learning framework. SAINT (Somepalli et al., 2021) combines CutMix (Yun et al., 2019) with contrastive learning to learn from tabular data. ExcelFormer (Chen et al., 2023) proposes two schemes, namely Feat-Mix and Hidden-Mix. They interpolate data in the input layer and the hidden layer, respectively.
7.2.2. Struture-based Tabular Combination
Contrary to Individual Tabular Augmentation, which lacks structural information to use, Multiple Tabular Augmentation suffers from strict and fixed table structure. Columns differ in their nomenclatures and representations, so it is hard to combine and learn from multiple distinct tables. If these restrictions are relaxed, models can learn across different tables, be more transferable, and better predict unseen tables.
Structure Relaxation. TransTab (Wang and Sun, 2022) suggests that multiple tables could share partially overlapped columns in the real world, but traditional methods often fail to fully utilize the data due to the removal of non-overlapping columns and mismatched samples when cleaning the data. TransTab first divides the table into three parts, each containing categorical, binary, and numerical data. These three types of data undergo different operations, are all converted to embeddings, and are again integrated into a new and encoded table. The processed tables can be joined in many ways, depending on the application scenarios. The augmented table contains information from many original tables and helps the model learn across them.
7.3. Populational Tabular Augmentation
7.3.1. Vanilla Tabular Generation
An Autoencoder (AE) learns from the dataset and extracts the essential structure, which can be used to generate a more robust dataset. Generative Adversarial Network (GAN) is another powerful method that can produce more augmented data and is especially useful for oversampling and alleviating the problem of class imbalance.
Autoencoder. (Delgado and Oyedele, 2021) tries out four types of autoencoder, concluding that all of these autoencoders can generate data sets that produce better task performance, while the variational autoencoder (VAE) provides the most robust result. Some other works (Fang et al., 2022; Tran et al., 2019) also use VAE as their data generation approach. SDAT (Fang et al., 2022) adds extra noise to the latent space of its VAE to obtain augmented samples. These samples, along with the original ones, are then trained under a semi-supervised framework. (Tran et al., 2019) combines VAE with active learning. The samples are iteratively selected by active learning, labeled by an oracle, reconstructed by the VAE, and discriminated and classified.
Generative Adversarial Networks. Table-GAN (Park et al., 2018) mainly considers the privacy issue, and experiments show that the augmented table given by GAN is compatible with downstream machine learning models and produces similar performance. CTGAN (Xu et al., 2019) focuses on the class imbalance problem and proposes a conditional generator that accounts for the imbalance in categorical columns. (Engelmann and Lessmann, 2021) shares similar goal and method with CTGAN (Xu et al., 2019) but employs a more sophisticated loss function that better incorporates the discriminator and the auxiliary classifier. ITS-GAN (Chen et al., 2019) proposes a framework that aims to preserve functional dependencies in tables. It models functional dependencies with autoencoders and includes a term to characterize the set of functional dependency constraints in the generator loss function. Arguing that GANs are difficult to train on small datasets, ConvGeN (Schultz et al., 2023) designed a new mechanism that requires the GAN’s generator and discriminator to cooperate with each other, that the generator generates batches of convex combination of minority neighborhood samples and the discriminator classifies these augmented minority samples against majority batches.
7.3.2. Exogenous Tabular Generation
The potential of tabular data is limited by its simple nature compared with data from other modalities. Due to the lack of spatial or relational relationships, many popular augmentation approaches and network configurations are not suitable for tabular datasets. A solution is to construct and introduce relational structure based on the tabular dataset.
Relational Structure Construction. GOGGLE (Liu et al., 2022c) proposes a solution to this problem. It learns the relational structure and corresponding functional relationships from tabular datasets and uses this basis to generate new samples. In other words, it learns a graph from the features in the table and obtains augmented data from that graph. PET (Du et al., 2022) depicts multiple tabular data instances are as a hyper graph and use the labels to help construct hyper edges. These methods are based on external assumptions and introduces extra structural information beyond the tabular dataset.
Table 8. A summary for tabular data augmentation
Tier 1: Quantity of source data Tier 2: Information type Augmentation method Typical literature
Individual Tabular Augmentation Value-based Transformation Table Masking VIME (Yoon et al., 2020) (NeurIPS 2020), SCARF (Bahri et al., 2021) (ICLR 2021), MET (Majmundar et al., 2022) (arXiv 2022), MTR (Onishi and Meguro, 2023) (arXiv 2023)
Feature Engineering RL1 (Khurana et al., 2018) (AAAI 2018), MARFS (Liu et al., 2019) (KDD 2019), AutoFS (Fan et al., 2020) (ICDM 2020), MCRFS (Liu et al., 2021b) (ICDM 2021), RAFT (Xiao et al., 2023a) (SDM 2023), GRFG (Wang et al., 2022) (KDD 2022), GAINS (Xiao et al., 2023b) (ICDM 2023), MOAT (Wang et al., 2023) (NeurIPS 2023), TFWT (Zhang et al., 2024) (IJCAI 2024)
Structure-based Transformation Table Subsetting SubTab (Ucar et al., 2021) (NeurIPS 2021)
Multiple Tabular Augmentation Value-based Mixture Tabular Mixup SMOTE (Chawla et al., 2002) (JAIR 2002), TAEI (Darabi and Elor, 2021) (arXiv 2021), Contrastive Mixup (Darabi et al., 2021) (arXiv 2021), SAINT (Somepalli et al., 2021) (arXiv 2021), ExcelFormer (Chen et al., 2023) (arXiv 2023)
Structure-based Combination Structure Relaxation TransTab (Wang and Sun, 2022) (NeurIPS 2022)
Populational Tabular Augmentation Vanilla Generation Autoencoder (Delgado and Oyedele, 2021) (Applied Soft Computing 2021), SDAT (Fang et al., 2022) (CIKM 2022), (Tran et al., 2019) (ICML 2019)
Generative Adversarial Networks Table-GAN (Park et al., 2018) (arXiv 2018), CTGAN (Xu et al., 2019) (NeurIPS 2019), (Engelmann and Lessmann, 2021) (Expert Systems with Applications 2021), ITS-GAN (Chen et al., 2019) (IJCAI 2019), ConvGeN (Schultz et al., 2023) (Pattern Recognition 2023)
Exogenous Generation Relational Structure Construction GOGGLE (Liu et al., 2022c) (ICLR 2022), PET (Du et al., 2022) (NeurIPS 2022)
8. Data Augmentation for Time-series Data
Time-series data represent sequential values within a period, and each timestamp is considered as an elemental component. Data augmentation methods for time-series data either perturb the value at each timestamp or perturb the sequence of timestamps.
8.1. Individual Time-series Augmentation
8.1.1. Value-based Time-series Transformation
A type of simple and straightforward value-based method is Value Perturbation, which changes the value represented by a sequence. There is also another type of method called Decomposition, which decomposes the value of a sequence into different parts that express the sequential values from different aspects.
Value Perturbation. Value Perturbation augments time-series data by altering the value on the Y-axis. It often perturbs the amplitude at a certain timestamp or in a certain frequency channel.
One of the most simple but popular augmentation methods is to add random noise to the value at each timestamp. For time-series data, this is referred to as jittering. In (Um et al., 2017), jittering is used to simulated additive noise $\epsilon$ of wearable sensors:
(12) $\tilde{x}(\epsilon)=\{x_{1}+\epsilon_{1},\cdots,x_{t}+\epsilon_{t},\cdots,x_{T}+\epsilon_{T}\}$
Scaling (Um et al., 2017) alters the data by multiplying the magnitude of the entire sequence with a scalar $\alpha$.
(13) $\tilde{x}(\alpha)=\{\alpha x_{1},\cdots,\alpha x_{t},\cdots,\alpha x_{T}\}$
Magnitude warping (Um et al., 2017) is similar to scaling, but it uses a smooth curve to warp the sequence so that magnitude at different timestamps is scaled with different parameters.
(14) $\tilde{x}(\alpha)=\{\alpha_{1}x_{1},\cdots,\alpha_{t}x_{t},\cdots,\alpha_{T}x_{T}\}$
Rotation (Um et al., 2017), or flipping, assumes the data is symmetric with respect to the X-axis, and makes the data upside down by applying a rotation matrix $R$.
(15) $\tilde{x}(R)=\{Rx_{1},\cdots,Rx_{t},\cdots,Rx_{T}\}$
Apart from the amplitude, the phase spectrum provides another option for data augmentation. RobustTAD (Gao et al., 2020) converts the time-series data from the time domain to the frequency domain, as shown in Eq. 16, and then increases all phases $\theta(\omega_{k})$ by a small perturbation.
(16) $F(\omega_{k})=\frac{1}{N}\sum^{N-1}_{t=0}{x_{t}e^{-j\omega_{k}t}}=A(\omega_{k})exp[j\theta(\omega_{k})]$
Decomposition. Empirical Mode Decomposition (EMD) (Huang et al., 1998) is used to produce a set of intrinsic mode functions (IMFs). This decomposition relies on extracting the energy linked to different time scales that are inherent to the data. EMD has been utilized as a method of data augmentation in (Nam et al., 2020) for the purpose of categorizing impact noise in vehicles. Another way to decompose time-series data is STL (Cleveland et al., 1990), which decomposes signals into seasonal, trend, and remainder:
(17) $y_{t}=\tau_{t}+s_{t}+r_{t},\quad t=1,2,\cdots N$
Other works such as (Wen et al., 2019) and (Gao et al., 2020) focus on improving robustness with the help of STL components.
8.1.2. Structure-based Time-series Transformation
Unlike Value Perturbation, sequence perturbation is implemented on the X-axis, either on the time domain or the frequency domain. Window Slicing splits the time-series data into pieces. It is also the basis of many Sequence Morphing methods because they apply varied operations on these slices.
Window Slicing. Inspired by image cropping, which is widely used for image data augmentation, MCNN (Cui et al., 2016) proposes window slicing. It extracts slices from the entire sequence as new training samples. Each slice is assigned with the label of the original sequence.
Sequence Morphing. Based on Window Slicing, (Le Guennec et al., 2016) extends or compresses the generated spans, which is called window wrapping. This is equivalent to speeding up or slowing down a span of the sequence. Vowel stretching (Nagano et al., 2019) is a special kind of window wrapping technique. It only perturbs the sequence by prolonging the vowels. This is meant to be consistent with how children speak. (Um et al., 2017) propose to randomly permeate these slices and concatenate them as a long sequence. Similar processing is also called Random Shuffling and is used in (Steven Eyobu and Han, 2018). VTLP (Jaitly and Hinton, 2013) applies window wrapping in the frequency domain. It adds a small distortion to the central frequency. SpecAugment (Park et al., 2019) also proposes to augment the frequency domain. It generates masks of length sampled from a uniform distribution and uses each of them to mask consecutive frequency channels.
8.2. Multiple Time-series Augmentation
8.2.1. Value-based Sequence Mixing
Most methods that combine information from different sequences only use the values, and they are referred to as sequence averaging or sequence mixing.
For example, DTW Barycenter Averaging (DBA) (Petitjean et al., 2011) revolves around the averaging problem for Dynamic Time Warping. It develops a global technique for computing the average of a set of sequences and avoids using iterative pairwise averaging. In Weighted-DBA (Forestier et al., 2017), instead of each time series contributing equally to the final average, some can contribute more than others, so that an infinite number of new examples from any set of given time series can be generated. (DeVries and Taylor, 2017a) applies both interpolation and extrapolation in the feature space to augment time-series data. It first uses a sequence autoencoder to learns a feature space from unlabeled data. The data is then encoded and augmented with additive noises, interpolation, and extrapolation. Finally, the data in feature spaces is decoded and classified. Extending sequence mixing to the frequency domain, (Nanni et al., 2020) applies the Equalized Mixture Data Augmentation (EMDA) to create new data. It computes the weighted average of two randomly chosen spectrograms with the same label.
8.3. Populational Time-series Augmentation
8.3.1. Vanilla Time-series Generation
GANs are widely used to generate virtual time-series data.
Generative Adversarial Networks. (Esteban et al., 2017) uses an RGAN and an RCGAN to produce realistic medical time-series data. TimeGAN (Yoon et al., 2019) combines supervised and unsupervised losses to train its autoencoding components and adversarial components. (Donahue et al., 2018) propose WaveGAN and SpecGAN to generate time-series data. WaveGAN is used on waveform, while SpecGAN is used on spectrogram. SpecGAN demonstrates how GAN is applied to generate time-series data on the frequency domain
8.3.2. Exogenous Time-series Generation
Some methods view time-series data outside the existing data and from a statistical perspective. They are based on human knowledge on statistics and can predict how the value changes in future timestamps.
Statistical Generation. (Smyl and Kuber, 2016) uses a statistical algorithm called LGT (Local and Global Trend) to forecast paths of the time-series data and samples from them to get extra data. GRATIS (Kang et al., 2020) employs mixture autoregressive (MAR) models to generate collections of time series and examine the variety and extent of the produced time series within a feature space for time series.
Table 9. A summary for time-series data augmentation
Tier 1: Quantity of source data Tier 2: Information type Augmentation method Typical literature
Individual Time-series Data Augmentation Value-based Transformation Value Perturbation Jittering, Scaling, Magnitude Wrapping, and Rotation (Um et al., 2017) (ICMI 2017), Phase Perturbation (Gao et al., 2020) (arXiv 2020)
Decomposition EMD (Nam et al., 2020) (ICASSP 2020), STL (Cleveland et al., 1990) (J. Off. Stat 1990), RobustSTL (Wen et al., 2019) (AAAI 2019), RobustTAD (Gao et al., 2020) (arXiv 2020)
Structure-based Transformation Window Slicing Window Slicing (Cui et al., 2016) (arXiv 2016)
Sequence Morphing Window Wrapping (Le Guennec et al., 2016) (ECML-PKDD 2016), Vowel Stretching (Nagano et al., 2019) (ICASSP 2019), Permutation (Um et al., 2017) (ICMI 2017), Random Shuffling (Steven Eyobu and Han, 2018) (Sensors 2018), VTLP (Jaitly and Hinton, 2013) (ICML 2013), SpecAugment (Park et al., 2019) (arXiv 2019)
Multiple Time-series Data Augmentation Value-based Mixture Sequence Mixing DTW Barycenter Averaging (Petitjean et al., 2011) (Pattern Recognition 2011), Weighted-DBA (Forestier et al., 2017) (ICDM 2017), (DeVries and Taylor, 2017a) (arXiv 2017), EMDA (Nanni et al., 2020) (Ecological Informatics 2020)
Populational Time-series Data Augmentation Vanilla Generation GAN (Esteban et al., 2017) (arXiv 2017), TimeGAN (Yoon et al., 2019) (NeurIPS 2019), WaveGAN & SpecGAN (Donahue et al., 2018) (ICLR 2018)
Exogenous Generation Statistical Generation Local and Global Trend (Smyl and Kuber, 2016) (ISF 2016), MAR (Kang et al., 2020) (Statistical Analysis and Data Mining 2020)
9. Discussion
Influence of Data Augmentation Operations. Some data augmentation methods aim to produce noisy or even “wrong” data to improve the AI models’ robustness by learning those low-quality or adversarial samples; in contrast, some are designed to improve data quality by removing noise or error from the data, so that the models can learn a clean representation of the data; some other methods do not intend to affect data quality, and they only make some semantic-preserving perturbation to the data, enlarge the dataset size, and cover more area in the feature space.
Evaluation Metrics Specific to Data Augmentation. Most research on data augmentation uses the models’ performance on downstream tasks to evaluate the effect of data augmentation methods. However, this criterion depends on too many extraneous factors and may fail to reflect the influence of data augmentation methods on the data. So, there is need for a set of evaluation metrics that can assess specific considerations of data augmentation, such as how much diversity is introduced and data distribution in the feature space changes.
Harnessing Latest Techniques. Large language models possess outstanding ability in manipulating text data, but up to now, most text data augmentation methods relies on symbolic analysis and rules. The question of how to utilize advanced techniques, such as LLM, to perform data augmentation tasks remains open and is rewarding to discover.
10. Conclusion
This survey presents a comprehensive summary of data augmentation techniques across five data modalities and proposes a modality-independent taxonomy from a data-centric perspective, which focuses on where the augmented data is derived from. It further enumerates related research papers on these data augmentation methods and annotates them with descriptive information in detail.

### v3

@@
 A Comprehensive Survey on Data Augmentation
-Abstract.
+Abstract
 Data augmentation is a series of techniques that generate high-quality artificial data by manipulating existing data samples.
@@
 Existing literature surveys only focus on a certain type of specific modality data, and categorize these methods from modality-specific and operation-centric perspectives, which lacks a consistent summary of data augmentation methods across multiple modalities and limits the comprehension of how existing data samples serve the data augmentation process.
-To bridge this gap, we propose a more enlightening taxonomy that encompasses data augmentation techniques for different common data modalities.
-Specifically, from a data-centric perspective, this survey proposes a modality-independent taxonomy by investigating how to take advantage of the intrinsic relationship between data samples, including single-wise, pair-wise, and population-wise sample data augmentation methods.
-Additionally, we categorize data augmentation methods across five data modalities through a unified inductive approach.
+To bridge this gap, this survey proposes a more enlightening taxonomy that encompasses data augmentation techniques for different common data modalities by investigating how to take advantage of the intrinsic relationship between and within instances.
+Additionally, it categorizes data augmentation methods across five data modalities through a unified inductive approach.
+Index Terms:
 Data augmentation, Data-centric taxonomy, Multi-modality
-copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXX
-1. Introduction
-With the rapid development of Artificial Intelligence (AI) in the past decades, AI methods have shown their superiority over human beings and other traditional approaches across most tasks, from universal tasks in our daily lives such as image recognition (He et al., 2016) and text translation (Bahdanau et al., 2015) to complicated scientific research tasks such as protein structure prediction (Jumper et al., 2021) and weather forecasting (Bi et al., 2023).
-Recently, powered by Latent Diffusion Models (Rombach et al., 2022) and Transformer (Vaswani et al., 2017), Stable Diffusion (Rombach et al., 2022) and ChatGPT (OpenAI, 2023) have greatly changed human beings’ working, living, and entertainment.
+I Introduction
+With the rapid development of Artificial Intelligence (AI) in the past decades, AI methods have shown their superiority over human beings and other traditional approaches across most tasks, from universal tasks in our daily lives such as image recognition [1] and text translation [2] to complicated scientific research tasks such as drug-drug interaction prediction [3] and single-cell RNA sequencing analysis [4].
+Recently, powered by Latent Diffusion Models [5] and Transformer [6], Stable Diffusion [5] and ChatGPT [7] have greatly changed human beings’ working, living, and entertainment.
 The success of AI products is usually attributed to AI models’ in-depth understanding of the accumulated data, thereby intrinsically uncovering data patterns and learning data-task correlations.
 The performances of AI models will be affected by the quantity and quality of training data, e.g.
-models trained with limited data suffer from over-fitting with significantly degenerative performance on testing datasets, and models trained with unbalanced samples will lead to poor generalization ability.
-In most situations, researchers have to overcome difficulties caused by limited data sample amounts and imbalanced distribution of data samples.
-An intuitive solution to these problems is to acquire more data, but data are scarce or difficult to collect in many cases, and labeling the data is yet another labor-intensive work.
-To solve these problems, data augmentation has been extensively applied and proven to be effective and efficient (Zha et al., 2023).
+models trained with limited data suffer from over-fitting with significantly degenerative performance on testing datasets, and models trained with imbalanced samples will lead to poor generalization ability.
+In most situations, researchers have to overcome difficulties caused by data scarcity and distribution imbalance.
+An intuitive solution to these problems is to acquire more data, but data are limited or difficult to collect in many cases, and labeling the data is yet another labor-intensive work.
+To solve these problems, data augmentation has been extensively applied and proven to be effective and efficient [8].
 The core idea of data augmentation is to artificially enlarge the training dataset by creating modified copies of existing data.
@@
 As technology continues to evolve, a variety of techniques are applied to augment data samples.
-Some simple approaches only randomly mask part of the data (DeVries and Taylor, 2017b; Zhong et al., 2020; Wei and Zou, 2019), while more sophisticated augmentation schemes involve generative adversarial networks (Bowles et al., 2018) or reinforcement learning agents (Cubuk et al., 2019).
+Some simple approaches only randomly mask part of the data [9, 10, 11], while more sophisticated augmentation schemes involve generative adversarial networks [12] or reinforcement learning agents [13].
 An ideal survey on data augmentation should be modality-independent, because only in this way can it focus on the intrinsic mechanisms of data augmentation, regardless of the data type, and thus provide insight into the nature of data augmentation.
-Existing literature surveys summarize data augmentation methods from different perspectives (Mikołajczyk and Grochowski, 2018; Shorten and Khoshgoftaar, 2019; Khalifa et al., 2022; Mumuni and Mumuni, 2022; Feng et al., 2021; Shorten et al., 2021; Bayer et al., 2022; Ding et al., 2022; Zhao et al., 2022; Zhou et al., 2022; Borisov et al., 2022; Sauber-Cole and Khoshgoftaar, 2022; Fonseca and Bacao, 2023; Wen et al., 2021; Iglesias et al., 2023; Iwana and Uchida, 2021).
+Existing literature surveys summarize data augmentation methods from different perspectives [14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29].
 As summarized in Table.
-1, each of them concentrates on a certain data modality.
+I, each of them concentrates on a certain data modality.
 Most of them categorize data augmentation methods with modality-specific (e.g.
-image modality and text modality), or operation-centric (e.g.
+image modality and text modality) or operation-centric (e.g.
 feature scaling operation, data perturbation operation) taxonomies.
@@
 However, they fail to cover data augmentation methods across data types, uncover their common patterns despite different modalities, and thus limit readers’ understanding of the essence of data augmentation.
-To fill this gap, we investigate data augmentation methods comprehensively, across five most popular data modalities.
+To fill this gap, we investigate data augmentation methods comprehensively, across the five most popular data modalities.
 We analyze how data augmentation methods utilize different sample numbers and how they leverage different components of the data information.
-Then, we propose a unified taxonomy that focuses on the way data augmentation derives new data samples.
+Then, we propose a unified taxonomy that focuses on the way data augmentation leverages information from samples.
 Finally, we investigate up-to-date literature and categorize them by our proposed taxonomy.
-The main contributions of this survey can be summarized as follows:
+The main contributions of this survey are as follows:
   • We propose a novel modality-independent taxonomy from a data-centric perspective that accommodates data augmentation techniques for all modalities consistently and inductively;
   • To the best of our knowledge, this is the first survey that covers data augmentation techniques across five data modalities, i.e. image, text, graph, tabular, and time-series modalities;
-  • We investigate how information is consistently contained in each of these data modalities and can be utilized for data augmentation.
+  • We investigate how information is consistently contained in each of these data modalities and can be utilized for data augmentation;
   • We include and categorize up-to-date literature in data augmentation.
-Table 1. A comparison of data modalities and taxonomies discussed by existing surveys and ours.
-Data Modality Taxonomies
-Image Text Graph Tabular Time-series Modality Technical Procedural Sample amount Information type
-IIPhDW 2018 (Mikołajczyk and Grochowski, 2018)
-J Big Data 2019 (Shorten and Khoshgoftaar, 2019)
-Artif Intell Rev 2022 (Khalifa et al., 2022)
-Array 2022 (Mumuni and Mumuni, 2022)
-IJCNLP 2021 (Feng et al., 2021)
-J Big Data 2021 (Shorten et al., 2021)
-Comput Surv 2022 (Bayer et al., 2022)
-KDD 2022 (Ding et al., 2022)
-arXiv 2022 (Zhao et al., 2022)
-arXiv 2022 (Zhou et al., 2022)
-TNNLS 2022 (Borisov et al., 2022)
-J Big Data 2022 (Sauber-Cole and Khoshgoftaar, 2022)
-J Big Data 2023 (Fonseca and Bacao, 2023)
-IJCAI 2021 (Wen et al., 2021)
-Neural Comput Appl 2023 (Iglesias et al., 2023)
-Plos one 2021 (Iwana and Uchida, 2021)
+TABLE I: A comparison of data modalities and taxonomies discussed by existing surveys and ours.
+Data Coverage Taxonomies
+ Image Text Graph Tabular Time-series  Modality-specific Modality-independent Operation-centric Data-centric
+IIPhDW 2018 [14]
+J Big Data 2019 [15]
+IJCNLP 2021 [18]
+J Big Data 2021 [19]
+IJCAI 2021 [27]
+Plos one 2021 [29]
+Artif Intell Rev 2022 [16]
+Array 2022 [17]
+Comput Surv 2022 [20]
+KDD 2022 [21]
+arXiv 2022 [22]
+arXiv 2022 [23]
+TNNLS 2022 [24]
+J Big Data 2022 [25]
+J Big Data 2023 [26]
+Neural Comput Appl 2023 [28]
+arXiv 2024 [30]
+arXiv 2024 [31]
+arXiv 2025 [32]
 Ours
-2. Background
-2.1. Early Evolution
+II Background
+II-A Early Evolution
 Despite the fact that data augmentation only became popular in no more than a decade, some of its embryonic forms were proposed quite early.
 The concept of data augmentation was conceived at the very beginning of Deep Learning.
-For example, the use of random distortion can be found in LeNet (LeCun et al., 1998), where ninefold the distorted images are added to the dataset to verify that increasing the size of the training set can effectively reduce test error.
-Since then, data augmentation has gradually been recognized as a best practice for training Convolutional Neural Networks (CNNs) (Simard et al., 2003; Cireşan et al., 2011; Ciregan et al., 2012).
-AlexNet (Krizhevsky et al., 2012) explicitly employs several data augmentation techniques to reduce overfitting.
+For example, the use of random distortion can be found in LeNet [33], where ninefold the distorted images are added to the dataset to verify that increasing the size of the training set can effectively reduce test error.
+Since then, data augmentation has gradually been recognized as a best practice for training Convolutional Neural Networks (CNNs) [34, 35].
+AlexNet [36] explicitly employs several data augmentation techniques to reduce overfitting.
 It augments the datasets by extracting patches from images and altering the color intensity.
 It also applies dropout, which randomly sets the output of some neurons to zero to prevent their co-adaption.
-Although dropout is often regarded as an approach to regularization, it has a similar idea to data augmentation and indeed inspires a set of augmentation methods (DeVries and Taylor, 2017b).
+Although dropout is often regarded as an approach to regularization, it has a similar idea to data augmentation and indeed inspires a set of augmentation methods [9].
 These applications of data augmentation mainly focus on increasing the size of the training dataset and introducing diversity to it.
-SMOTE (Chawla et al., 2002) is another early and representative work on data augmentation but from a different perspective.
+SMOTE [37] is another early and representative work on data augmentation but from a different perspective.
 Focusing on addressing the class imbalance problem, it suggests that oversampling the minority class can achieve better classification performance when categories are not equally represented in the dataset.
-2.2. Existing Surveys
+II-B Existing Surveys
 Given the rapid development of CNNs and their wide application in image-related tasks, most early data augmentation methods are proposed for image data.
 Also, it is natural to augment image data because many data augmentation methods suit the traits of CNNs very well, such as translation invariance.
-Published in 2018, (Mikołajczyk and Grochowski, 2018) is one of the first surveys on data augmentation.
-It gives a brief discussion of the effect of traditional and neural-based augmentation methods.
-In 2019, the use of data augmentation on image data is thoroughly studied in (Shorten and Khoshgoftaar, 2019), which is up to now the most extensive work on this track.
+Published in 2018, [14] is one of the first surveys on data augmentation.
+It briefly discusses the effects of traditional and neural-based augmentation methods.
+In 2019, data augmentation usage on image data is thoroughly studied in [15].
 It categorizes image data augmentation methods into basic image manipulation and deep learning approaches, with a wide coverage.
-It also introduces how meta-learning is used to wisely choose and combine various basic operations for data augmentation.
-(Khalifa et al., 2022) is a recent survey on image data augmentation published in 2022.
+It also introduces how meta-learning selects and combines basic data augmentation operations.
+[16] is a recent survey on image data augmentation published in 2022.
 It formalizes commonly used operations for image data augmentation with equations and illustrates their effects with examples.
-In the same year, (Mumuni and Mumuni, 2022) also elaborates on image data augmentation with plenty of examples.
-It adopts taxonomies similar to those of (Shorten and Khoshgoftaar, 2019) but includes more modern approaches and evaluates the improvements in model performance when these augmentation methods are used.
-Data augmentation on text data is not thoroughly researched as early as image data, possibly due to the discrete and correlated nature of text components.
+In the same year, [17] also elaborates on image data augmentation with plenty of examples.
+It adopts taxonomies similar to [15], but includes more modern approaches and evaluates the improvements in model performance.
+Data augmentation on text data has not been thoroughly researched as early as image data, possibly due to the discrete and correlated nature of text components.
 Still, several augmentation methods are proposed for or transplanted to text data.
-(Feng et al., 2021) is among the first surveys on text data augmentation.
+[18] is among the first surveys on text data augmentation.
 It adopts a simple taxonomy but gives more emphasis on the applications.
-Built on previous work on image data augmentation, (Shorten et al., 2021) continues to investigate how data augmentation is used to enrich text datasets, and methods are classified as symbolic augmentation and neural augmentation.
-(Bayer et al., 2022) categorizes text data augmentation methods by identifying whether they are used on the data space or feature space.
+Built on previous work on image data augmentation, [19] continues to investigate how data augmentation is used to enrich text datasets, and methods are classified as symbolic augmentation and neural augmentation.
+[20] categorizes text data augmentation methods by identifying whether they are used on the data space or feature space.
 It also presents the performance improvements of different methods.
+[31] focuses on the application of data augmentation for training Large Language Models.
+[32] discusses how Multimodal Large Language Models can be used for augmenting text, image, and speech data.
 Following the success of data augmentation methods for image and text data, there is a growing trend in augmenting graph data, too.
@@
 This also leads to more possible ways to categorize graph data augmentation methods.
-For example, (Ding et al., 2022) views graph data augmentation from the employed techniques and application scenarios, while (Zhao et al., 2022) discusses graph data augmentation in terms of its data modality, task level, and whether the augmentation operation is rule-based or learned.
-(Zhou et al., 2022) tries to elaborate on all feasible taxonomies for graph data augmentation and uses detailed schematic illustrations to explain the overview framework as well as the idea of some typical methods.
+For example, [21] views graph data augmentation from the employed techniques and application scenarios, while [22] discusses graph data augmentation in terms of its data modality, task level, and whether the augmentation operation is rule-based or learned.
+[23] tries to elaborate on all feasible taxonomies for graph data augmentation and uses detailed schematic illustrations to explain the overview framework as well as the idea of some typical methods.
 Tabular data is common in reality, but there is not as much room for data augmentation as other data types because tabular data lacks some features for augmentation to utilize.
-Some surveys only discuss augmentation techniques in the context of deep learning on tabular data (Borisov et al., 2022), and others mainly focus on how to generate tabular data (Sauber-Cole and Khoshgoftaar, 2022; Fonseca and Bacao, 2023).
+Some early surveys only discuss augmentation techniques in the context of deep learning on tabular data [24], and others mainly focus on how to generate tabular data [25, 26].
+Recently, [30] presented a dedicated and thorough review on tabular data augmentation, focusing on how generative models and retrieval techniques contribute to tabular data augmentation and discussing their future trends.
 With its growing application in scenarios such as speech recognition and IoT, time-series data is attracting more attention.
-(Wen et al., 2021) discusses basic and advanced augmentation methods.
-(Iglesias et al., 2023) focuses on generative approaches.
-(Iwana and Uchida, 2021) gives more discussion on mixing time-series data.
+[27] discusses basic and advanced augmentation methods.
+[28] focuses on generative approaches.
+[29] gives more discussion on mixing time-series data.
 Till now, data augmentation techniques for different data types have been exhaustively discussed; however, there has not been a comprehensive survey that summarizes data augmentation for all types of data.
@@
 Summarizing these similarities can help reveal the common pattern of data augmentation.
-A comprehensive survey does mean simply combining existing surveys because the proposed taxonomies in these surveys view data augmentation from different angles and are incompatible with each other.
-Some taxonomies contrast basic operations to advanced approaches; some categorize methods by their application scenarios.
+A comprehensive survey does not mean simply combining existing surveys because the proposed taxonomies in these surveys view data augmentation from different angles and are incompatible with each other.
+Some taxonomies contrast basic operations to advanced approaches; some categorize methods with modality-specific terms.
 Besides, these taxonomies provide limited insight into what happens to the information.
-Indeed, huge gaps lie between operations for data modalities due to their distinctive nature; these operation-based or application-based taxonomies reflect the development of data augmentation techniques.
+Indeed, huge gaps lie between operations for data modalities due to their distinctive nature; these operation-based or modality-specific taxonomies reflect the development of data augmentation techniques.
 However, from such taxonomies, we cannot grasp the essence of data augmentation and lack perception of the big picture of data augmentation.
-Hence there is a need for a comprehensive survey that covers data augmentation techniques for all data modalities with an all-embracing and information-centric taxonomy.
-3. Taxonomy
+Hence there is a need for a comprehensive survey that covers data augmentation techniques for all data modalities with an all-embracing and data-centric taxonomy.
+III Taxonomy
 Before presenting the taxonomy, we formalize data augmentation as follows.
 Given a labeled dataset $\mathcal{D}_{L}=\{\mathbf{X},\mathbf{y}\}$, where $\mathbf{X}$ stands for the data, and $\mathbf{y}$ stands for the labels, data augmentation can be represented by a certain function $f_{\theta}$, such that the augmented dataset $\mathcal{\tilde{D}}_{L}=\{\mathbf{\tilde{X}},\mathbf{\tilde{y}}\}$ is derived by:
-(1) $f_{\theta}:\mathcal{D}_{L}=\{\mathbf{X},\mathbf{y}\}\rightarrow\mathcal{\tilde{D}}_{L}=\{\mathbf{\tilde{X}},\mathbf{\tilde{y}}\}$
-This representation also applies to an unlabeled dataset $\mathcal{D}_{U}=\{\mathbf{X}\}$ or a partially labeled dataset $\mathcal{D}=\{\mathbf{X}_{L}\cup\mathbf{X}_{U},\mathbf{y}_{L}\}$, but for simplicity, in this section, we only use a labeled dataset to explain the taxonomy.
+$f_{\theta}:\mathcal{D}_{L}=\{\mathbf{X},\mathbf{y}\}\rightarrow\mathcal{\tilde{D}}_{L}=\{\mathbf{\tilde{X}},\mathbf{\tilde{y}}\}$ (1)
+This representation also applies to an unlabeled dataset $\mathcal{D}_{U}=\{\mathbf{X}\}$ or a partially labeled dataset $\mathcal{D}_{P}=\{\mathbf{X}_{L}\cup\mathbf{X}_{U},\mathbf{y}_{L}\}$, but for simplicity, we only use a labeled dataset to explain the taxonomy.
 Some data augmentation approaches are used in the input data space, while others are used in the latent feature space.
 This section uses $\mathbf{X}$ as a space-agnostic representation.
-In this comprehensive survey, we try to propose a two-tier taxonomy from an information-centric perspective that can be applied to all data modalities.
+In this survey, we try to propose a two-tier taxonomy from a data-centric perspective that can be applied to all data modalities.
 The main consideration of our taxonomy is whence the information in the augmented data comes.
@@
 The answers to these questions constitute the taxonomy hierarchy which we present below.
-By answering RQ1, we can divide data augmentation approaches into Individual, Multiple, and Populational Augmentation, which build the first tier of the taxonomy in our survey.
-Individual Augmentation.
+By answering RQ1, we can divide data augmentation approaches into Single-instance Level Augmentation, Multi-instance Level Augmentation, and Dataset Level Augmentation, which build the first tier of the taxonomy in our survey.
+Figure 1: A taxonomy for data augmentation (left) and how information in each data modality is utilized (right).
+Single-instance Level Augmentation.
 This type of data augmentation method transforms one data sample without referring to others to create new data.
@@
 The information of new data comes from exactly one sample in the original dataset, and each new sample often lies around its original sample in the feature space.
-We formalize Individual Augmentation as follows, where $\epsilon(\mathbf{x}_{i})$ stands for a certain perturbation that is related to the original data $\mathbf{x}_{i}$:
-(2) $\mathbf{\tilde{x}}=\mathbf{x}_{i}+\epsilon(\mathbf{x}_{i}),\qquad\tilde{y}=y_{i}$
+We formalize single-instance level augmentation as follows, where $\epsilon(\mathbf{x}_{i})$ stands for a certain perturbation related to the original data $\mathbf{x}_{i}$:
+$\mathbf{\tilde{x}}=\mathbf{x}_{i}+\epsilon(\mathbf{x}_{i}),\qquad\tilde{y}=y_{i}$ (2)
 We further answer RQ2 to build the second tier of the taxonomy.
@@
 For most data, there is also some structural information that connects these elements in a certain way, such as the positional relationship of pixels or the syntax of a sentence.
-The key to designing data augmentation methods is to think of how to properly leverage the two types of information.
-So, we can categorize Individual Augmentation according to which part of the information is used for data augmentation:
+The key to designing data augmentation methods is to think of how to wisely leverage the two types of information.
+So, we can categorize single-instance level augmentation according to which part of the information is used for data augmentation:
   • Value-based Transformation, which perturbs the value that an element carries.
   • Structure-based Transformation, which perturbs elements’ structural relationships.
-Multiple Augmentation.
+Multi-instance Level Augmentation.
 This type of data augmentation method makes use of multiple data to acquire new ones.
@@
 New samples derived from these methods often lie between the source data in the feature space.
-We formalize Multiple Augmentation as follows, where $i$ and $j$ are two data samples from the original dataset, and $\lambda$ is a scalar factor that often lies between $[0,1]$:
-(3) $\mathbf{\tilde{x}}=\lambda\cdot\mathbf{x}_{i}+(1-\lambda)\cdot\mathbf{x}_{j},\qquad\tilde{y}=\lambda\cdot y_{i}+(1-\lambda)\cdot y_{j}$
-Now we answer RQ2 for Multiple Augmentation.
-Similarly to Individual Augmentation, this type of data augmentation is also conducted utilizing either value of structure information, and the detailed classification is as follows:
+We formalize multi-instance level augmentation as follows, where $i$ and $j$ are two data samples from the original dataset, and $\lambda$ is a scalar factor that often lies between $[0,1]$:
+$\mathbf{\tilde{x}}=\lambda\cdot\mathbf{x}_{i}+(1-\lambda)\cdot\mathbf{x}_{j},\qquad\tilde{y}=\lambda\cdot y_{i}+(1-\lambda)\cdot y_{j}$ (3)
+Now we answer RQ2 for multi-instance level augmentation.
+Similarly to single-instance level augmentation, this type of data augmentation is also conducted utilizing either value or structure information, and the detailed classification is:
   • Value-based Mixture, which arithmetically mixes the values of multiple data.
   • Structure-based Combination, which pieces multiple data (or parts of them) together.
-Populational Augmentation.
-This type of data augmentation method does not explicitly make use of one or multiple data but rather uses the population of the data.
-It comprehensively learns the features and abstractions of the entire dataset and samples completely new data from a learned distribution of the data.
+Dataset Level Augmentation.
+This type of data augmentation method does not explicitly make use of one or multiple data but rather uses the entire dataset.
+It comprehensively learns the features and abstractions of the dataset, and samples completely new data from a learned distribution of the dataset.
 The generated data should also fall into this distribution.
-We formalize Populational Augmentation as follows, where $P(\cdot)$ stands for a probability distribution learned from the original dataset $\mathcal{D}_{L}=\{\mathbf{X},\mathbf{y}\}$:
-(4) $\vspace{-1.5mm}\mathbf{\tilde{x}}\sim P(\mathbf{X}),\qquad\tilde{y}\sim P(\mathbf{y}|\mathbf{\tilde{x}})$
-The answer to RQ2 for Populational Augmentation is different from the two categories mentioned above.
-When generating new data based on population distributions, most methods simultaneously consider value and structural information.
-Some methods also introduce external knowledge when learning from the data population.
+We formalize dataset level augmentation as follows, where $P(\cdot)$ stands for a probability distribution learned from the original dataset $\mathcal{D}_{L}=\{\mathbf{X},\mathbf{y}\}$:
+$\vspace{-1.5mm}\mathbf{\tilde{x}}\sim P(\mathbf{X}),\qquad\tilde{y}\sim P(\mathbf{y}|\mathbf{\tilde{x}})$ (4)
+The answer to RQ2 for dataset level augmentation is different from the two categories mentioned above.
+When generating new data based on dataset distributions, most methods simultaneously consider value and structural information.
+Some methods also introduce external knowledge when learning from the source dataset.
 So, we have the following classification:
   • Vanilla Generation, where the generation process only resorts to the existing dataset.
-  • Exogenous Generation, where the generation process refers to external resources such as other datasets or expert knowledge.
-We have presented a general taxonomy for data augmentation, whereas the category is at a high level, and each data modality has its specific data augmentation techniques depending on the nature of that modality.
-In addition, the equations that we use as a general formalization are only abstract and conceptual.
-They do not literally represent all specific data augmentation techniques in these categories.
-From paragraph 4 to paragraph 8, we will elaborate on how this taxonomy applies to each data modality and review some typical data augmentation methods.
-Figure 1. A taxonomy for data augmentation (left) and how information in each data modality is utilized (right).
-4. Data Augmentation for Image Data
-The elemental components of an image are pixels.
+  • Exogenous Generation, which uses external resources like other datasets or expert knowledge.
+We have presented a general taxonomy for data augmentation at a high level with abstract and conceptual formalization.
+We organize this survey in accordance with this taxonomy, from Section IV to Section VI.
+We elaborate on how the proposed taxonomy applies to each data modality and review some typical data augmentation methods.
+IV Single-instance Level Augmentation
+Single-instance level augmentation methods perturb one data sample’s value or structure information to obtain the augmented data.
+Different data modalities represent value information with different elemental components and organize these components with different types of structure.
+For image data, the elemental components are pixels.
 A pixel carries unit information that uses a numerical value to represent the color.
 Each pixel also has spatial relationships with other pixels, so, images are well-structured.
-Data augmentation for image data revolves around the manipulation of the colors and spatial relationships of image pixels.
-4.1. Individual Image Augmentation
-4.1.1. Value-based Image Transformation
-Two kinds of methods are used to perturb the value a pixel carries, and thus change the color it displays.
-One is called Pixel Erasing, which completely masks out the color information in an area of the data; another is called Photometric Transformation, which only imposes a soft adjustment to change the appearance of the image, but the essential information remains visible.
-Pixel Erasing.
-Pixel Erasing is one of the simplest yet effective types of method in data augmentation.
-The idea is to mask some information from the original input space or the feature space.
-Typical work like Cutout (DeVries and Taylor, 2017b) and random erasing (Zhong et al., 2020) first selects a rectangle region and erases all pixels in that region.
-The erased region is then filled with a certain color or Gaussian noises.
-Table 2. Region selection strategy for Pixel Erasing methods
-Name Ref Augmentation Strategy
-Random Location-based
-Cutout (DeVries and Taylor, 2017b)
-Image-aware Random Erasing (Zhong et al., 2020)
-Object-aware Random Erasing (Zhong et al., 2020)
-Region-aware Random Erasing (Yang et al., 2019)
-Hide-and-seek (Kumar Singh and Jae Lee, 2017)
-Grid-Mask (Chen et al., 2020a)
-Random Erasing in the frequency domain (Hwang et al., 2023)
-If erasing is performed by randomly selecting an area in the image, it is called Image-aware Random Erasing (IRE) (Zhong et al., 2020).
-In some tasks, the location of the target in the image is known, so random erasing can be restricted in the object’s bounding box, known as Object-aware random erasing (ORE) (Zhong et al., 2020).
-Further argued in (Yang et al., 2019), the erased area within the bounding box should be limited to a manageable size and a part of the less informative background should also be erased.
-(Yang et al., 2019) proposed Region-aware Random Erasing (RRE), which first introduces ORE with upper bounds for width and height in the bounding boxes, and then randomly erases the background.
-Random Erasing typically adds one or two big rectangular masks to the image, and some other methods deploy more flexible implementations.
-Hide-And-Seek (Kumar Singh and Jae Lee, 2017) proposes to randomly remove multiple parts of the image so that the model has to learn from multiple parts of the image other than the most discriminating part.
-Similarly, the Grid-Mask method (Chen et al., 2020a) generates a set of squared masks under the grid layout.
-It can prevent excessive deletion and reservation and achieve a reasonable balance between deleted and reserved regional information.
-In most research, random erasing is used in the spatial domain, while REF (Hwang et al., 2023) extends the use of random erasing to the frequency domain.
-It uses the discrete Fourier transform to process the image, applies random erasing on the image’s frequency domain, and uses the inverse discrete Fourier transform to recover the image to the spatial domain.
-Photometric Transformation. In contrast to Pixel Erasing, which put masks on an area of the image, Photometric Transformation changes the overall color and style of the image.
-Some methods alter the value of each pixel without considering other pixels.
-The simplest way is to adjust the brightness of the image.
-Color inversion replaces the pixels with their “opposite” color.
-Color casting applies different operations in different channels.
-For example, red color casting increases the image’s red channel while keeping the green channel and blue channel untouched.
-Noise injection adds Gaussian or other types of noises to the images.
-It helps the model learn from various distorted features and become robust.
-In addition, this method can produce a large amount of augmented data with low cost and complexity.
-Some other methods decide how to adjust the value of pixels according to other pixels.
-Histogram equalization (Cheng and Shi, 2004; Abdullah-Al-Wadud et al., 2007), for example, enhances the image by enhancing its contrast, and white balancing simulates alternative light cast on the object.
-Kernel filters are matrices that slide across and are convoluted with original the image to generate new images.
-Depending on the value of the matrix, a kernel filter can be used to sharpen or blur the image, or make other changes.
-A sharpening filter helps extract the features, and a blurring filter can improve the robustness of the model.
-(Reichenbach et al., 1990), for instance, uses kernel filters to improve edge detection performance.
-Greyscaling transforms a colored picture into black-and-white by combining channels for different colors.
-This change forces the model to focus on the shapes or edges of target objects instead of their colors.
-Apart from this, greyscaling would be especially useful in compressing data and improving the model’s efficiency.
-A counter-operation of grey-scaling is Grayscale-to-Color Conversion (GCC) (Gupta et al., 2020).
-Since most pre-trained models for image-related tasks are trained on colored images, single-channeled greyscale images should be colored to fit the models.
-4.1.2. Structure-based Image Transformation
-This type of Individual Image Augmentation manipulates the spatial relationship between pixels.
-Image Cropping breaks the original image into pieces, and Geometric Transformation perturbs the structure within the original data.
-Image Cropping.
-Image Cropping refers to selecting an area from the original image and removing other parts of the image.
-Some cropping techniques are discussed in the context of reducing the cost of image storage and transmission, improving image quality, and reducing computation complexity and cost when training artificial intelligence models.
-However, it can also be used as a data augmentation technique, in that image cropping can produce multiple samples by selecting different regions from the original image, and prevent the models from being affected by irrelevant background information in the images.
-Primitive cropping methods make use of manual selection or random cropping, while some methods employ more advanced techniques to perform this task, including attention-based approaches and aesthetics-based approaches (Chen et al., 2016).
-Attention-based approaches use certain attention scores to evaluate a subimage.
-Regions that contain the main topic or other types of information often have high attention scores.
-(Ma and Guo, 2004) proposes a general solution, which uses the entropy, area ratio, and center distance of the image to evaluate the performance of the image segmentation and accordingly adjust the cropping parameters.
-(Ciocca et al., 2007) uses an adaptive cropping method that applies different processes depending classification result of the image.
-These studies discuss the way to evaluate the importance of each part of the image, while they overlook how to search for the optimal cropping rectangle.
-Formulations of this problem and algorithms with low computational complexity are given by (Chen et al., 2016).
-Apart from these automatic methods, there are also some interactive image cropping methods.
-(Santella et al., 2006), a gaze-based image cropping approach, uses eye tracking techniques to capture human fixations and thus infer the location of important contents on the image.
-Unlike fully hand-crafted image cropping, the implicit interaction applied in eye tracking is less burdensome and less time-consuming.
-Aesthetics-based approaches (Nishiyama et al., 2009; Cheng et al., 2010; Zhang et al., 2012) are similar to quality evaluation.
-For example, (Nishiyama et al., 2009) detects several subject regions and one background region, extracts features from the regions, estimates and combines posterior probabilities, and determines the quality score according to these probabilities.
-A2-RL (Li et al., 2018b) integrates image cropping into a sequential decision-making process and an aesthetics-aware reinforcement learning model to solve the best cropping problem.
-Beyond attention-based and aesthetic-based cropping, (Yan et al., 2013) proposed a change-based approach.
-It takes into account the content that is removed and extracts two sets of specially designed features to model what is changed after cropping.
-Geometric Transformation.
-As one of the most classical image data augmentation methods, Geometric Transformation changes the spatial relationship of pixels within an image while keeping the value of each pixel untouched to create more data without extra sampling or labeling.
-It increases the number of training samples, extends the input space, and better represents the real world by deriving more possible views from the original images.
-For example, flipping the image of a cat can be understood as taking another picture of the cat from a different orientation.
-Affine transformations like flipping, shearing, rotation, and translation manipulate the image while preserving the colinearity and ratios of distances.
-Image flipping reflects the image around its vertical axis, horizontal axis, or both axes, rotation turns the image by an angle about a fixed point.
-Image shearing fixs all pixels along a given line, and other pixels are shifted parallel to that line by a distance proportional to their perpendicular distance from the line.
-Translation means moving an image in a certain direction.
-It can introduce more variance in the position of target objects in the image.
-In non-affine transformation, the spatial relationship between pixels is further changed compared with affine transformations because the colinearity and ratio of distances are broken.
-One example of non-affine transformation is perspective transformation (Wang et al., 2019), which mimics images taken from other views of the object.
-Image stretching (Wang et al., 2008) is a kind of content-aware image resizing.
-It preserves or even enlarges the prominent contents in the image and compresses the background or unimportant foreground information.
-For example, if there is a person in the middle of an image, the middle part of the image is made wider, and the left and right parts are made narrower.
-Distortion (Simard et al., 2003; Jaderberg et al., 2015) also stretches the pixels but often focuses on a certain region.
-It aims at fixing distortion in the original images and is especially useful in digit classification tasks where the handwriting is poor or when the images are distorted by a camera lens (e.g.
-wide-angle lens) when collecting them.
-PatchShuffle (Kang et al., 2017) divides an image into block matrices and shuffles the pixels within each block under a certain probability.
-This transformation introduces more variance in the image and extends the input space or feature space of the training data.
-4.1.3. Image Value-structure Transformation
-Automatic Data Augmentation methods search for the optimized augmentation policy.
-The policy is derived from a search space that includes both value-based transformations and structure-based transformations.
-When implementing the augmentation, both pixels’ coloring and the positions are changed.
-Automatic Data Augmentation.
-AutoAugment (Cubuk et al., 2019) is among the most significant achievements in Automatic Data Augmentation.
-It represents data augmentation as a decision-making process and uses reinforcement learning to search for the best operations.
-In AutoAugment, an operation can be any basic value-based transformation like shearing, or structure-based transformation like histogram equalization.
-A sub-policy consists of two operations along with their probabilities and magnitudes with which the operations are applied.
-A policy consists of many sub-policies.
-All sub-policies and their parameters form the search space (action space), and a controller RNN (agent) picks the sub-polices from the search space and gets a policy (action).
-Images are processed under the policy, which produces several augmented images.
-The augmented images are fed to downstream networks (environment) and then AutoAugment uses the network’s performance (reward) to update the controller.
-As a result, the controller is trained to pick the policies and parameters that can yield the best network performance.
-This optimal policy will be used on all images and is also transferable so that it can be used on other datasets, too.
-Fast AutoAugment (Lim et al., 2019) proposes a more efficient searching algorithm for augmentation policies, which works significantly faster than AutoAugment and thus allows a larger number of sub-policies, which should benefit the generalization performance.
-Similarly, RandAugment (Cubuk et al., 2020) removes the need for a separate search phase on a proxy task, and makes the training process less complicated and less expensive.
-Population Based Augmentation (PBA) (Ho et al., 2019) learns a augmentation schedule rather than a fixed policy as in AutoAugment to improve efficiency.
-These methods combine the benefits of both value-based transformations and structure-based transformations and further discover the potential of data augmentation with individual image data.
-4.2. Multiple Image Augmentation
-4.2.1. Value-based Image Mixture
-This type of Multiple Image Augmentation computes an arithmetical mixture (e.g. interpolation) of the color from different images.
-Image MixUp.
-Image mixup is a category of data augmentation methods that combine two images pixel-wise.
-Each pixel of the augmented image is produced by interpolating the pixels from the original images.
-Mixup (Zhang et al., 2017) mixes two random images on the input layer.
-Pixels from two original images each contribute a portion to the new image:
-(5) $\mathbf{\tilde{x}}=\lambda\mathbf{x}_{i}+(1-\lambda)\mathbf{x}_{j},\qquad\mathbf{\tilde{y}}=\lambda\mathbf{y}_{i}+(1-\lambda)\mathbf{y}_{j}$
-where $(\mathbf{x}_{i},\mathbf{y}_{i})$ and $(\mathbf{x}_{j},\mathbf{y}_{j})$ are two labeled images from the dataset, and $\lambda\in[0,1]$.
-As a variant, manifold mixup (Verma et al., 2019) is operated on the hidden layers of the neural network.
-Most mixup methods combine two different images, while AugMix (Hendrycks et al., 2019) applies three augmentation operations to one source image, and combines the three augmented images to generate a new sample.
-In the above-mentioned mixup approaches, the mixing factor is even across the entire image, so they are region-consistent.
-In region-variant mixup approaches, the mixing operation varies in strength across regions.
-For example, SmoothMix (Lee et al., 2020) applies a soft-edged mask when mixing.
-Two more sophisticated methods are Co-mixup (Kim et al., 2021) and Puzzle mix (Kim et al., 2020), which both use saliency to decide the mixing strength of different regions, potentially leading to better mixing performances.
-Most mixup methods create new samples with soft labels, as in Eq.5, while there are some exceptions.
-Sample Paring (Inoue, 2018), for example, uses the label of one of the original images, and that of the other image is left unused.
-ReMix (Chou et al., 2020) is designed to address the class imbalance problem, and the label of the augmented data is assigned with the label from the minority class.
-Neural Blending. In mixup, we explicitly and arithmetically blend different images. However, this process can benefit from the use of neural networks.
-AutoMix (Liu et al., 2022a), for example, integrates a Mix Block in the training process.
-It builds a bridge between the selection of mixup policy and the optimization of the model.
-In AutoMix, the augmentation module is optimized along with the classification feature map, and the mixup functions and mixing ratio will updated according to the classification loss.
-Another image blending technique is proposed as Smart Augmentation (Lemley et al., 2017).
-At the input end of Smart Augmentation, two random samples are fed into the network through two channels, and a CNN sub-network yields a one-channel output with the same shape as the input, which means one new image is generated based on the information of the two input images.
-The augmented image is then used for downstream tasks, and the augmentation sub-network adjusts to the task performance accordingly.
-4.2.2. Structure-based Image Combination
-Methods that combine images based on the position of pixels and image substructure are called Image Patching.
-It refers to putting together small fragments of many images or putting a part of an image over a second image.
-Image Patching.
-RICAP (Takahashi et al., 2019) randomly crops four images, and patches them to construct a new image.
-CutMix (Yun et al., 2019) makes use of two images.
-It is developed based on Cutout (DeVries and Taylor, 2017b), but the removed region is replaced by a patch from another image.
-Attentive CutMix (Walawalkar et al., 2020) further enhances the CutMix method by adding a feature extraction module that selects an important or representative region, and pastes this attentive patch onto the other image.
-Image patching is not limited to rectangular shapes.
-CowMask (French et al., 2020), for example, introduces patching with irregular shape rather than rectangular.
-Figure 2.
-A conceptual demonstration of (a) Value-based Individual Image Transformation, (b) Structure-based Individual Image Transformation, (c) Value-based Multiple Image Mixture, and (d) Structure-based Multiple Image Combination.
-4.3. Populational Image Augmentation
-4.3.1. Vanilla Image Generation
-Generative Adversarial Networks.
-The rise of Generative Adversarial Networks (GAN) (Goodfellow et al., 2014) makes it possible for humans to generate new images on a large scale using computers.
-This nature also leads to a new path of data augmentation: creating new images instead of modifying or combining old ones.
-GANs have now been widely used as a data augmentation technique (Bowles et al., 2018), especially in the medical field (Frid-Adar et al., 2018; Kaur et al., 2021), where data is to some extent scarce.
-Apart from simply enlarging the dataset, GAN can be used in other way such as removing background noise and cleaning the dataset (Kupyn et al., 2018; Sharma et al., 2019; Ashraf et al., 2021).
-It can also generate different views or poses of the object based on the original image (Zhang et al., 2022), and thus add more variation to the dataset.
-4.3.2. Exogenous Image Generation
-Neural Style Transfer introduces style information from external dataset when augmenting an entire dataset.
-Computer Graphic Modelling is to leverage expert knowledge on how objects in the existing dataset looks like and the power of computer software to generate new image samples.
-Neural Style Transfer. Neural Style Transfer (Gatys et al., 2016) learns the artistic style from a group of image data and uses this artistic style to depict the topic of images from another dataset.
-As is (Gatys et al., 2016), Neural Style Transfer can be used to mimic the drawing styles of particular artists.
-Such an augmentation process can help the model learn a more general representation of the topics.
-(Li et al., 2018a) proposes a more realistic use of neural style transfer.
-It transfers the weather or illumination from reference images to the target dataset.
-This is particularly useful because it is hard to collect sufficient images under certain weather.
-Computer Graphics Modelling.
-Computer Graphics Modelling refers to artificially creating images by computer software, and it can be used to produce datasets for model training (Gaidon et al., 2016; McCormac et al., 2017; Barbosa et al., 2018).
-These datasets have several advantages over ordinary image datasets.
-First, it requires little human efforts to collect data.
-Second, by applying different aspects, poses, and lighting condition, computer graphics modelling can generate images with high diversity.
-Third, these images are all fully labelled with accurate ground truth.
-Table 3. A summary for image data augmentation
-Tier 1: Quantity of source data Tier 2: Information type Augmentation method Typical literature
-Individual Image Augmentation Value-based Transformation Pixel Erasing Cutout (DeVries and Taylor, 2017b) (arXiv 2017), Random Erasing (Zhong et al., 2020) (AAAI 2020), Region-aware Random Erasing (Yang et al., 2019) (ICCT 2019), Hide-and-seek (Kumar Singh and Jae Lee, 2017) (ICCV 2017), Grid-Mask (Chen et al., 2020a) (arXiv 2020), REF (Hwang et al., 2023) (ICEIC 2023)
-Photometric Transformation multi-peak GHE (Cheng and Shi, 2004) (Digital Signal Processing 2024), DHE (Abdullah-Al-Wadud et al., 2007) (TCE 2007), (Reichenbach et al., 1990) (ICPR 1990), GCC (Gupta et al., 2020) (JDI 2020)
-Structure-based Transformation Image Cropping (Ma and Guo, 2004) (CCNC 2004), Self-adaptive Image Cropping (Ciocca et al., 2007) (TCE 2007), Automatic Image Cropping (Chen et al., 2016) (CVPR 2016), Gaze-based Cropping (Santella et al., 2006) (CHI 2006), Sensation-based Cropping (Nishiyama et al., 2009) (MM 2009), (Cheng et al., 2010) (MM 2010), (Zhang et al., 2012) (TIP 2012), Change-based Image Cropping (Yan et al., 2013) (CVPR 2013), A2-RL (Li et al., 2018b) (CVPR 2018)
-Geometric Transformation Perspective Transformation (Wang et al., 2019) (IEEE Access 2019), Stretching (Wang et al., 2008) (SIGGRAPH 2008), Distortion (Simard et al., 2003) (ICDAR 2003), STN (Jaderberg et al., 2015) (NeurIPS 2015), PatchShuffle (Kang et al., 2017) (arXiv 2017)
-Value-structure Transformation Automatic Augmentation AutoAugment (Cubuk et al., 2019) (CVPR 2019), Fast AutoAugment (Lim et al., 2019) (NeurIPS 2019), RandAugment (Cubuk et al., 2020) (CVPR 2020), PBA (Ho et al., 2019) (ICML 2019)
-Multiple Image Augmentation Value-based Mixture Image Mixup Mixup (Zhang et al., 2017) (arXiv 2017), Manifold Mixup (Verma et al., 2019) (ICML 2019), AugMix (Hendrycks et al., 2019) (arXiv 2019), SmoothMix (Lee et al., 2020) (CVPR 2020), Co-Mixup (Kim et al., 2021) (arXiv 2021), Puzzle Mix (Kim et al., 2020) (ICML 2020), ReMix (Chou et al., 2020) (ECCV 2020), Sample Paring (Inoue, 2018) (arXiv 2018)
-Neural Blending AutoMix (Liu et al., 2022a) (ECCV 2022), Smart Augmentation (Lemley et al., 2017) (IEEE Access 2017)
-Structure-based Combination Image Patching RICAP (Takahashi et al., 2019) (TCSVT 2019), CutMix (Yun et al., 2019) (ICCV 2019), Attentive CutMix (Walawalkar et al., 2020) (ICASSP 2020), CowMask (French et al., 2020) (arXiv 2020)
-Populational Image Augmentation Vanilla Generation Generative Adversarial Networks GAN Augmentation (Bowles et al., 2018) (arXiv 2018), (Frid-Adar et al., 2018) (Neurocomputing 2018), (Kaur et al., 2021) (ICAIA 2021), DeblurGAN (Kupyn et al., 2018) (CVPR 2018), (Sharma et al., 2019) (ACCV 2019), (Ashraf et al., 2021) (IEEE Access 2021), Re-ID (Zhang et al., 2022) (Neural Computing and Applications 2022)
-Exogenous Generation Neural Style Transfer Neural Style Transfer (Gatys et al., 2016) (CVPR 2016), Photorealistic Image Stylization (Li et al., 2018a) (ECCV 2018)
-Computer Graphics Modelling Virtual KITTI (Gaidon et al., 2016) (CVPR 2016), SceneNet RGB-D (McCormac et al., 2017) (CVPR 2017), SOMAnet (Barbosa et al., 2018) (CVIU 2018)
-5. Data Augmentation for Text Data
-The elemental components of text data are words, and sometimes the words are processed in the form of tokens for later convenience.
+Single-instance level augmentation for image data revolves around the perturbation of the image pixels’ colors and spatial relationships.
+For text data, the elemental components of text data are words (or sometimes, tokens).
+A word often expresses minimal meanings.
 Words are organized into sentences following syntactic rules.
 The sentence is a sequential representation since one word comes after another, but there is also a tree-structured grammatical dependency behind the sentence.
-Data augmentation for text data often changes the wording or reshapes the structure of a sentence.
-5.1. Individual Text Augmentation
-5.1.1. Value-based Text Transformation
-Text data augmentation of this type uses word-level manipulation. It applies changes to the choice of words or adds or removes some words.
-Invariant Replacement. Invariant Replacement refers to those augmentations that try to use alternative wording in a given sentence but convey the same meaning.
-EDA (Wei and Zou, 2019), for example, proposes a set of text augmentations, including synonym replacement.
-It randomly selects several words from a sentence, substituting them with their synonyms, i.e.
-words that mean exactly or nearly the same.
-There are several ways to implement Synonym Replacement.
-(Marivate and Sefara, 2020) uses WordNet (Miller, 1995), a lexical database to find synonyms for verbs and nouns in sentences and thus make replacements.
-Apart from WordNet, (Marivate and Sefara, 2020) also derives word embeddings by Word2Vec (Mikolov et al., 2013) and replaces a word with others who have high cosine similarity.
-Another possibility is to use a knowledge graph.
-A knowledge graph describes the relationships of entities, and by utilizing the “is equivalent” relation, we can easily find a series of synonyms for a given word.
-A special case of synonym replacement is discussed in (Liu et al., 2011).
-Non-standard tokens are frequently used in people’s daily lives, especially for text messages and social network websites, posing challenges to language processing tasks.
-(Liu et al., 2011) models the process of generating non-standard tokens from standard dictionary words by letter removal or substitution and proposes a letter transformation approach to handle non-standard tokens.
-Although used as a normalization method, this technique could also be applied in data augmentation tasks to transform non-standard tokens to standard form, or vice versa.
-Token Addition. Token Addition refers to adding extra tokens to the original text, either introducing junk tokens as noise or enriching the sentences and making them more informative.
-The random insertion, proposed in EDA (Wei and Zou, 2019), inserts a random synonym of a random word at a random position in the sentence.
-The words added in the EDA play the role of data noising and may carry either little or harmful information.
-As a variant of EDA, the technique of AEDA (Karimi et al., 2021) only inserts punctuation marks.
-It only adds limited noise and avoids the potentially negative effect of random insertion.
-INSNET (Lu et al., 2022) is an insertion-based text generation model that can sequentially insert new tokens into an incomplete sentence, and with consecutive insertion steps, the sentence is enriched and contains more information.
-LeCA (Chen et al., 2021) introduces lexical constraints to improve the performance of neural machine translation.
-In LeCA, a set of words is supposed to appear in the translation result.
-These words are called constraints whose embeddings are appended to those of the original tokens.
-The embeddings are then fed to the transformers’ encoding layer, and these constraint words are very likely to be included in the output text.
-A very similar research is (Hokamp and Liu, 2017), which also uses lexical constraints.
-The difference is that in LeCA, the constraints are given a soft suggestion, rather than being forced to appear as in (Hokamp and Liu, 2017).
-Token Deletion.
-Random deletion (Wei and Zou, 2019) randomly removes some words from a sentence.
-In (Xie et al., 2017), noising techniques are used to avoid overfitting, and the randomly selected word is replaced by a blank placeholder token “_”, rather than simply being removed.
-In (Duan et al., 2023), apart from word dropout which is similar to random deletion, it also uses word blanking to replace a word with a special “¡BLANK¿” placeholder.
-On Word Selection.
-Most methods randomly select the word to be replaced, added, and deleted.
-But some method considers the role a word plays in the sentence when deciding whether to impose perturbation.
-For example, TEXTFOOLER (Jin et al., 2020a) ranks words by their importance and replaces the important words with their synonyms.
-Syntax-aware data augmentation (Duan et al., 2023) argues that most previous text data augmentation approaches are only computation-motivated and cannot assure the good quality of the generated text.
-Instead of simply modifying words with the same probability, Syntax-aware data augmentation calculates the probability of being modified for each word based on its syntactic role in the sentence.
-Specifically, it utilizes a dependency tree to calculate the importance of each word, i.e.
-the depth in the tree.
-A word far from the root is considered less important syntactically and is more likely to be removed (by means of dropout and blanking) or replaced during the augmentation process.
-Table 4. Word selection strategy
-Name Ref Augmentation Strategy
-Random Role-based Learned
-EDA (Wei and Zou, 2019)
-- (Marivate and Sefara, 2020)
-TEXTFOOLER (Jin et al., 2020a)
-LetterTran (Liu et al., 2011)
-AEDA (Karimi et al., 2021)
-INSNET (Lu et al., 2022)
-LeCA (Chen et al., 2021)
-GBS (Hokamp and Liu, 2017)
-- (Xie et al., 2017)
-STA (Guo et al., 2022)
-Instead of using a syntax tree to assess word importance, Selective Text Augmentation (STA) (Guo et al., 2022) focuses more on the statistical and semantic features of words to describe the role of words.
-It proposes two perspectives to view words in text classification tasks, namely, statistical correlation, which means how frequently a word co-occurs with a certain category, and semantic similarity, which measures the semantic overlap of the given words and the category label.
-Using the metrics as axes, STA divides the roles of words into four quadrants: gold, venture, bonus, and trivial.
-Words with different roles are treated differently when imposing augmentation operations on the sentence.
-For example, gold words will be kept in selective deletion and synonyms of gold words are added in selective insertion.
-5.1.2. Structure-based Text Transformation
-We categorize transformation that perturbs sentence structure into two types: Sentence Cropping that splits a sentence into many pieces, and Sentence Morphing that swaps the location of words or grammatical components.
-Sentence Cropping.
-Borrowing the idea of image cropping, which selects segments from the images, Sentence Cropping (Şahin and Steedman, 2018) divides sentences into several small pieces.
-It often uses dependency trees to extract the relationships between subjects and objects of the sentence and produces multiple smaller sentences that focus on different parts of the original sentence.
-This operation breaks the syntactic solidity and semantic comprehensiveness of the entire sentence but preserves local syntactic tags and shallow semantic labels.
-Sentence Morphing. These methods manipulate sentence structure by altering the position of words or terms and thus changing the syntax of the sentence.
-Random swap (Artetxe et al., 2018; Wei and Zou, 2019) randomly chooses two words in a sentence and swaps their position.
-This perturbs the structure and adds noise to the text data across paragraphs.
-(Min et al., 2020) studies two types of syntactic augmentations: inversion and passivization.
-Inversion swaps the subject and object in a sentence, and the semantics might be different in most cases.
-Passivization converts a sentence to its passive form, and the meaning of the new sentence is typically consistent with the original one.
-Sentence rotation (Şahin and Steedman, 2018) borrows the idea of image rotation, where the image is rotated around its center.
-In sentence rotation, the root of the sentence is chosen as the center of the sentence, and other sentence fragments are connected to their center via the dependency tree.
-For some languages, such an operation is semantic-preserving, while for other languages the sentence’s meaning is subject to change.
-The DRNNs model (Xu et al., 2016) also utilizes the dependency tree and subject-object knowledge for data augmentation.
-It samples the shortest dependency path from a sentence’s dependency parse tree and then extracts two sub-paths, i.e.
-the subject-predicate path and the object-predicate path.
-Augmentation is done by changing the order of the two paths.
-Table 5. Demonstration of some elementary operations for text data augmentation
-Operation Sentence
-Original I am an art student and I paint a lot of pictures. Many people pretend that they understand modern art.
-Synonym Replacement (Wei and Zou, 2019) I am an art student and I draw a lot of pictures. Many people pretend that they understand contemporary art.
-Similarity Replacement (Marivate and Sefara, 2020) I am an art student and I paint a lot of photos. Many people pretend that they understand modern literature.
-Random Insertion (Wei and Zou, 2019) I am an art student and picture I paint a lot of pictures. Many appreciate people pretend that they understand modern art.
-Punctuation Insertion (Karimi et al., 2021) I am an, art student and: I paint a lot of pictures. Many people? pretend that they. understand modern art.
-Random Swap (Wei and Zou, 2019) I am an of student and I paint a lot art pictures. Many people they that pretend understand modern art.
-Random Deletion (Wei and Zou, 2019) I am an art student and I paint a lot of pictures. Many people pretend that they understand modern art.
-Placeholder Replacement (Xie et al., 2017) I am an art _ and I paint a lot of pictures. Many _ pretend that they understand modern art.
-Inversion (Min et al., 2020) I am an art student and a lot of pictures paint me. Many people pretend that they understand modern art.
-Passivization (Min et al., 2020) I am an art student and I paint a lot of pictures. Many people pretend that modern art are understood by them.
-5.1.3. Text Value-structure Manipulation
-This type of method perturbs both value and structure information of a sentence. And two typical methods are Back-translation Augmentation and Hierarchical Data Augmentation.
-Back-translation Augmentation.
-Back-translation (Sennrich et al., 2016) means translating text data from one language to another, or a sequence of other languages, and finally translating back to the original language.
-The semantic information is preserved during the translation process, but the process can introduce some variance in syntax and wording.
-Such transformation can help the model learn more flexible expressions that convey the same meaning.
-Hierarchical Data Augmentation.
-Hierarchical Data Augmentation (HDA) (Yu et al., 2019) augments text data at both word and sentence levels.
-It evaluates the roles of words and sentences by training a hierarchical attention network to obtain the attention values of the text data at both levels.
-The attention value is used to help decide the importance of components and augment the data by cropping and concatenating these components.
-5.2. Multiple Text Augmentation
-5.2.1. Value-based Text Mixture
-Mixup (Zhang et al., 2017) has been a widely used data augmentation approach especially used for image data.
-However, unlike image data, whose coloring information is expressed by numerical values and can be interpolated at the input layer, it is hard to calculate an interpolation for the words.
-So, text data are usually interpolated at embedding layers.
-Text Mixup.
-The adaptation of Mixup in sentence classification is studied in (Guo et al., 2019), which proposes wordMixup and senMixup.
-WordMixup interpolates a pair of word embeddings and then feeds the interpolation to the sentence encoder.
-SenMixup, on the other hand, interpolates a pair of sentence embeddings.
-The embedding vector produced by WordMixup or SenMixup is then passed to the softmax layer to produce the target distribution of the classification task.
-The use of mixup in text data augmentation has also been extensively studied in many research, such as nonlinear Mixup (Guo, 2020) and DoubleMix (Chen et al., 2022).
-A special example is GPT3Mix (Yoo et al., 2021), which makes use of the latest large language model to mix sentences.
-By inserting two source sentences in a prompt template that describes the instruction to the model, we can get a mixed sentence of high quality with a soft label.
-5.2.2. Structure-based Text Combination
-Some methods combine text data by merging text fragments from one sentence into another.
-Text Fragment Merging.
-Substructure Substitution (Shi et al., 2021) and Subtree Swapping (Dehouck and Gómez-Rodríguez, 2020) both use tree structures to analyze the part of speech or role in the sentences and thus make reasonable swaps.
-Some simpler method like GECA (Andreas, 2019) aims to remove a fragment from a sentence and populate it with another fragment that occurs in some common environment.
-5.3. Populational Text Augmentation
-5.3.1. Vanilla Text Generation
-We have discussed how augmented text data samples can be produced based on existing text materials.
-While modifying existing data is not the only solution to data augmentation.
-By leveraging various generative models, we can easily acquire high-quality text data on a large scale, and such abundance in data availability could greatly improve the performance of language models especially in terms of their generalization and robustness.
-Data augmentation by text generation previously relied on Traditional Models, while Large Language Models have now become the dominant technique.
-Traditional Models.
-The use in data augmentation of some early generative language models, such as seq2seq (Prakash et al., 2016), VAE (Bowman et al., 2015), and CVAE (Zhou and Wang, 2017) has been investigated in (Malandrakis et al., 2019), which shows improvement in the accuracy of an extrinsic low resource classification task.
-Large Language Models.
-Recent advances in large language models (LLM) and generative pre-training (GPT) (Radford et al., 2018) models have demonstrated unprecedented power in text understanding, classification, and generation tasks.
-Apart from supporting intelligent chat-bots that help people in their daily lives, the capacity of these state-of-the-art language generation models establishes a new paradigm of text data augmentation.
-LAMBADA (Anaby-Tavor et al., 2020) is among the first attempts to use a state-of-the-art language generator to synthesize new text data for data augmentation, and its procedure is as follows.
-LAMBADA first trains a baseline classifier from a small dataset, and then it fine-tunes the GPT-2 (Radford et al., 2019) model.
-A set of labels is thereafter fed to the fine-tuned language model, which generates a large dataset of labeled sentences.
-These sentences are evaluated by the classifier trained prior according to its confidence score, and top-ranked sentences are retained to compose the synthesized dataset.
-Similarly, (Quteineh et al., 2020) integrates GPT-2 in its text data augmentation process and improves the performance of active learning on tiny datasets.
-It uses the language model to generate a tree of tokens as candidates and construct a sentence from them.
-The set of new sentences is repeatedly refined by active learning, and the outcome is an augmented text dataset.
-Some other research also applies GPT-2 to effectively augment low-resource text datasets, such as (Yang et al., 2020) and (Edwards et al., 2022).
-Despite their success, both of them point out the need for more sophisticated methods to generate text data.
-(Yang et al., 2020) for example, argues that human-in-the-loop for relabeling might be rewarding.
-(Edwards et al., 2022) adopts human-in-the-loop for seed selection and further discusses the potential benefits of injecting expert knowledge for specialized domains.
-Notably, Data Boost (Liu et al., 2020) proposes a reinforcement learning framework to guide the process of text data augmentation with GPT-2 and succeeds in generating data of high quality and high similarity to the original data.
-The rise of GPT-3 (Brown et al., 2020) and later LLMs is now providing more possibilities for text generative augmentation.
-Their use and improvements are studied in the latest research such as (Balkus and Yan, 2022) and (Cochran et al., 2023).
-Table 6. A summary for text data augmentation
-Tier 1: Quantity of source data Tier 2: Information type Augmentation method Typical literature
-Individual Text Augmentation Value-based Transformation Invariant Replacement EDA (Wei and Zou, 2019) (EMNLP-IJCNLP 2019), (Marivate and Sefara, 2020) (CD-MAKE 2020), TEXTFOOLER (Jin et al., 2020a) (AAAI 2020), LetterTran (Liu et al., 2011) (ACL 2011), (Duan et al., 2023) (TASLP 2023), STA (Guo et al., 2022) (arXiv 2022)
-Token Addition EDA (Wei and Zou, 2019) (EMNLP-IJCNLP 2019), AEDA (Karimi et al., 2021) (EMNLP 2021), INSNET (Lu et al., 2022) (NeurIPS 2022), LeCA (Chen et al., 2021) (IJCAI 2021), GBS (Hokamp and Liu, 2017) (ACL 2017), STA (Guo et al., 2022) (arXiv 2022)
-Token Deletion EDA (Wei and Zou, 2019) (EMNLP-IJCNLP 2019), (Xie et al., 2017) (ICLR 2017), (Duan et al., 2023) (TASLP 2023), STA (Guo et al., 2022) (arXiv 2022)
-Structure-based Transformation Sentence Cropping Sentence cropping (Şahin and Steedman, 2018) (EMNLP 2018), STA (Guo et al., 2022) (arXiv 2022)
-Sentence Morphing EDA (Wei and Zou, 2019) (EMNLP-IJCNLP 2019), (Artetxe et al., 2018) (ICLR 2018), Sentence Rotation (Şahin and Steedman, 2018) (EMNLP 2018), Inversion & Passivization (Min et al., 2020) (ACL 2020), DRNNs (Xu et al., 2016) (COLING 2016)
-Value-structure Transformation Back-translation Back-translation (Sennrich et al., 2016) (ACL 2016)
-Hierarchical HDA (Yu et al., 2019) (IEEE Access 2019)
-Multiple Text Augmentation Value-based Mixture Text Mixup wordMixup & senMixup (Guo et al., 2019) (arXiv 2019), Nonlinear Mixup (Guo, 2020) (AAAI 2020), DoubleMix (Chen et al., 2022) (COLING 2022), GPT3Mix (Yoo et al., 2021) (EMNLP 2021)
-Structure-based Combination Text Fragment Merging Substructure Substitution (Shi et al., 2021) (IJCNLP 2021), GECA (Andreas, 2019) (arXiv 2019), Subtree Swapping (Dehouck and Gómez-Rodríguez, 2020) (COLING 2020)
-Populational Text Augmentation Vanilla Generation Traditional Models (Malandrakis et al., 2019) (EMNLP-IJCNLP 2019)
-LLMs LAMBADA (Anaby-Tavor et al., 2020) (AAAI 2020), (Quteineh et al., 2020) (EMNLP 2020), (Yang et al., 2020) (EMNLP 2020), (Edwards et al., 2022) (DaSH 2021), Data Boost (Liu et al., 2020) (EMNLP 2020), (Balkus and Yan, 2022) (Natural Language Engineering 2022), (Cochran et al., 2023) (ICAIE 2023)
-6. Data Augmentation for Graph Data
+Single-instance level augmentation for text data often changes the wording or reshapes the syntax of a sentence.
 A graph is represented by $\mathbf{G}=\{\mathcal{V},\mathcal{E}\}$, where $\mathcal{V}$ stands for the vertex set, and $\mathcal{E}$ stands for the edge set.
@@
 Hence another representation for a graph is $\mathbf{G}=\{\mathbf{A},\mathbf{X}\}$.
-Graph data augmentation is to perturb the values in the feature matrix $\mathbf{X}$ of nodes or the topology structure represented by the adjacency matrix $\mathbf{A}$.
-6.1. Individual Graph Augmentation
-6.1.1. Value-based Graph Transformation
-Graph nodes often contain rich attributes in the feature matrix. This matrix is generally similar to tabular data, so augmenting node attributes does not need to consider the specialty of graph data.
-Node Attributes Augmentation.
-Various typical augmentation approaches on node attributes from random masking (You et al., 2020; Feng et al., 2020; Thakoor et al., 2021) to perturbing attributes with random noises (Yang et al., 2021) can be applied, as we do for common tabular data.
-These methods are unified in Eq.
-6, where $\mathbf{M}\in\{0,1\}^{N\times N}$ is a binary mask, $\odot$ denotes the Hadamard product, and $\mathbf{N_{W}}\sim\mathcal{N}(0,\Sigma)$ is the Gaussian white noise.
-(6) $\mathbf{\tilde{X}}=\mathbf{X}\odot\mathbf{M}+\mathbf{N}_{W}$
-Some other methods are more sophisticated.
-For example, (Song et al., 2021) first learns the topology information of the graph by random walk, and uses the topology embedding together with the original features as inputs to a dual graph neural network.
-Similarly, LA-GNN (Liu et al., 2022d) applies a local augmentation mechanism.
-It samples features from nodes’ neighborhoods as additional information when augmenting the attributes.
-A different approach is proposed in DropMessage (Fang et al., 2023).
-Instead of directly modifying the attributes on the nodes, it applies random dropping in the message-passing phase.
-6.1.2. Structure-based Graph Transformation
-Graph topology structure has immense information representation capacity.
-Many augmentation methods designed for graph data focus on modifying the topology structure.
-Naive Topology Perturbation methods modify a single node or edge component or consider a set of nodes as a whole and modify the substructure.
-There is also Subgraphing that extracts subgraphs from the complete graph.
-Graph Diffusion, Graph Rewiring and Graph Structure Learning are more sophisticated techniques that focus on adjusting the edges to obtain a better graph.
-Naive Topology Perturbation.
-Due to its straightforward idea and easy implementation, node dropping and addition are among the first applied augmentation approaches in graph learning tasks.
-As given in Eq.
-7, Node dropping (You et al., 2020) randomly discards some nodes from the graph.
-Such an operation will also remove the attributes that these nodes carry and the edges on these nodes.
-(7) $\tilde{\mathcal{V}}=\text{Drop}(\mathcal{V}),\qquad\tilde{\mathbf{G}}=\{\tilde{\mathbf{A}},\tilde{\mathbf{X}}\}=\{\mathbf{A}[\tilde{\mathcal{V}},\tilde{\mathcal{V}}],\mathbf{X}[\tilde{\mathcal{V}},:]\}$
-In some special cases where node dropping is adopted, the edges on the removed nodes are not simply dropped.
-For example, in MoCL (Sun et al., 2021c), when dropping a general carbon, the two sides of the carbon atom are connected to ensure the completeness of the molecule.
-So, in MoCL, rather than being removed, the node is more of being “skipped”.
-As for node addition, an example is MPNN (Gilmer et al., 2017).
-It introduces a latent master node that is connected to all other nodes, which helps messages pass through the graph network.
-An edge describes the relationship between two nodes, and by perturbing the edges, we also introduce diversity to the structural information of a graph.
-DropEdge (Rong et al., 2019) perturbs the edge set by randomly setting a portion of the adjacency matrix to zero.
-It first produces a sparse matrix from a random subset of the original edges, then derives the augmented edge set by subtracting the sparse matrix from the original adjacency matrix.
-Edge masking in (Thakoor et al., 2021) constructs a Bernoulli distribution, and randomly masks edges under that probability.
-FairDrop (Spinelli et al., 2021) computes a mask according to the network’s homophily, and use this mask to guide the edge dropping process to remove unfair connections.
-AD-GCL (Suresh et al., 2021) applies a GNN-augmenter to learn how to drop edges so that the mutual information of the augmented graph and that of the original graph is minimized.
-Graph substructure perturbation is beneficial for molecular graphs.
-A molecule graph represents atoms with nodes and chemical bonds with edges.
-Functional groups, which often decide the chemical characteristics of molecules, can be represented by substructures.
-So, by wisely sampling substructures, the global semantics of a molecule graph can be well contained.
-MoCL (Sun et al., 2021c), for example, utilizes substructure information as one of its augmentation scheme.
-It performs graph augmentation by replacing functional groups, effectively introducing reasonable noises to the graph.
-GEAR (Liu et al., 2022e) also employs such technique.
-It augment molecule graphs by so-called environment replacement, which introduces noise from the environment subgraphs.
-Subgraphing.
-Subgraphing is based on the idea that graph semantics are partially represented in local structures (You et al., 2020), and some subgraphing methods, such as those used in GCC (Qiu et al., 2020) and SUGAR (Sun et al., 2021a) use random walk to sample a subgraph from the original graph.
-GrpahCrop (Wang et al., 2020) applies a node-centric strategy to crop connectivity-preserved subgraphs and expand the dataset with these augmented subgraphs.
-For each subgraph, the rest of the graph is removed, but compared with edge dropping and node dropping methods, GraphCrop (Wang et al., 2020) better maintains the topology characteristics of the original graph.
-Similarly, SUBG-CON (Jiao et al., 2020) uses the Personalized PageRank algorithm to select the important nodes.
-NeuralSparse (Zheng et al., 2020) proposes to obtain the subgraph by learning.
-It uses a sparsification network to generate sparsified subgraphs.
-Graph Diffusion.
-Traditional message passing in graph neural networks only leverages one-hop connections, while the edges of a real graph are often noisy or arbitrarily defined, and many relationships in the networks are indirectly contained in higher-order neighborhoods.
-To address this issue, GDC (Gasteiger et al., 2019), as in Eq.
-$\mathbf{\tilde{A}}=\sum_{k=0}^{\infty}\theta_{k}\mathbf{T}^{k}$ , proposes to calculate a series of transition matrices $\mathbf{T}^{k}$ by means such as random walk and combines these transition matrices with weights $\theta_{k}$ to obtain a new adjacency matrix $\mathbf{\tilde{A}}$ and thus generate a new graph that directly contains some node connection hidden in the original graph.
-The new graph can also be understood as being processed by a denoising filter because graph diffusion smooths out the neighborhood over the graph (Gasteiger et al., 2019).
-This technique is widely used by self-supervised approaches to transform the graph and generate different views for contrastive learning, such as MVGRL (Hassani and Khasahmadi, 2020), SelfGNN (Kefato and Girdzijauskas, 2021), and MERIT (Jin et al., 2021).
-Graph Rewiring.
-The topology structure has significant influences on the performance of GNN models.
-Message passing plays an indispensable role in GNN models.
-Still, this process suffers from the over-squashing problem when information is spreading over long distances in a graph with the bottleneck (Topping et al., 2021).
-Also, the over-smoothing problem leads to GNNs’ performance degeneration on heterophilous graphs (Guo et al., 2023).
-Moreover, the graph’s robustness to perturbations poses the stability problem to GNN models (Kenlay et al., 2021).
-To address these issues, a series of methods referred to as Graph Rewiring is proposed, which adjust the structure of the graphs to help models learn better graph representations and yield better inference in downstream tasks (Guo et al., 2023).
-Graph rewiring, as presented in Eq.
-8, is initially proposed as randomly swapping edges (Schneider et al., 2011) to improve a network’s robustness and mitigate malicious attacks.
-Some later research such as smart rewiring (Louzada et al., 2013) takes the degrees of chosen nodes when swapping their edges.
-Stochastic Discrete Ricci Flow (Topping et al., 2021) is a curvature-based graph rewiring method, which in each step adds an edge that can improve the most negatively curved edge, and then removes the edge with the most positive curvature.
-Its results demonstrate the ability of graph rewiring to alleviate the over-squashing problem.
-DiffWire (Arnaiz-Rodríguez et al., 2022) introduces the CT-LAYER and GAP-LAYER in a GNN model, which respectively learns the commute times and spectral gap of the graph.
-It rewires the adjacency graph so that the graph’s spectral gap is minimized, and by doing so tries to predict the optimal graph structure for the downstream task.
-DHGR (Bi et al., 2022) first learns a similarity matrix from the original graph.
-Then it rewires the graph by adding edges between high-similarity node pairs and removing edges from low-similarity ones.
-Such an operation can effectively improve the homophily of the graph and thus improve GNN performance.
-HDHGR (Guo et al., 2023) is a method that follows DHGR (Bi et al., 2022).
-It computes multiple similarity matrices under different meta-paths and thus generates several rewired meta-path subgraphs.
-(8) $\vspace{-2mm}\mathbf{\tilde{A}}=\overbrace{\mathbf{A}\odot(\mathbf{1}-\mathbbm{1}_{r})}^{\text{edge removal}}+\overbrace{(\mathbf{1}-\mathbf{A})\odot\mathbbm{1}_{r}}^{\text{edge addition}}$
-Graph Structure Learning.
-Graph Neural Networks are subject to noisy, wrong, or incomplete graph structures.
-Similar to graph rewiring, graph structure learning (GSL) also aims at simultaneously optimizing graph structure and its corresponding representations (Zhu et al., 2021) to deal with the above-mentioned problem, however, GSL is often implemented in a more complicated manner and discussed in a broader context than graph rewiring.
-GSL applies various techniques to learn and model graph structures, and it is often depicted with a pipeline that includes iterative learning and updating (Zhu et al., 2021).
-Also, unlike graph rewiring, which only rewires a graph, GSL can be used to learn and construct a graph structure from non-graph datasets (Zhu et al., 2021), so that many models specific to graph datasets, e.g.
-GNNs can be trained on these data.
-Based on how the graph structure is modeled, GSL can be categorized into three classes: metric-based, neural, and direct modeling (Zhu et al., 2021).
-Metric-based modeling evaluates the similarity between node pairs through some kernel functions, and in the new graph, the weights of the edges are decided by the similarities of the node pairs.
-Many early GSL methods fall into this group.
-For example, IDGL (Chen et al., 2020b) and HGSL (Zhao et al., 2021) both use cosine function for similarity learning to evaluate the optimal adjacency of nodes.
-Neural approaches use neural networks for both graph learning and graph convolution tasks, as presented in the work of GLCN (Jiang et al., 2019), which introduces a simple single-layer neural network to learn a new graph from the original node attributes before the graph data is learned by the graph convolution module.
-PTDNet (Luo et al., 2021) focuses on dropping noisy, task-irrelevant edges from the graph structure.
-VIB-GSL (Sun et al., 2022) learns graph structure from an information theory perspective.
-SUBLIME (Liu et al., 2022f) proposes an unsupervised learning paradigm for GSL that benefits more types of downstream task.
-Direct GSL approaches do not involve estimating edge weights from the vertex or generating a graph by training neural networks.
-Pro-GNN (Jin et al., 2020b) for example, directly updates the learned adjacency matrix itself during the training process.
-To avoid sub-optimal results, it jointly learns the graph structure and downstream GNN model.
-The learned graph and model parameters are updated sequentially in each training epoch.
-LDS-GNN (Franceschi et al., 2019) treats graph learning as a probabilistic distribution problem.
-When learning from the objectives, it updates the parameter of a Bernoulli distribution, and samples the learned graph from this distribution.
-6.2. Multiple Graph Augmentation
-6.2.1. Value-based Graph Mixture
-Feature or label information from different nodes can be mixed by Graph Propagation, which is often used for tasks such as node classification.
-Graph Propagation. Propagation refers to spreading information from one node to other nodes along the graph. This type of method can be divided into feature propagation and label propagation.
-Feature propagation shares node features with other nodes.
-After feature vectors of some nodes are masked out, GRAND (Feng et al., 2020) further augments the node features by randomly propagating them through edges.
-MV-GCN (Yuan et al., 2021) first creates three different views from the original graph and propagates the features of the three views respectively before their representations are combined.
-Label propagation algorithm (LPA) (Zhu and Ghahramani, 2002) propagates labels from labeled nodes to all unlabeled nodes according to their proximity through areas with dense unlabeled nodes.
-AutoGRL (Sun et al., 2021b) applies LPA in their semi-supervised framework to label more nodes utilizing a few ground truth nodes.
-(Huang et al., 2020) performs two types of label propagation, correcting the base predictions and smoothing the final prediction, respectively.
-Though not in a typical graph-based scenario, (Iscen et al., 2019) constructs a nearest neighbor graph on the classifier’s manifolds.
-The affinity matrix, which is the analogy to the adjacency matrix in this case, is constructed based on the $k$ nearest neighbors of each node’s embeddings, and this matrix decides how labels propagate.
-6.2.2. Graph Value-structure Mixutre
-Graph Mixup refers to mixing graph nodes or graph instances. This often involves both node attribute values and the topology structure information.
-Graph Mixup.
-The technique of mixup has been extensively applied in various augmentation approaches for image and text tasks.
-By interpolating feature vectors and labels, mixup expands the feature space of training data and effectively improves the model’s robustness and generalization ability.
-However, the application of mixup in graph data augmentation is challenged by the special structure of graph datasets.
-Due to the irregular connectivity and various topological structures of graph data, it is hard to directly combine nodes or graphs, as we do when combining two images.
-For node-level mixup, one solution to this problem is temporarily ignoring the edges and mixing node features in latent space.
-For example, GraphMix (Verma et al., 2021) interpolates the hidden states and labels of input nodes to train a better FCN layer, whose parameters are shared with the GNN layer.
-A topology-involved node mixup method is proposed in (Wang et al., 2021).
-It randomly pairs nodes and includes the nodes’ local topology representations when interpolating them.
-Notwithstanding the topology information’s involvement in (Wang et al., 2021), it dose not mentions the construction of edges for the new node.
-S-Mixup (Structural Mixup) (Kim et al., 2023) also considers neighboring edges when pairing nodes but works in the input space and discusses how to connect the node with the original graph with edges.
-It matches nodes based on their classification confidence of a GNN classifier and selects edges with high gradients to connect with the mixed node.
-(Wang et al., 2021) also proposes a method for graph-level mixup.
-It uses GNNs to embed the complex and irregular graph structures and conduct mixup in the embedding space.
-However, this method cannot mix graphs in the input layer.
-To directly mix two inputs at the graph level, ifMixup (Guo and Mao, 2021) adds dummy nodes so that two graphs have the same number of nodes.
-These dummy nodes have all-zero feature vectors and are not connected to any other nodes, or in other words, the weights of edges on dummy nodes are zeros.
-When mixing the graph, the interpolation of a normal, one-hot labeled edge and a zero-weighted edge yields an edge with a soft label.
-Similarly, S-Mixup (Graph Mixup with Soft Alignments) (Ling et al., 2023) also aligns graphs first before mixing them.
-It obtains an assignment matrix by a soft assignment and uses the assignment matrix to transform a graph so that every node corresponds with a node in the other graph.
-Another approach that mixes graph-structured datasets at the graph level and input space is Graph Transplant (Park et al., 2022).
-It extracts a salient subgraph from the source graph and a random subgraph from the destination graph.
-The augmented graph is produced by transplanting the source subgraph to the destination subgraph and adaptively interpolating the labels of the original graphs.
-Compared with the above-mentioned approaches, G-Mixup (Han et al., 2022) proposes a more sophisticated way to interpolate graph data.
-It first uses graphs within the same class to estimate a graphon, i.e.
-graph generator:
-(9) $\mathbf{\mathcal{G}}\rightarrow W_{\mathbf{\mathcal{G}}},\mathbf{\mathcal{H}}\rightarrow W_{\mathbf{\mathcal{H}}}$
-The graphon $W:[0,1]^{2}\rightarrow[0,1]$ is a matrix, where $W(i,j)$ describes the probability of an edge between nodes $i$ and $j$, and graphs of the same class are considered to be generated from the same graphon.
-Graphons of two classes of graphs are then interpolated to produce a mixed graphon:
-(10) $W_{\mathbf{\mathcal{I}}}=\lambda W_{\mathbf{\mathcal{G}}}+(1-\lambda)W_{\mathbf{\mathcal{H}}}$
-By sampling from the graphon mixup $W_{\mathbf{\mathcal{I}}}$, we acquire a series of graphs with the same soft labels:
-(11) $\mathbf{y}_{\mathbf{\mathcal{I}}}=\lambda\mathbf{y}_{\mathbf{\mathcal{G}}}+(1-\lambda)\mathbf{y}_{\mathbf{\mathcal{H}}}$
-At the graphon estimation phase, G-Mixup aligns the original node features and obtains the graphon node features average pooling, which are used as node features of the generated graphs.
-DAGAD (Liu et al., 2022b) mixes graph representations by performing random permutations and it shows that data augmentation is very helpful in scenarios such as graph anomaly detection.
-6.3. Populational Graph Augmentation
-6.3.1. Vanilla Graph Generation
-Despite that Variational Autoencoders (VAEs) and Generative Adversarial Networks (GANs) can all be used to generate graph data, they face many problems in practice.
-VAE models require things like a massive graph-matching process, and GANs on graphs easily fall into mode collapse (Liu et al., 2023a).
-So, graph data augmentation resort to another recently proposed technique, namely, Generative Diffusion Models.
-Generative Diffusion Models
-The denoising diffusion model (Ho et al., 2020) is a generative paradigm that has become popular in recent years.
-Though not yet widely applied as an approach to graph data augmentation, the Generative Diffusion Model boasts enormous potential in graph generation and can possibly be adopted as a promising data augmentation technique.
-To implement diffusion models on graph data, some properties of the graph data must be properly treated.
-For example, diffusion models often rely on a continuous Gaussian noise process, but the graph structures are discrete (Zhang et al., 2023).
-DiGress (Vignac et al., 2023) designs a discrete denoising diffusion model for graph generation and achieves state-of-the-art performance on molecular and non-molecular datasets.
-The use of Generative Diffusion Models for data augmentation especially benefits studies on molecules, proteins, and materials, whose data are often in the form of graph instances.
-Table 7. A summary for graph data augmentation
-Tier 1: Quantity of source data Tier 2: Information type Augmentation method Typical literature
-Individual Graph Augmentation Value-based Transformation Node Attributes Augmentation Attribute Masking (You et al., 2020) (NeurIPS 2020), GRAND (Feng et al., 2020) (NeurIPS 2020), BGRL (Thakoor et al., 2021) (ICLR 2021), GASSL (Yang et al., 2021) (NeurIPS 2021), Topological Regularization (Song et al., 2021) (arXiv 2021), LA-GNN (Liu et al., 2022d) (ICML 2022), DropMessage (Fang et al., 2023) (AAAI 2023)
-Structure-based Transformation Naive Topology Perturbation Node Dropping & Edge perturbation (You et al., 2020) (NeurIPS 2020), MPNN (Gilmer et al., 2017) (ICML 2017), BGRL (Thakoor et al., 2021) (ICLR 2021), DropEdge (Rong et al., 2019) (ICLR 2019), AD-GCL (Suresh et al., 2021) (NeurIPS 2021), FairDrop (Spinelli et al., 2021) (TAI 2021)
-Subgraphing Subgraph (You et al., 2020) (NeurIPS 2020), GCC (Qiu et al., 2020) (KDD 2020), SUGAR (Sun et al., 2021a) (WWW 2021), GrpahCrop (Wang et al., 2020) (arXiv 2020), SUBG-CON (Jiao et al., 2020) (ICDM 2020), NeuralSparse (Zheng et al., 2020) (ICML 2020), MoCL (Sun et al., 2021c) (KDD 2021), GEAR (Liu et al., 2022e) (KDD 2021)
-Graph Diffusion GDC (Gasteiger et al., 2019) (NeurIPS 2019), MVGRL (Hassani and Khasahmadi, 2020) (ICML 2020), SelfGNN (Kefato and Girdzijauskas, 2021) (arXiv 2021), MERIT (Jin et al., 2021) (arXiv 2021)
-Graph Rewiring (Schneider et al., 2011) (PNAS 2011), Smart Rewiring (Louzada et al., 2013) (J.
-Complex Netw.
-2013), SDRF (Topping et al., 2021) (ICLR 2021), DiffWire (Arnaiz-Rodríguez et al., 2022) (LoG 2022), DHGR (Bi et al., 2022) (arXiv 2022), HDHGR (Guo et al., 2023) (WWW 2023)
-Graph Structure Learning IDGL (Chen et al., 2020b) (NeurIPS 2020), HGSL (Zhao et al., 2021) (AAAI 2021), GLCN (Jiang et al., 2019) (CVPR 2019), PTDNet (Luo et al., 2021) (WSDM 2021), VIB-GSL (Sun et al., 2022) (AAAI 2022), SUBLIME (Liu et al., 2022f) (WWW 2022), Pro-GNN (Jin et al., 2020b) (KDD 2020), LDS-GNN (Franceschi et al., 2019) (ICML 2019)
-Multiple Graph Augmentation Value-based Mixture Propagation Label propagation (Zhu and Ghahramani, 2002) (2002), GRAND (Feng et al., 2020) (NeurIPS 2020), MV-GCN (Yuan et al., 2021) (CIKM 2021), AutoGRL (Sun et al., 2021b) (IJCNN 2021), C&S (Huang et al., 2020) (ICLR 2020), (Iscen et al., 2019) (CVPR 2019)
-Value-structure Mixture Graph Mixup GraphMix (Verma et al., 2021) (AAAI 2021), (Wang et al., 2021) (WWW 2021), Structural Mixup (Kim et al., 2023) (CIKM 2023), ifMixup (Guo and Mao, 2021) (arXiv 2021), Graph Mixup with Soft Alignments (Ling et al., 2023) (arXiv 2023), Graph Transplant (Park et al., 2022) (AAAI 2022), G-Mixup (Han et al., 2022) (ICML 2022), DAGAD (Liu et al., 2022b) (ICDM 2022)
-Populational Graph Augmentation Vanilla Generation Diffusion Models DiGress (Vignac et al., 2023) (ICLR 2023)
-7. Data Augmentation for Tabular Data
+Single-instance level augmentation for graph data is to perturb the values in the feature matrix $\mathbf{X}$ of nodes or the topology structure represented by the adjacency matrix $\mathbf{A}$.
 A table is made up of cells, which store numerical or categorical values that represent attributes of the described entities.
@@
 So, most data augmentation methods leverage the value information.
-7.1. Individual Tabular Augmentation
-7.1.1. Value-based Table Transformation
+Still, it is also possible to perturb the structure information of a table.
+Time-series data represent sequential values within a period, and each timestamp is considered an elemental component.
+Data augmentation methods for time-series data either perturb the value at each timestamp or perturb the sequence of timestamps.
+That is, single-instance level augmentation methods focus on utilizing the values and structures of these components.
+IV-A Value-based Transformation
+IV-A1 Color Perturbation
+Pixel Erasing masks out all color information in an area of the data; Photometric Transformation imposes a soft adjustment to change the appearance of the image, but the essential information remains visible.
+Pixel Erasing
+As illustrated in Fig.
+2b, the idea of pixel erasing is to mask some information from the original input space or the feature space.
+Cutout [9] and random erasing [10] first selects a rectangle region and erases all pixels in that region.
+The erased region is then filled with a color or Gaussian noise.
+Random Erasing typically adds one or two big rectangular masks to the image, and some other methods deploy more flexible implementations.
+Hide-And-Seek [38] randomly removes multiple parts of the image so that the model has to learn from multiple parts of the image other than the most discriminating part.
+Similarly, the Grid-Mask method [39] generates a set of squared masks in the grid layout.
+It can prevent excessive deletion and reservation and achieve a reasonable balance between deleted and reserved regional information.
+In most research, random erasing is used in the spatial domain, while REF [40] extends the use of random erasing to the frequency domain.
+It uses the discrete Fourier transform to process the image, applies random erasing on the image’s frequency domain, and uses the inverse discrete Fourier transform to recover the image to the spatial domain.
+Photometric Transformation
+Augmentation that changes the overall color and style of the image is called photometric transformation, as shown in Fig 2c.
+The simplest way is to adjust the brightness of the image.
+Color inversion replaces the pixels with their “opposite” color.
+Color casting applies different operations in different channels, for example, increasing the image’s red channel while keeping the green channel and blue channel untouched.
+Noise injection adds Gaussian or other types of noises to the images.
+Greyscaling transforms a colored picture into black-and-white by combining channels for different colors, forcing the model to focus on the shapes or edges of target objects instead of their colors.
+Histogram equalization [41, 42] enhances the image by enhancing its contrast, and white balancing simulates alternative light cast on the object.
+Kernel filters are matrices that slide across and are convoluted with the original image to generate new images, either sharpening or blurring the image.
+Figure 2: A conceptual demonstration of (a) Original Image, (b) Pixel Erasing, (c) Photometric Transformation, (d) Image Cropping, (e) Geometric Transformation, (f) Policy-based Data Augmentation (g) Prompt-based Image Editing.
+IV-A2 Wording Perturbation
+Text data augmentation of this type uses word-level manipulation. It applies changes to the choice of words or adds or removes words. Some examples are presented in Table II.
+Token Replacement
+Token replacement refers to using alternative wording in a given sentence.
+EDA [11] proposes a set of text augmentations, including synonym replacement.
+It randomly selects several words from a sentence, substituting them with their synonyms, i.e.
+words that mean exactly or nearly the same.
+[43] uses WordNet [44], a lexical database to find synonyms for verbs and nouns in sentences and thus make replacements.
+It also derives word embeddings by Word2Vec [45] and replaces a word with others that have high cosine similarity.
+Another possibility is to use a knowledge graph.
+A knowledge graph describes the relationships of entities, and by utilizing the “is equivalent” relation, we can find a series of synonyms for a given word.
+Token Addition
+Token addition refers to adding tokens to the original text, either introducing junk tokens as noise or enriching the sentences and making them more informative.
+The random insertion, proposed in EDA [11], inserts a random synonym of a random word at a random position in the sentence.
+The words added in the EDA play the role of data noising and may carry either little or harmful information.
+As a variant of EDA, the technique of AEDA [46] only inserts punctuation marks.
+Token Deletion
+Random deletion [11] randomly removes some words from a sentence.
+In [47], noising techniques are used to avoid overfitting, and the randomly selected word is replaced by a blank placeholder token “_”, rather than simply being removed.
+In [48], apart from word dropout which is similar to random deletion, it also uses word blanking to replace a word with a special “¡BLANK¿” placeholder.
+IV-A3 Node Attributes Perturbation
+Graph nodes often contain rich attributes in the feature matrix. This matrix is generally similar to tabular data, so augmenting node attributes does not need to consider the specialty of graph data.
+Attribute Masking and Noise Injection
+Various typical augmentation approaches on node attributes from random masking [49, 50, 51] to perturbing attributes with random noises [52] can be applied, as we do for common tabular data.
+These methods are unified in Eq.
+5, where $\mathbf{M}\in\{0,1\}^{N\times N}$ is a binary mask, $\odot$ denotes the Hadamard product, and $\mathbf{N_{W}}\sim\mathcal{N}(0,\Sigma)$ is the Gaussian white noise.
+$\mathbf{\tilde{X}}=\mathbf{X}\odot\mathbf{M}+\mathbf{N}_{W}$ (5)
+Some other methods are more sophisticated.
+For example, [53] first learns the topology information of the graph by random walk, and uses the topology embedding together with the original features as inputs to a dual graph neural network.
+Similarly, LA-GNN [54] applies a local augmentation mechanism.
+It samples features from nodes’ neighborhoods as additional information when augmenting the attributes.
+A different approach is proposed in DropMessage [55].
+Instead of directly modifying the attributes on the nodes, it applies random dropping in the message-passing phase.
+IV-A4 Cell Value Perturbation
 Tabular data is much simpler compared with other modalities, so it is hard to design complex augmentation methods that make use of spatial or sequential relationships within a data sample as we do for image or text data, and most data augmentation for tabular data perturbs the value stored in the table cells.
@@
 Another approach to tabular data augmentation borrows ideas from Feature Engineering.
-Some work on Feature Engineering design automatic feature selection and transformation frameworks to generate optimal feature space, and the operations can also be used to augment tabular datasets.
-Table Masking.
-VIME (Yoon et al., 2020) uses a mask generator to corrupt features.
+Some work on Feature Engineering automates feature selection and transformation to generate optimal feature space, and the operations can also be used to augment tabular datasets.
+Table Masking
+VIME [56] uses a mask generator to corrupt features.
 In a self-supervised scenario, two pretext tasks respectively reconstruct the original data and estimate the mask vector.
@@
 The two losses jointly contribute to the back-propagation phase.
-SCARF (Bahri et al., 2021) applies data augmentation in a contrastive learning framework.
+SCARF [57] applies data augmentation in a contrastive learning framework.
 Its corruption process is similar to that of VIME but it uses InfoNCE loss to evaluate the similarity between the embeddings of the corrupted and original views and then update the model accordingly.
-MET (Majmundar et al., 2022) works in a pattern similar to VIME and SCARF, which calculates the reconstruction loss of a partially masked sample.
-MTR (Onishi and Meguro, 2023) randomly replaces part of the tokenized embeddings, instead of working directly in the input space.
-Feature Engineering.
-(Khurana et al., 2018) continuously generates new features and records them in a transformation graph.
-It uses reinforcement learning to select the feature set with best cross-validation performance.
-Some methods formulate the feature selection problem into a multi-agent reinforcement learning task (Liu et al., 2019, 2021a; Fan et al., 2020).
-They assign each feature to an agent, and the agents learn from the reward whether to select or unselect the features.
-However, the use of multi-agent reinforcement learning increases the computational burden and hardware cost, and some single-agent frameworks are further proposed for this problem (Liu et al., 2021b, 2023b).
-In these methods, only one agent is employed to traverse the whole feature set using the Monte Carlo algorithm and early stopping.
-Compared to the above-mentioned automatic feature selection methods, RAFT (Wang et al., 2022) and GRFG (Xiao et al., 2023a) emphasize feature transformation and generation.
-Unlike latent representation learning methods, they use arithmetic operations to generate new features, providing better traceability and explainability.
-GAINS (Xiao et al., 2023b) and MOAT (Wang et al., 2023) extend feature selection and transformation to continuous space.
-In GAINS (Xiao et al., 2023b), the selected feature subset is encoded into a continuous embedding space.
-It searches for better embeddings in the learned space and uses the decoder to generate an optimal selection of the dataset.
-MOAT (Wang et al., 2023) adapts this process to a feature transformation scenario, where features go through arithmetic operations transformation before the continuous embedding and searching stage, and the decoded features are also processed by a series of transformations.
-TFWT (Zhang et al., 2024) computes weights for samples and features, and the weights are combined with the data for later tasks.
+MET [58] works in a pattern similar to VIME and SCARF, which calculates the reconstruction loss of a partially masked sample.
+Feature Engineering
+Feature transformation [59, 60] and feature generation techniques can effectively enlarge tabular datasets by generating tables with different feature sets [61].
+Powered by reinforcement learning, GRFG [62] transforms existing features with cascading agents.
+It uses arithmetic operations to provide better traceability and explainability.
+MOAT [63] extends feature transformation to continuous space.
+It trains an encoder to project the feature transformation process to continuous embeddings and decodes the optimized embeddings for the best feature transformation strategies.
+LFG [64] and TIFG [65] leverage LLM agents to generate new features that are both explainable and effective.
+TFWT [66] computes weights for samples and features, and the weights are combined with the data for later tasks.
 Although these studies focus on feature engineering, their way of thinking and application of reinforcement learning is very enlightening for data augmentation tasks.
-7.1.2. Structure-based Table Transformation
+IV-A5 Sequence Value Perturbation
+The value information of a sequence is carried on the Y-axis, represented by the amplitude at specific timestamps or frequency channels.
+Amplitude Perturbation augments time-series data by altering the Y-axis values, and Decomposition decomposes the value of a sequence into different parts that express the sequential values from different aspects.
+Amplitude Perturbation
+Jittering adds random noise to the value at each timestamp of a sequence. In [67], jittering is used to simulated additive noise $\epsilon$ of wearable sensors:
+$\tilde{x}(\epsilon)=\{x_{1}+\epsilon_{1},\cdots,x_{t}+\epsilon_{t},\cdots,x_{T}+\epsilon_{T}\}$ (6)
+Scaling [67] alters the data by multiplying the magnitude of the entire sequence with a scalar $\alpha$.
+$\tilde{x}(\alpha)=\{\alpha x_{1},\cdots,\alpha x_{t},\cdots,\alpha x_{T}\}$ (7)
+Magnitude warping [67] is similar to scaling, but it uses a smooth curve to warp the sequence so that magnitude at different timestamps is scaled with different parameters.
+$\tilde{x}(\alpha)=\{\alpha_{1}x_{1},\cdots,\alpha_{t}x_{t},\cdots,\alpha_{T}x_{T}\}$ (8)
+Rotation [67], or flipping, assumes the data is symmetric with respect to the X-axis, and makes the data upside down by applying a rotation matrix $R$.
+$\tilde{x}(R)=\{Rx_{1},\cdots,Rx_{t},\cdots,Rx_{T}\}$ (9)
+Apart from the amplitude, the phase spectrum provides another option for data augmentation.
+RobustTAD [68] converts the time-series data from the time domain to the frequency domain, as shown in Eq.
+10, and then increases all phases $\theta(\omega_{k})$ by a small perturbation.
+$F(\omega_{k})=\frac{1}{N}\sum^{N-1}_{t=0}{x_{t}e^{-j\omega_{k}t}}=A(\omega_{k})exp[j\theta(\omega_{k})]$ (10)
+Decomposition
+Empirical Mode Decomposition (EMD) [69] is used to produce a set of intrinsic mode functions (IMFs).
+This decomposition relies on extracting the energy linked to different time scales that are inherent to the data.
+EMD has been utilized as a method of data augmentation in [70] for the purpose of categorizing impact noise in vehicles.
+Another way to decompose time-series data is STL [71], which decomposes signals into seasonal, trend, and remainder:
+$y_{t}=\tau_{t}+s_{t}+r_{t},\quad t=1,2,\cdots N$ (11)
+Other works such as [72] and [68] focus on improving robustness with the help of STL components.
+IV-B Structure-based Transformation
+IV-B1 Shape Transformation
+This type of data augmentation manipulates the spatial relationship between pixels.
+Image Cropping breaks the original image into pieces (Fig.
+2d), and Geometric Transformation perturbs the structure within the original data (Fig.
+2e).
+Image Cropping
+Image cropping selects an area from the original image and removes other parts of the image.
+It includes manual selection, random cropping, attention-based, aesthetics-based, and change-based approaches [73].
+Attention-based approaches use certain attention scores to evaluate a subimage.
+Regions that contain the main topic or other types of information often have high attention scores.
+[74] uses the entropy, area ratio, and center distance of the image to evaluate the performance of the image segmentation and accordingly adjust the cropping parameters.
+[75] uses eye-tracking techniques to capture human fixations and thus infer the location of important contents on the image.
+Aesthetics-based approaches [76, 77, 78] are similar to quality evaluation.
+For example, [76] detects several subject regions and one background region, extracts features from the regions, estimates and combines posterior probabilities, and determines the quality score according to these probabilities.
+A2-RL [79] integrates image cropping into a sequential decision-making process and an aesthetics-aware reinforcement learning model to solve the best cropping problem.
+Change-based approaches such as [80] take into account the content that is removed and extract two sets of specially designed features to model what is changed after cropping.
+Geometric Transformation
+As one of the most classical image data augmentation methods, geometric transformation changes the spatial relationship of pixels within an image while keeping the value of each pixel untouched to create more data without extra sampling or labeling.
+Affine transformations like flipping, shearing, rotation, and translation manipulate the image while preserving the colinearity and ratios of distances.
+Image flipping reflects the image around its vertical axis, horizontal axis, or both axes, rotation turns the image by an angle about a fixed point.
+Translation means moving an image in a certain direction.
+It can introduce more variance in the position of target objects in the image.
+In non-affine transformation, the spatial relationship between pixels is further changed compared with affine transformations because the colinearity and ratio of distances are broken.
+Perspective transformation [81] mimics images taken from other views of the object.
+Image stretching [82] preserves or even enlarges the prominent contents in the image and compresses the background or unimportant foreground information.
+Distortion [34, 83] also stretches the pixels but often focuses on a certain region.
+PatchShuffle [84] divides an image into block matrices and shuffles the pixels within each block under a certain probability.
+IV-B2 Syntax Transformation
+We categorize syntax transformation as: Sentence Cropping that splits a sentence into many pieces, and Sentence Morphing that swaps the location of words or grammatical components.
+Some examples are presented in Table II.
+Sentence Cropping
+Sentence cropping [85] divides sentences into several small pieces.
+It often uses dependency trees to extract the relationships between subjects and objects of the sentence and produces multiple smaller sentences that focus on different parts of the original sentence.
+This operation breaks the syntactic solidity and semantic comprehensiveness of the entire sentence but preserves local syntactic tags and shallow semantic labels.
+Sentence Morphing
+These methods manipulate sentence structure by altering the position of words or terms.
+Random swap [11] randomly chooses two words in a sentence and swaps their position.
+This perturbs the structure and adds noise to the text data across paragraphs.
+[86] studies two types of syntactic augmentations: inversion and passivization.
+Inversion swaps the subject and object in a sentence, and the semantics might be different in most cases.
+Passivization converts a sentence to its passive form, and the meaning of the new sentence is typically consistent with the original one.
+Sentence rotation [85] borrows the idea of image rotation, where the image is rotated around its center.
+In sentence rotation, the root of the sentence is chosen as the center of the sentence, and other sentence fragments are connected to their center via the dependency tree.
+TABLE II: Demonstration of some elementary operations for text data augmentation
+Operation Sentence
+Original I am an art student and I paint a lot of pictures. Many people pretend that they understand modern art.
+Synonym Replacement [11] I am an art student and I draw a lot of pictures. Many people pretend that they understand contemporary art.
+Similarity Replacement [43] I am an art student and I paint a lot of photos. Many people pretend that they understand modern literature.
+Random Insertion [11] I am an art student and picture I paint a lot of pictures. Many appreciate people pretend that they understand modern art.
+Punctuation Insertion [46] I am an, art student and: I paint a lot of pictures. Many people? pretend that they. understand modern art.
+Random Swap [11] I am an of student and I paint a lot art pictures. Many people they that pretend understand modern art.
+Random Deletion [11] I am an art student and I paint a lot of pictures. Many people pretend that they understand modern art.
+Placeholder Replacement [47] I am an art _ and I paint a lot of pictures. Many _ pretend that they understand modern art.
+Inversion [86] I am an art student and a lot of pictures paint me. Many people pretend that they understand modern art.
+Passivization [86] I am an art student and I paint a lot of pictures. Many people pretend that modern art are understood by them.
+IV-B3 Topology Transformation
+Graph topology structure has immense information representation capacity.
+Naive Topology Perturbation methods modify a single node or edge component or consider a set of nodes as a whole and modify the substructure.
+Subgraphing extracts subgraphs from the complete graph.
+Graph Diffusion and Graph Rewiring are more sophisticated techniques that focus on adjusting the edges to obtain a better graph.
+Naive Topology Perturbation
+Due to their straightforward idea and easy implementation, node dropping and node addition are among the earliest augmentation approaches in graph learning.
+As Eq.
+12 shows, Node dropping [49] randomly discards some nodes from the graph.
+Such an operation will also remove the attributes that these nodes carry and the edges on these nodes.
+$\tilde{\mathcal{V}}=\text{Drop}(\mathcal{V}),\qquad\tilde{\mathbf{G}}=\{\tilde{\mathbf{A}},\tilde{\mathbf{X}}\}=\{\mathbf{A}[\tilde{\mathcal{V}},\tilde{\mathcal{V}}],\mathbf{X}[\tilde{\mathcal{V}},:]\}$ (12)
+As for node addition, an example is MPNN [87].
+It introduces a latent master node that is connected to all other nodes, which helps messages pass through the graph network.
+An edge describes the relationship between two nodes, and by perturbing these relationships through edge dropping and edge addition, we also introduce diversity to the structural information of a graph.
+DropEdge [88] perturbs the edge set by randomly setting a portion of the adjacency matrix to zero.
+It first produces a sparse matrix from a random subset of the original edges, then derives the augmented edge set by subtracting the sparse matrix from the original adjacency matrix.
+Edge masking in [51] constructs a Bernoulli distribution, and randomly masks edges under that probability.
+FairDrop [89] computes a mask according to the network’s homophily, and uses this mask to guide the dropping process to remove unfair connections.
+AD-GCL [90] applies a GNN-augmenter to learn how to drop edges so that the mutual information of the augmented graph and that of the original graph is minimized.
+Subgraphing
+Subgraph cropping is based on the idea that graph semantics are partially represented in local structures [49].
+Some subgraphing methods, such as those used in GCC [91] and SUGAR [92] use random walk to sample a subgraph from the original graph.
+GrpahCrop [93] applies a node-centric strategy to crop connectivity-preserved subgraphs and expand the dataset with these augmented subgraphs.
+These subgraphs are able to maintain some topology characteristics of the original graph.
+Similarly, SUBG-CON [94] uses the Personalized PageRank algorithm to select the important nodes.
+NeuralSparse [95] proposes to obtain the subgraph by learning.
+It uses a sparsification network to generate sparsified subgraphs.
+Subgraph Substitution perturbs the substructure as a whole of a graph, which is especially beneficial for molecular graphs.
+A molecule graph represents atoms with nodes and chemical bonds with edges.
+Functional groups, which often decide the chemical characteristics of molecules, can be represented by substructures.
+So, by wisely sampling substructures, the global semantics of a molecule graph can be well contained.
+MoCL [96], for example, utilizes substructure information as one of its augmentation schemes.
+It performs graph augmentation by replacing functional groups, effectively introducing reasonable noise.
+GEAR [97] also employs such a technique.
+It augments molecule graphs by environment replacement, introducing noise from environment subgraphs.
+Graph Diffusion
+Traditional message passing in graph neural networks only leverages one-hop connections, while the edges of a real graph are often noisy or arbitrarily defined, and many relationships in the networks are indirectly contained in higher-order neighborhoods.
+To address this issue, GDC [98], as in Eq.
+$\mathbf{\tilde{A}}=\sum_{k=0}^{\infty}\theta_{k}\mathbf{T}^{k}$ , proposes to calculate a series of transition matrices $\mathbf{T}^{k}$ by means such as random walk and combines these transition matrices with weights $\theta_{k}$ to obtain a new adjacency matrix $\mathbf{\tilde{A}}$ and thus generate a new graph that directly contains some node connection hidden in the original graph.
+The new graph can also be understood as being processed by a denoising filter because graph diffusion smooths out the neighborhood over the graph [98].
+This technique is widely used by self-supervised approaches to transform the graph and generate different views for contrastive learning, such as MVGRL [99], SelfGNN [100], and MERIT [101].
+Graph Rewiring
+Graph rewiring, as presented in Eq.
+13, is initially proposed as randomly swapping edges [102] to improve a network’s robustness and mitigate malicious attacks.
+Some later research such as smart rewiring [103] takes the degrees of chosen nodes when swapping their edges.
+Stochastic Discrete Ricci Flow [104] is a curvature-based graph rewiring method, which in each step adds an edge that can improve the most negatively curved edge, and then removes the edge with the most positive curvature.
+Its results demonstrate the ability of graph rewiring to alleviate the over-squashing problem.
+DiffWire [105] introduces the CT-LAYER and GAP-LAYER in a GNN model, which respectively learns the commute times and spectral gap of the graph.
+It rewires the adjacency graph so that the graph’s spectral gap is minimized, and by doing so tries to predict the optimal graph structure for the downstream task.
+DHGR [106] first learns a similarity matrix from the original graph.
+Then it rewires the graph by adding edges between high-similarity node pairs and removing edges from low-similarity ones.
+Such an operation can effectively improve the homophily of the graph and thus improve GNN performance.
+Following DHGR [106], HDHGR [107] computes multiple similarity matrices under different meta-paths and thus generates several rewired meta-path subgraphs.
+$\vspace{-2mm}\mathbf{\tilde{A}}=\overbrace{\mathbf{A}\odot(\mathbf{1}-\mathbbm{1}_{r})}^{\text{edge removal}}+\overbrace{(\mathbf{1}-\mathbf{A})\odot\mathbbm{1}_{r}}^{\text{edge addition}}$ (13)
+IV-B4 Table Structure Transformation
 Table Subsetting
-Inspired by image cropping (Chen et al., 2016; Yan et al., 2013), SubTab (Ucar et al., 2021) divides a table into multiple subsets by splitting its columns.
+Inspired by image cropping [73, 80], SubTab [108] divides a table into multiple subsets by splitting its columns.
 Each subset contains only a percentage of all features.
 The model is forced to learn the representation of the samples by these features and is required to reconstruct all the features from the subset.
-7.2. Multiple Tabular Augmentation
-7.2.1. Value-based Tabular Mixture
-Mixup is the technique to interpolate samples to produce new data. Tabular data, due to its will-aligned structure, is suitable for this augmentation approach.
-Table Mixup.
-SMOTE (Chawla et al., 2002) is among the first works to study data interpolation, even before the proposal of Mixup (Zhang et al., 2017).
+IV-B5 Sequence Transformation
+Unlike Value Perturbation, sequence perturbation is implemented on the X-axis, either on the time domain or the frequency domain.
+Window Slicing splits the time-series data into pieces.
+It is also the basis of many Sequence Morphing methods because they apply varied operations on these slices.
+Window Slicing
+Inspired by image cropping, which is widely used for image data augmentation, MCNN [109] proposes window slicing.
+It extracts slices from the entire sequence as new training samples.
+Each slice is assigned with the label of the original sequence.
+Sequence Morphing
+Based on Window Slicing, [110] extends or compresses the generated spans, which is called window wrapping.
+This is equivalent to speeding up or slowing down a span of the sequence.
+Vowel stretching [111] is a special kind of window wrapping technique.
+It only perturbs the sequence by prolonging the vowels.
+This is meant to be consistent with how children speak.
+[67] propose to randomly permeate these slices and concatenate them as a long sequence.
+Similar processing is also called Random Shuffling and is used in [112].
+VTLP [113] applies window wrapping in the frequency domain.
+It adds a small distortion to the central frequency.
+SpecAugment [114] also proposes to augment the frequency domain.
+It generates masks of length sampled from a uniform distribution and uses each of them to mask consecutive frequency channels.
+IV-C Value-structure Transformation
+IV-C1 Image Value-structure Transformation
+Policy-based methods search for the optimized augmentation policy that includes both value-based transformations and structure-based transformations, as demonstrated in Fig.
+2f.
+Prompt-based Image Editing employs a text-to-image model to change the content, background, or style of the image as humans desire.
+When implementing the augmentation, both coloring and the positions of pixels are changed (Fig.
+2g).
+Policy-based Data Augmentation
+AutoAugment [13] represents data augmentation as a decision-making process and uses reinforcement learning to search for the best operations.
+In AutoAugment, a policy includes a group of basic data augmentation operations with probabilities and magnitudes.
+AutoAugment processes images under the policy, evaluates these augmented images on downstream networks, and rewards a policy generator for better performance.
+Fast AutoAugment [115] proposes a more efficient searching algorithm for augmentation policies, which works significantly faster than AutoAugment and thus allows a larger number of sub-policies, which should benefit the generalization performance.
+RandAugment [116] removes the need for a separate search phase on a proxy task, and makes the training process less complicated and less expensive.
+Population Based Augmentation [117] learns an augmentation schedule rather than a fixed policy as in AutoAugment to improve efficiency.
+These methods combine the benefits of both value-based transformations and structure-based transformations and further discover the potential of data augmentation with individual image data.
+Prompt-based Image Editing
+Visual generative models have been rapidly evolving in recent years.
+These models are typically conditioned on text embeddings, allowing flexible image editing by modifying textual descriptions of images and prompting the models.
+DA-Fusion [118] learns fine-grained concepts, replaces the original category labels, and uses the Stable Diffusion model to generate new images.
+SaSPA [119] uses GPT-4 to generate prompts based on the sample’s meta-class, then inputs the prompts, reference images, and detected edges to the Diffusion model to generate augmented images.
+IV-C2 Text Value-structure Manipulation
+Hierarchical Data Augmentation explicitly augments both value and structure information of text data.
+Back-translation Augmentation and Paraphrase Augmentation use neural networks to generate sentences that are semantic-preserving but different from original sentences in both wording and syntax.
+Hierarchical Data Augmentation
+Hierarchical Data Augmentation (HDA) [120] augments text data at both word and sentence levels.
+It evaluates the roles of words and sentences by training a hierarchical attention network to obtain the attention values of the text data at both levels.
+Attention values help decide the importance of components and augment the data by cropping and concatenating these components.
+Back-translation Augmentation
+Back-translation Augmentation.
+Back-translation [121] means translating text data from one language to another, or a sequence of other languages, and finally translating back to the original language.
+Semantic information is preserved, but the translation process can introduce some variance in syntax and wording.
+Such transformation can help the model learn more flexible expressions that convey the same meaning.
+Paraphrase Augmentation
+AugGPT [122], for example, prompts the state-of-the-art ChatGPT model to rewrite sentences in the dataset and preserves dataset coherence after data augmentation.
+Taboo [123] requires paraphrasing without using some frequently used words, Chain [124] paraphrase one sentence multiple times, and Hint [124] provide some examples, to encourage the LLM response otherwise.
+DoAug [125] constructs a preference dataset based on the diversity of paraphrased sentences and trains the LLM with the DPO to boost its generation diversity.
+V Multi-instance Level Data Augmentation
+Multi-instance augmentation combines value and structure information from different samples. Specific operation varies depending on the data modality in which it is applied.
+Multi-instance level image augmentation combines the visual information, i.e., pixels, from different images.
+It either mixes the colors of images or puts their patches together.
+Multi-instance level text augmentation integrates the linguistic information from different sentences.
+It either mixes the sentence at the word level or combines fragments from multiple sentences.
+Multi-instance level graph augmentation mixes the network information of different graph components or graph instances.
+It either mixes the attributes of different nodes or combines both the topology and attributes of graphs.
+Multi-instance level tabular augmentation combines the information of different lines or tables.
+Most methods mix the value representations across different lines, while it is also possible to combine two heterogeneous tables.
+Multi-instance time-series augmentation mixes values from different sequences, with few methods combining sequences using structural relationships.
+V-A Value-based Mixture
+V-A1 Pixel Mixture
+Image Mixup computes an arithmetical mixture of the color from different images and Neural Blending resorts to neural networks to enhance such process.
+Image Mixup
+Mixup [126] mixes two random images on the input layer by calculating their interpolation pixel-wise. Pixels from two original images each contribute a portion to the new image:
+$\vspace{-2mm}\mathbf{\tilde{x}}=\lambda\mathbf{x}_{i}+(1-\lambda)\mathbf{x}_{j},\qquad\mathbf{\tilde{y}}=\lambda\mathbf{y}_{i}+(1-\lambda)\mathbf{y}_{j}$ (14)
+where $(\mathbf{x}_{i},\mathbf{y}_{i})$ and $(\mathbf{x}_{j},\mathbf{y}_{j})$ are two labeled images from the dataset, and $\lambda\in[0,1]$.
+As a variant, manifold mixup [127] is operated on the hidden layers of the neural network.
+Most mixup methods combine two different images, while AugMix [128] applies three operations to one image, and combines the three augmented images to generate a new sample.
+In some approaches, the mixing operation varies in strength across regions.
+For example, SmoothMix [129] applies a soft-edged mask when mixing.
+Most mixup methods create new samples with soft labels, as in Eq.14, while there are some exceptions.
+Sample Paring [130], for example, uses the label of one of the original images, and that of the other image is left unused.
+ReMix [131] addresses the class imbalance problem by assigning the minority class label to augmented data.
+Neural Blending
+AutoMix [132] integrates a Mix Block in the training process.
+It builds a bridge between the selection of mixup policy and the optimization of the model.
+In AutoMix, the augmentation module is optimized along with the classification feature map, and the mixup functions and mixing ratio will updated according to the classification loss.
+Another image blending technique is proposed as Smart Augmentation [133].
+At the input end of Smart Augmentation, two random samples are fed into the network through two channels, and a CNN sub-network yields a one-channel output with the same shape as the input, which means one new image is generated based on the two input images.
+The augmented image is then used for downstream tasks, and the augmentation sub-network adjusts to the task performance accordingly.
+V-A2 Text Representation Mixture
+Unlike image data, whose coloring information is expressed by numerical values and can be interpolated at the input layer, words are difficult to interpolate directly, so text interpolation is usually done at the embedding layer.
+Text Mixup
+The adaptation of Mixup in sentence classification is studied in [134], which proposes wordMixup and senMixup.
+WordMixup interpolates a pair of word embeddings and then feeds the interpolation to the sentence encoder.
+SenMixup, on the other hand, interpolates a pair of sentence embeddings.
+The embedding vector produced by WordMixup or SenMixup is then passed to the softmax layer to produce the target distribution of the classification task.
+The use of mixup in text data augmentation has also been extensively studied in many research, such as nonlinear Mixup [135] and DoubleMix [136].
+GPT3Mix [137] makes use of the latest large language model to mix sentences.
+By inserting two source sentences in a prompt template that describes the instruction to the model, we can get a mixed sentence of high quality with a soft label.
+V-A3 Node Mixture
+Feature or label information from different nodes can be mixed by Graph Propagation, which is often used for tasks such as node classification.
+Graph Propagation
+Propagation refers to spreading information from one node to other nodes along the graph.
+These methods include feature propagation and label propagation.
+Feature propagation shares node features with other nodes.
+After feature vectors of some nodes are masked out, GRAND [50] further augments the node features by randomly propagating them through edges.
+MV-GCN [138] first creates three different views from the original graph, propagates the features in each, and then combines the three representations.
+Label propagation algorithm (LPA) [139] propagates labels from labeled nodes to all unlabeled nodes according to their proximity through areas with dense unlabeled nodes.
+AutoGRL [140] applies LPA in their semi-supervised framework to label more nodes utilizing a few ground truth nodes.
+[141] performs two types of label propagation, correcting the base predictions and smoothing the final prediction, respectively.
+Though not in a typical graph-based scenario, [142] constructs a nearest neighbor graph on the classifier’s manifolds.
+The affinity matrix, analogous to the adjacency matrix in this case, is constructed based on the $k$ nearest neighbors of each node’s embeddings, and this matrix decides how labels propagate.
+V-A4 Cell Value Mixture
+Due to its well-aligned structure, tabular data is very suitable for interpolation.
+Tabular Mixup
+SMOTE [37] is among the first works to study data interpolation, even before the proposal of Mixup [126].
 It addresses the problem of class imbalance by oversampling from the minority class.
 The samples are selected within the $k$ nearest neighbors in the feature space and new synthetic data are produced by a random interpolation operation.
-Since tabular features are often nominal, two variants, namely SMOTE-NC and SMOTE-N, are also proposed along with the initial version of SMOTE (Chawla et al., 2002).
-As argued in (Darabi and Elor, 2021), SMOTE is subject to unrealistic samples when interpolating in a sparse feature space.
-To solve this problem, TAEI (Darabi and Elor, 2021) first encodes the sparse data into a dense latent space, interpolates the new data, and then decodes them to obtain the original data plus the augmented data.
-Contrastive Mixup (Darabi et al., 2021) applies mixup for tabular data in a self- and semi-supervised learning framework.
-SAINT (Somepalli et al., 2021) combines CutMix (Yun et al., 2019) with contrastive learning to learn from tabular data.
-ExcelFormer (Chen et al., 2023) proposes two schemes, namely Feat-Mix and Hidden-Mix.
+Since tabular features are often nominal, two variants, namely SMOTE-NC and SMOTE-N, are also proposed along with the initial version of SMOTE [37].
+As argued in [143], SMOTE is subject to unrealistic samples when interpolating in a sparse feature space.
+To solve this problem, TAEI [143] first encodes the sparse data into a dense latent space, interpolates the new data, and then decodes them to obtain the original data plus the augmented data.
+Contrastive Mixup [144] applies mixup for tabular data in a self- and semi-supervised learning framework.
+SAINT [145] combines CutMix [146] with contrastive learning to learn from tabular data.
+ExcelFormer [147] proposes two schemes, namely Feat-Mix and Hidden-Mix.
 They interpolate data in the input layer and the hidden layer, respectively.
-7.2.2. Struture-based Tabular Combination
-Contrary to Individual Tabular Augmentation, which lacks structural information to use, Multiple Tabular Augmentation suffers from strict and fixed table structure.
+V-A5 Sequence Value Mixing
+Methods that combine information from different sequences are referred to as sequence averaging or sequence mixing.
+DTW Barycenter Averaging (DBA) [148] revolves around the averaging problem for Dynamic Time Warping.
+It develops a global technique for computing the average of a set of sequences and avoids using iterative pairwise averaging.
+In Weighted-DBA [149], instead of each time series contributing equally to the final average, some can contribute more than others, so that an infinite number of new examples from any set of given time series can be generated.
+[150] applies both interpolation and extrapolation in the feature space to augment time-series data.
+It first uses a sequence autoencoder to learn a feature space from unlabeled data.
+The data is then encoded and augmented with additive noises, interpolation, and extrapolation.
+Finally, the data in feature spaces is decoded and classified.
+Extending sequence mixing to the frequency domain, [151] applies the Equalized Mixture Data Augmentation (EMDA) to create new data.
+It computes the weighted average of two randomly chosen spectrograms with the same label.
+V-B Structure-based Combination
+V-B1 Image Structural Combination
+Image Patching puts together small fragments of many images or puts a part of an image over a second image.
+Image Patching
+RICAP [152] randomly crops four images, and patches them to construct a new image.
+CutMix [146] makes use of two images.
+It is developed based on Cutout [9], but the removed region is replaced by a patch from another image.
+Attentive CutMix [153] further enhances the CutMix method by adding a feature extraction module that selects an important or representative region, and pastes this attentive patch onto the other image.
+CowMask [154] introduces patching with irregular shape rather than rectangular.
+V-B2 Text Structural Combination
+Text Fragment Merging combines text data by merging text fragments from one sentence into another.
+Text Fragment Merging
+GECA [155] aims to remove a fragment from a sentence and populate it with another fragment that occurs in some common environment.
+Subtree Swapping [156], Substructure Substitution [157], and TreeMix [158] use tree structures to analyze the part of speech or role in the sentences and thus make reasonable swaps.
+SSMix [159] replaces a low saliency segment in a sentence with a segment with high saliency from another sentence and mixes sentence labels according to the ratio of their span length in the combined sentence.
+V-B3 Table Concatenation
+Contrary to single-instance level tabular augmentation, which lacks structural information to use, multi-instance level tabular augmentation suffers from strict and fixed table structures.
 Columns differ in their nomenclatures and representations, so it is hard to combine and learn from multiple distinct tables.
 If these restrictions are relaxed, models can learn across different tables, be more transferable, and better predict unseen tables.
-Structure Relaxation.
-TransTab (Wang and Sun, 2022) suggests that multiple tables could share partially overlapped columns in the real world, but traditional methods often fail to fully utilize the data due to the removal of non-overlapping columns and mismatched samples when cleaning the data.
+Structure Relaxation
+TransTab [160] suggests that multiple tables could share partially overlapped columns in the real world, but traditional methods often fail to fully utilize the data due to the removal of non-overlapping columns and mismatched samples when cleaning the data.
 TransTab first divides the table into three parts, each containing categorical, binary, and numerical data.
@@
 The augmented table contains information from many original tables and helps the model learn across them.
-7.3. Populational Tabular Augmentation
-7.3.1. Vanilla Tabular Generation
+V-C Value-structure Mixture
+V-C1 Image Value-structure Mixture
+Diffusion Models take textual prompts as inputs and can perform a flexible and smart mixture of existing images.
+Prompt-based Mixing
+Conditioned on textual inputs, Diffusion Models can generate combined images by using textual descriptions of multiple images.
+Diff-Mix [161] uses text embeddings of different categories to replace the foreground to align with the target classes.
+Diff-II [162] calculates the inversion noise of two images, conducts random circle interpolation, and then denoises with different prompts.
+V-C2 Graph Value-structure Mixutre
+Graph Mixup refers to mixing graph nodes or graph instances. This often involves both node attribute values and the topology structures.
+Graph Mixup
+Due to the irregular connectivity and various topological structures of graph data, it is hard to directly combine nodes or graphs.
+For node-level mixup, one solution to this problem is temporarily ignoring the edges and mixing node features in latent space.
+GraphMix [163] interpolates the hidden states and labels of input nodes.
+A topology-involved node mixup method is proposed in [164].
+It randomly pairs nodes and includes the nodes’ local topology representations when interpolating them.
+S-Mixup (Structural Mixup) [165] considers neighboring edges when pairing nodes but works in the input space and discusses how to connect the node with the original graph with edges.
+It matches nodes based on GNN classification confidence and connects with high-gradient edges to the mixed node.
+For graph-level mixup, ifMixup [166] adds dummy nodes so that two graphs have the same number of nodes.
+These dummy nodes have all-zero feature vectors and are not connected to any other nodes, or in other words, the weights of edges on dummy nodes are zeros.
+S-Mixup (Graph Mixup with Soft Alignments) [167] also aligns graphs first before mixing them.
+It obtains an assignment matrix by a soft assignment and uses the assignment matrix to transform a graph so that every node corresponds with a node in the other graph.
+Compared with the above approaches, G-Mixup [168] proposes a more sophisticated way to interpolate graph data.
+It first uses graphs within the same class to estimate a graphon, i.e.
+graph generator:
+$\vspace{-2mm}\mathbf{\mathcal{G}}\rightarrow W_{\mathbf{\mathcal{G}}},\mathbf{\mathcal{H}}\rightarrow W_{\mathbf{\mathcal{H}}}$ (15)
+The graphon $W:[0,1]^{2}\rightarrow[0,1]$ is a matrix, where $W(i,j)$ describes the probability of an edge between nodes $i$ and $j$, and graphs of the same class are considered to be generated from the same graphon.
+Graphons of two classes of graphs are then interpolated to produce a mixed graphon:
+$\vspace{-2mm}W_{\mathbf{\mathcal{I}}}=\lambda W_{\mathbf{\mathcal{G}}}+(1-\lambda)W_{\mathbf{\mathcal{H}}}$ (16)
+By sampling from the graphon mixup $W_{\mathbf{\mathcal{I}}}$, we acquire a series of graphs with the same soft labels:
+$\vspace{-2mm}\mathbf{y}_{\mathbf{\mathcal{I}}}=\lambda\mathbf{y}_{\mathbf{\mathcal{G}}}+(1-\lambda)\mathbf{y}_{\mathbf{\mathcal{H}}}$ (17)
+During graphon estimation, G-Mixup aligns the original node features and obtains the graphon node features average pooling, which are used as node features of the generated graphs.
+DAGAD [169] mixes graph representations by performing random permutations, demonstrating that data augmentation is helpful in scenarios such as graph anomaly detection.
+VI Dataset Level Data Augmentation
+Dataset-level data augmentation methods learn the overall data distribution and often generate new data that is consistent but diverse in terms of both value and structure information.
+VI-A Vanilla Generation
+A straightforward way to produce augmented data from the entire dataset is to learn the data distribution of existing datasets with neural networks and generate new samples that fall in that distribution.
+Some typical solutions are Autoencoder (AE), Generative Adversarial Networks (GAN), Large Language Models (LLMs), and Generative Diffusion Models.
+VI-A1 Autoencoder
 An Autoencoder (AE) learns from the dataset and extracts the essential structure, which can be used to generate a more robust dataset.
-Generative Adversarial Network (GAN) is another powerful method that can produce more augmented data and is especially useful for oversampling and alleviating the problem of class imbalance.
-Autoencoder.
-(Delgado and Oyedele, 2021) tries out four types of autoencoder, concluding that all of these autoencoders can generate data sets that produce better task performance, while the variational autoencoder (VAE) provides the most robust result.
-Some other works (Fang et al., 2022; Tran et al., 2019) also use VAE as their data generation approach.
-SDAT (Fang et al., 2022) adds extra noise to the latent space of its VAE to obtain augmented samples.
+It is widely used for tabular data augmentation.
+[170] tries out four types of autoencoder, concluding that all of these autoencoders can generate data sets that produce better task performance, while the variational autoencoder (VAE) provides the most robust result.
+Some other works [171, 172] also use VAE as their data generation approach.
+SDAT [171] adds extra noise to the latent space of its VAE to obtain augmented samples.
 These samples, along with the original ones, are then trained under a semi-supervised framework.
-(Tran et al., 2019) combines VAE with active learning.
+[172] combines VAE with active learning.
 The samples are iteratively selected by active learning, labeled by an oracle, reconstructed by the VAE, and discriminated and classified.
-Generative Adversarial Networks.
-Table-GAN (Park et al., 2018) mainly considers the privacy issue, and experiments show that the augmented table given by GAN is compatible with downstream machine learning models and produces similar performance.
-CTGAN (Xu et al., 2019) focuses on the class imbalance problem and proposes a conditional generator that accounts for the imbalance in categorical columns.
-(Engelmann and Lessmann, 2021) shares similar goal and method with CTGAN (Xu et al., 2019) but employs a more sophisticated loss function that better incorporates the discriminator and the auxiliary classifier.
-ITS-GAN (Chen et al., 2019) proposes a framework that aims to preserve functional dependencies in tables.
+VI-A2 Generative Adversarial Networks
+The rise of Generative Adversarial Networks (GANs) [173] makes it possible for humans to generate new images on a large scale using computers.
+This nature also leads to a new path of data augmentation: creating new data samples instead of modifying or combining old ones.
+It can produce considerably more augmented data and is especially useful for oversampling and alleviating the problem of class imbalance.
+For image data, GANs have now been widely used as a data augmentation technique [12], especially in the medical field [174, 175], where data is to some extent scarce.
+Apart from simply enlarging the dataset, GAN can be used in other ways such as removing background noise and cleaning the dataset [176, 177, 178].
+It can also generate different views or poses of the object based on the original image [179], and thus add more variation to the dataset.
+Tabular data augmentation is another case where GANs play an important role.
+Table-GAN [180] mainly considers the privacy issue, and experiments show that the augmented table given by GAN is compatible with downstream machine learning models and produces similar performance.
+CTGAN [181] focuses on the class imbalance problem and proposes a conditional generator that accounts for the imbalance in categorical columns.
+[182] shares a similar goal and method with CTGAN [181] but employs a more sophisticated loss function that better incorporates the discriminator and the auxiliary classifier.
+ITS-GAN [183] proposes a framework that aims to preserve functional dependencies in tables.
 It models functional dependencies with autoencoders and includes a term to characterize the set of functional dependency constraints in the generator loss function.
-Arguing that GANs are difficult to train on small datasets, ConvGeN (Schultz et al., 2023) designed a new mechanism that requires the GAN’s generator and discriminator to cooperate with each other, that the generator generates batches of convex combination of minority neighborhood samples and the discriminator classifies these augmented minority samples against majority batches.
-7.3.2. Exogenous Tabular Generation
-The potential of tabular data is limited by its simple nature compared with data from other modalities.
+Arguing that GANs are difficult to train on small datasets, ConvGeN [184] designed a new mechanism that requires the GAN’s generator and discriminator to cooperate, that the generator generates batches of a convex combination of minority neighborhood samples and the discriminator classifies these augmented minority samples against majority batches.
+GANs are also widely used to generate virtual time-series data.
+[185] uses an RGAN and an RCGAN to produce realistic medical time-series data.
+TimeGAN [186] combines supervised and unsupervised losses to train its autoencoding components and adversarial components.
+[187] propose WaveGAN and SpecGAN to generate time-series data.
+WaveGAN is used on the waveform, while SpecGAN is used on the spectrogram.
+SpecGAN demonstrates how GAN is applied to generate time-series data on the frequency domain.
+VI-A3 Large Language Models
+Recent advances in large language models (LLM) and generative pre-training (GPT) [188] models have demonstrated unprecedented power in text understanding, classification, and generation tasks.
+Apart from supporting intelligent chat-bots that help people in their daily lives, the capacity of these state-of-the-art language generation models establishes a new paradigm of text data augmentation.
+LAMBADA [189] is among the first attempts to use a state-of-the-art language generator to synthesize new text data for data augmentation, and its procedure is as follows.
+LAMBADA first trains a baseline classifier from a small dataset, and then it fine-tunes the GPT-2 [190] model.
+A set of labels is thereafter fed to the fine-tuned language model, which generates a large dataset of labeled sentences.
+These sentences are evaluated by the classifier trained prior according to its confidence score, and top-ranked sentences are retained to compose the synthesized dataset.
+Similarly, [191] integrates GPT-2 in its text data augmentation process and improves the performance of active learning on tiny datasets.
+It uses the language model to generate a tree of tokens as candidates and construct a sentence from them.
+The set of new sentences is repeatedly refined by active learning, and the outcome is an augmented text dataset.
+Some other research also applies GPT-2 to effectively augment low-resource text datasets, such as [192] and [193].
+Despite their success, both of them point out the need for more sophisticated methods to generate text data.
+[192] for example, argues that human-in-the-loop for relabeling might be rewarding.
+[193] adopts human-in-the-loop for seed selection and further discusses the potential benefits of injecting expert knowledge for specialized domains.
+Notably, Data Boost [194] proposes a reinforcement learning framework to guide the process of text data augmentation with GPT-2 and succeeds in generating data of high quality and high similarity to the original data.
+The rise of GPT-3 [195] and later LLMs is now providing more possibilities for text generative augmentation. Their use and improvements are studied in the latest research such as [196] and [197].
+VI-A4 Generative Diffusion Models
+Despite that AEs and GANs can all be used to generate graph data, they face many problems in practice.
+VAE models require things like a massive graph-matching process, and GANs on graphs easily fall into mode collapse [198].
+So, graph data augmentation resorts to another recently proposed technique, namely, Generative Diffusion Models.
+The denoising diffusion model [199] is a generative paradigm that has become popular in recent years.
+Though not yet widely applied as an approach to graph data augmentation, the Generative Diffusion Model boasts enormous potential in graph generation and can possibly be adopted as a promising data augmentation technique.
+To implement diffusion models on graph data, some properties of the graph data must be properly treated.
+For example, diffusion models often rely on a continuous Gaussian noise process, but the graph structures are discrete [200].
+DiGress [201] designs a discrete denoising diffusion model for graph generation and achieves state-of-the-art performance on molecular and non-molecular datasets.
+The use of Generative Diffusion Models for data augmentation especially benefits studies on molecules, proteins, and materials, whose data are often in the form of graph instances.
+VI-B Exogenous Generation
+Apart from the source dataset, we can also refer to exogenous information for data augmentation.
+Exogenous information is perceived by introducing external datasets or using knowledge learned by human experts.
+VI-B1 Introducing Exogenous Datasets
+Neural Style Transfer
+Image datasets vary in artistic styles, and neural style transfer [202] introduces style information from an external dataset when augmenting an entire dataset.
+It learns the artistic style from a group of image data and uses this artistic style to depict the topic of images from another dataset.
+As is [202], neural style transfer can be used to mimic the drawing styles of particular artists.
+Such an augmentation process can help the model learn a more general representation of the topics.
+[203] proposes a more realistic use of neural style transfer.
+It transfers the weather or illumination from reference images to the target dataset.
+This is particularly useful because it is hard to collect sufficient images under certain weather.
+Table Retrieval
+Tabular data populates the web and by retrieving consistent and related tables, it is possible to generate an augmented table that incorporates extended information.
+A key consideration of Table Retrieval Augmentation is to design appropriate ways to search for ideal tables.
+Santos [204] leverages external knowledge bases, YAGO 4, and focuses on designing a union search technique to match desired target tables.
+HYTREL [205] predicts table similarities by building hypergraphs on top of the tables and uses a hypergraph encoder to calculate the encoding.
+LEKA [206] uses an LLM to extract the table’s descriptive information, retrieve it in the Kaggle database, and harmonize tables from different sources.
+VI-B2 Leveraging Expert Knowledge
+Here we introduce three examples of how humans utilize expert knowledge for image, tabular, and time-series data augmentation
+Computer Graphic Modeling is to leverage expert knowledge on how objects in the existing dataset look like and the power of computer software to generate new image samples [207, 208, 209].
+These datasets have several advantages over ordinary image datasets.
+First, it requires little human effort to collect data.
+Second, by applying different aspects, poses, and lighting conditions, computer graphics modeling can generate images with high diversity.
+Third, these images are all fully labeled with accurate ground truth.
+Relational Structure Construction can enhance the potential of tabular data, which is limited by its simple nature compared with data from other modalities.
 Due to the lack of spatial or relational relationships, many popular augmentation approaches and network configurations are not suitable for tabular datasets.
-A solution is to construct and introduce relational structure based on the tabular dataset.
-Relational Structure Construction.
-GOGGLE (Liu et al., 2022c) proposes a solution to this problem.
+A solution is to construct and introduce a relational structure based on the tabular dataset.
+GOGGLE [210] proposes a solution to this problem.
 It learns the relational structure and corresponding functional relationships from tabular datasets and uses this basis to generate new samples.
 In other words, it learns a graph from the features in the table and obtains augmented data from that graph.
-PET (Du et al., 2022) depicts multiple tabular data instances are as a hyper graph and use the labels to help construct hyper edges.
-These methods are based on external assumptions and introduces extra structural information beyond the tabular dataset.
-Table 8. A summary for tabular data augmentation
-Tier 1: Quantity of source data Tier 2: Information type Augmentation method Typical literature
-Individual Tabular Augmentation Value-based Transformation Table Masking VIME (Yoon et al., 2020) (NeurIPS 2020), SCARF (Bahri et al., 2021) (ICLR 2021), MET (Majmundar et al., 2022) (arXiv 2022), MTR (Onishi and Meguro, 2023) (arXiv 2023)
-Feature Engineering RL1 (Khurana et al., 2018) (AAAI 2018), MARFS (Liu et al., 2019) (KDD 2019), AutoFS (Fan et al., 2020) (ICDM 2020), MCRFS (Liu et al., 2021b) (ICDM 2021), RAFT (Xiao et al., 2023a) (SDM 2023), GRFG (Wang et al., 2022) (KDD 2022), GAINS (Xiao et al., 2023b) (ICDM 2023), MOAT (Wang et al., 2023) (NeurIPS 2023), TFWT (Zhang et al., 2024) (IJCAI 2024)
-Structure-based Transformation Table Subsetting SubTab (Ucar et al., 2021) (NeurIPS 2021)
-Multiple Tabular Augmentation Value-based Mixture Tabular Mixup SMOTE (Chawla et al., 2002) (JAIR 2002), TAEI (Darabi and Elor, 2021) (arXiv 2021), Contrastive Mixup (Darabi et al., 2021) (arXiv 2021), SAINT (Somepalli et al., 2021) (arXiv 2021), ExcelFormer (Chen et al., 2023) (arXiv 2023)
-Structure-based Combination Structure Relaxation TransTab (Wang and Sun, 2022) (NeurIPS 2022)
-Populational Tabular Augmentation Vanilla Generation Autoencoder (Delgado and Oyedele, 2021) (Applied Soft Computing 2021), SDAT (Fang et al., 2022) (CIKM 2022), (Tran et al., 2019) (ICML 2019)
-Generative Adversarial Networks Table-GAN (Park et al., 2018) (arXiv 2018), CTGAN (Xu et al., 2019) (NeurIPS 2019), (Engelmann and Lessmann, 2021) (Expert Systems with Applications 2021), ITS-GAN (Chen et al., 2019) (IJCAI 2019), ConvGeN (Schultz et al., 2023) (Pattern Recognition 2023)
-Exogenous Generation Relational Structure Construction GOGGLE (Liu et al., 2022c) (ICLR 2022), PET (Du et al., 2022) (NeurIPS 2022)
-8. Data Augmentation for Time-series Data
-Time-series data represent sequential values within a period, and each timestamp is considered as an elemental component.
-Data augmentation methods for time-series data either perturb the value at each timestamp or perturb the sequence of timestamps.
-8.1. Individual Time-series Augmentation
-8.1.1. Value-based Time-series Transformation
-A type of simple and straightforward value-based method is Value Perturbation, which changes the value represented by a sequence.
-There is also another type of method called Decomposition, which decomposes the value of a sequence into different parts that express the sequential values from different aspects.
-Value Perturbation. Value Perturbation augments time-series data by altering the value on the Y-axis. It often perturbs the amplitude at a certain timestamp or in a certain frequency channel.
-One of the most simple but popular augmentation methods is to add random noise to the value at each timestamp.
-For time-series data, this is referred to as jittering.
-In (Um et al., 2017), jittering is used to simulated additive noise $\epsilon$ of wearable sensors:
-(12) $\tilde{x}(\epsilon)=\{x_{1}+\epsilon_{1},\cdots,x_{t}+\epsilon_{t},\cdots,x_{T}+\epsilon_{T}\}$
-Scaling (Um et al., 2017) alters the data by multiplying the magnitude of the entire sequence with a scalar $\alpha$.
-(13) $\tilde{x}(\alpha)=\{\alpha x_{1},\cdots,\alpha x_{t},\cdots,\alpha x_{T}\}$
-Magnitude warping (Um et al., 2017) is similar to scaling, but it uses a smooth curve to warp the sequence so that magnitude at different timestamps is scaled with different parameters.
-(14) $\tilde{x}(\alpha)=\{\alpha_{1}x_{1},\cdots,\alpha_{t}x_{t},\cdots,\alpha_{T}x_{T}\}$
-Rotation (Um et al., 2017), or flipping, assumes the data is symmetric with respect to the X-axis, and makes the data upside down by applying a rotation matrix $R$.
-(15) $\tilde{x}(R)=\{Rx_{1},\cdots,Rx_{t},\cdots,Rx_{T}\}$
-Apart from the amplitude, the phase spectrum provides another option for data augmentation.
-RobustTAD (Gao et al., 2020) converts the time-series data from the time domain to the frequency domain, as shown in Eq.
-16, and then increases all phases $\theta(\omega_{k})$ by a small perturbation.
-(16) $F(\omega_{k})=\frac{1}{N}\sum^{N-1}_{t=0}{x_{t}e^{-j\omega_{k}t}}=A(\omega_{k})exp[j\theta(\omega_{k})]$
-Decomposition.
-Empirical Mode Decomposition (EMD) (Huang et al., 1998) is used to produce a set of intrinsic mode functions (IMFs).
-This decomposition relies on extracting the energy linked to different time scales that are inherent to the data.
-EMD has been utilized as a method of data augmentation in (Nam et al., 2020) for the purpose of categorizing impact noise in vehicles.
-Another way to decompose time-series data is STL (Cleveland et al., 1990), which decomposes signals into seasonal, trend, and remainder:
-(17) $y_{t}=\tau_{t}+s_{t}+r_{t},\quad t=1,2,\cdots N$
-Other works such as (Wen et al., 2019) and (Gao et al., 2020) focus on improving robustness with the help of STL components.
-8.1.2. Structure-based Time-series Transformation
-Unlike Value Perturbation, sequence perturbation is implemented on the X-axis, either on the time domain or the frequency domain.
-Window Slicing splits the time-series data into pieces.
-It is also the basis of many Sequence Morphing methods because they apply varied operations on these slices.
-Window Slicing.
-Inspired by image cropping, which is widely used for image data augmentation, MCNN (Cui et al., 2016) proposes window slicing.
-It extracts slices from the entire sequence as new training samples.
-Each slice is assigned with the label of the original sequence.
-Sequence Morphing.
-Based on Window Slicing, (Le Guennec et al., 2016) extends or compresses the generated spans, which is called window wrapping.
-This is equivalent to speeding up or slowing down a span of the sequence.
-Vowel stretching (Nagano et al., 2019) is a special kind of window wrapping technique.
-It only perturbs the sequence by prolonging the vowels.
-This is meant to be consistent with how children speak.
-(Um et al., 2017) propose to randomly permeate these slices and concatenate them as a long sequence.
-Similar processing is also called Random Shuffling and is used in (Steven Eyobu and Han, 2018).
-VTLP (Jaitly and Hinton, 2013) applies window wrapping in the frequency domain.
-It adds a small distortion to the central frequency.
-SpecAugment (Park et al., 2019) also proposes to augment the frequency domain.
-It generates masks of length sampled from a uniform distribution and uses each of them to mask consecutive frequency channels.
-8.2. Multiple Time-series Augmentation
-8.2.1. Value-based Sequence Mixing
-Most methods that combine information from different sequences only use the values, and they are referred to as sequence averaging or sequence mixing.
-For example, DTW Barycenter Averaging (DBA) (Petitjean et al., 2011) revolves around the averaging problem for Dynamic Time Warping.
-It develops a global technique for computing the average of a set of sequences and avoids using iterative pairwise averaging.
-In Weighted-DBA (Forestier et al., 2017), instead of each time series contributing equally to the final average, some can contribute more than others, so that an infinite number of new examples from any set of given time series can be generated.
-(DeVries and Taylor, 2017a) applies both interpolation and extrapolation in the feature space to augment time-series data.
-It first uses a sequence autoencoder to learns a feature space from unlabeled data.
-The data is then encoded and augmented with additive noises, interpolation, and extrapolation.
-Finally, the data in feature spaces is decoded and classified.
-Extending sequence mixing to the frequency domain, (Nanni et al., 2020) applies the Equalized Mixture Data Augmentation (EMDA) to create new data.
-It computes the weighted average of two randomly chosen spectrograms with the same label.
-8.3. Populational Time-series Augmentation
-8.3.1. Vanilla Time-series Generation
-GANs are widely used to generate virtual time-series data.
-Generative Adversarial Networks.
-(Esteban et al., 2017) uses an RGAN and an RCGAN to produce realistic medical time-series data.
-TimeGAN (Yoon et al., 2019) combines supervised and unsupervised losses to train its autoencoding components and adversarial components.
-(Donahue et al., 2018) propose WaveGAN and SpecGAN to generate time-series data.
-WaveGAN is used on waveform, while SpecGAN is used on spectrogram.
-SpecGAN demonstrates how GAN is applied to generate time-series data on the frequency domain
-8.3.2. Exogenous Time-series Generation
+PET [211] depicts multiple tabular data instances as a hyper-graph and uses the labels to help construct hyper-edges.
+These methods are based on external assumptions and introduce extra structural information beyond the tabular dataset.
+Statistical Generation is a technique that predicts values in the future.
 Some methods view time-series data outside the existing data and from a statistical perspective.
-They are based on human knowledge on statistics and can predict how the value changes in future timestamps.
-Statistical Generation.
-(Smyl and Kuber, 2016) uses a statistical algorithm called LGT (Local and Global Trend) to forecast paths of the time-series data and samples from them to get extra data.
-GRATIS (Kang et al., 2020) employs mixture autoregressive (MAR) models to generate collections of time series and examine the variety and extent of the produced time series within a feature space for time series.
-Table 9. A summary for time-series data augmentation
-Tier 1: Quantity of source data Tier 2: Information type Augmentation method Typical literature
-Individual Time-series Data Augmentation Value-based Transformation Value Perturbation Jittering, Scaling, Magnitude Wrapping, and Rotation (Um et al., 2017) (ICMI 2017), Phase Perturbation (Gao et al., 2020) (arXiv 2020)
-Decomposition EMD (Nam et al., 2020) (ICASSP 2020), STL (Cleveland et al., 1990) (J. Off. Stat 1990), RobustSTL (Wen et al., 2019) (AAAI 2019), RobustTAD (Gao et al., 2020) (arXiv 2020)
-Structure-based Transformation Window Slicing Window Slicing (Cui et al., 2016) (arXiv 2016)
-Sequence Morphing Window Wrapping (Le Guennec et al., 2016) (ECML-PKDD 2016), Vowel Stretching (Nagano et al., 2019) (ICASSP 2019), Permutation (Um et al., 2017) (ICMI 2017), Random Shuffling (Steven Eyobu and Han, 2018) (Sensors 2018), VTLP (Jaitly and Hinton, 2013) (ICML 2013), SpecAugment (Park et al., 2019) (arXiv 2019)
-Multiple Time-series Data Augmentation Value-based Mixture Sequence Mixing DTW Barycenter Averaging (Petitjean et al., 2011) (Pattern Recognition 2011), Weighted-DBA (Forestier et al., 2017) (ICDM 2017), (DeVries and Taylor, 2017a) (arXiv 2017), EMDA (Nanni et al., 2020) (Ecological Informatics 2020)
-Populational Time-series Data Augmentation Vanilla Generation GAN (Esteban et al., 2017) (arXiv 2017), TimeGAN (Yoon et al., 2019) (NeurIPS 2019), WaveGAN & SpecGAN (Donahue et al., 2018) (ICLR 2018)
-Exogenous Generation Statistical Generation Local and Global Trend (Smyl and Kuber, 2016) (ISF 2016), MAR (Kang et al., 2020) (Statistical Analysis and Data Mining 2020)
-9. Discussion
-Influence of Data Augmentation Operations.
-Some data augmentation methods aim to produce noisy or even “wrong” data to improve the AI models’ robustness by learning those low-quality or adversarial samples; in contrast, some are designed to improve data quality by removing noise or error from the data, so that the models can learn a clean representation of the data; some other methods do not intend to affect data quality, and they only make some semantic-preserving perturbation to the data, enlarge the dataset size, and cover more area in the feature space.
-Evaluation Metrics Specific to Data Augmentation.
-Most research on data augmentation uses the models’ performance on downstream tasks to evaluate the effect of data augmentation methods.
-However, this criterion depends on too many extraneous factors and may fail to reflect the influence of data augmentation methods on the data.
-So, there is need for a set of evaluation metrics that can assess specific considerations of data augmentation, such as how much diversity is introduced and data distribution in the feature space changes.
-Harnessing Latest Techniques.
-Large language models possess outstanding ability in manipulating text data, but up to now, most text data augmentation methods relies on symbolic analysis and rules.
-The question of how to utilize advanced techniques, such as LLM, to perform data augmentation tasks remains open and is rewarding to discover.
-10. Conclusion
-This survey presents a comprehensive summary of data augmentation techniques across five data modalities and proposes a modality-independent taxonomy from a data-centric perspective, which focuses on where the augmented data is derived from.
+They are based on human knowledge of statistics and can predict how the value changes in future timestamps.
+[212] uses a statistical algorithm called LGT (Local and Global Trend) to forecast paths of the time-series data and samples from them to get extra data.
+GRATIS [213] employs mixture auto-regressive (MAR) models to generate collections of time series and examine the variety and extent of the produced time series within a feature space for time series.
+VII Discussions
+Up to now this survey focuses on wherefrom the information in the augmented samples is derived. And there are many other perspectives to discuss data augmentation and understand how it works.
+VII-A Augmentation Evaluation Metrics
+Most research on data augmentation uses the models’ performance on downstream tasks to evaluate the effect of data augmentation [13].
+However, this criterion depends on many extraneous factors and may fail to reflect the influence of data augmentation methods on the data.
+So, there is a need for a set of evaluation metrics that can assess specific considerations of data augmentation.
+Two of the most critical metrics are affinity and diversity [161, 162, 214, 215, 216, 217, 125].
+Affinity is also referred to as consistency, correctness, faithfulness, quality, similarity, or validity.
+It reflects if the augmented sample is semantically consistent with the original sample and preserves the original label.
+Model-centric evaluation [215] defines affinity as the ratio of model $m$’s performance on augmented validation dataset $D^{\prime}_{val}$ and that on the original validation dataset $D_{val}$, given by
+[…]
+where $\text{ACC}(m,D)$ stands for model $m$’s performance on evaluation dataset $D$.
+Data-centric evaluation, on the other hand, defines affinity as how much the augmented dataset deviates from the original dataset in the feature space.
+For example, [214] proposes Fréchet BERT Distance (FBD) to evaluate the deviation of two textual datasets, defined as:
+[…]
+where $m$ is the mean vector and $C$ is the covariance matrix.
+Diversity reflects the breadth of patterns in the dataset and exposing the model to a diverse training dataset can improve its robustness.
+[218] considers a dataset as an ellipsoid and defines its diversity as the isocontour radius of the ellipsoid:
+[…]
+where $H$ is the dimension of the feature vectors and $\sigma_{i}$ is the standard deviation along the $i$-th feature. [219] proposes to evaluate dataset diversity by calculating the dispersion:
+[…]
+where $d(x_{i},x_{j})=1-\frac{f(x_{i}){\cdot}f(x_{j})}{\|f(x_{i})\|\|f(x_{j})\|}$.
+Some empirical studies including [219], [217], and [215] investigate how affinity and/or diversity influence the effectiveness of data augmentation, and their real-world experimental results demonstrate the model performance takes the most advantage of data augmentation that achieves both high affinity and diversity.
+VII-B Selecting the Target
+VII-B1 On Selecting Pixels
+For Pixel Erasing, selecting target pixels is a necessity because it is impossible to mask out all pixels in the image.
+If erasing is performed by randomly selecting an area in the image, it is called Image-aware Random Erasing (IRE) [10].
+In some tasks, the location of the target in the image is known, so, random erasing can be restricted in the object’s bounding box, known as Object-aware random erasing (ORE) [10].
+Further argued in [220], the erased area within the bounding box should be limited to a manageable size and a part of the less informative background should also be erased.
+[220] proposed Region-aware Random Erasing (RRE), which introduces ORE with upper bounds for width and height in the bounding boxes, and then randomly erases the background.
+The same consideration occurs when performing Image Cropping and Mixup, and some importance-based strategies are used to preserve important contents in the images.
+For Image Cropping (IC), [74] uses entropy as the importance score to select the most importance regions, and [75] uses eye-tracking technique to locate important content.
+For Image Mixup, Co-mixup [221] and Puzzle mix [222] both use saliency to decide the mixing strength of different regions, potentially leading to better mixing performances.
+A summary of how these methods select target pixels are presented in Table III.
+TABLE III: Region selection strategy for Pixel Erasing, Image Cropping and Mixing methods.
+Name Ref Augmentation Strategy
+Random Location-based Importance-based
+Cutout [9]
+Image-aware RE [10]
+Object-aware RE [10]
+Region-aware RE [220]
+Hide-and-seek [38]
+Grid-Mask [39]
+Entropy-bases IC [74]
+Gaze-based IC [75]
+Co-mixup [221]
+Puzzle mix [222]
+VII-B2 On Selecting Words
+Most value-based text transformation methods randomly select the word to be replaced, added, and deleted, but some methods considers the role a word plays in the sentence when deciding whether to impose perturbation.
+For example, TEXTFOOLER [223] ranks words by their importance and replaces the important words with their synonyms.
+Syntax-aware data augmentation [48] argues that most previous text data augmentation approaches are only computation-motivated and cannot assure the good quality of the generated text.
+Instead of simply modifying words with the same probability, Syntax-aware data augmentation calculates the probability of being modified for each word based on its syntactic role in the sentence.
+Specifically, it utilizes a dependency tree to calculate the importance of each word, i.e.
+the depth in the tree.
+A word far from the root is considered less important syntactically and is more likely to be removed (by means of dropout and blanking) or replaced during the augmentation process.
+Instead of using a syntax tree to assess word importance, Selective Text Augmentation (STA) [224] focuses more on the statistical and semantic features of words to describe the role of words.
+It proposes two perspectives to view words in text classification tasks, namely, statistical correlation, which means how frequently a word co-occurs with a certain category, and semantic similarity, which measures the semantic overlap of the given words and the category label.
+Using the metrics as axes, STA divides the roles of words into four quadrants: gold, venture, bonus, and trivial.
+Words with different roles are treated differently when imposing augmentation operations on the sentence.
+For example, gold words will be kept in selective deletion and synonyms of gold words will be added to the sentences in selective insertion.
+A summary of how these methods select target words is presented in Table IV.
+TABLE IV: Word selection strategy of text data selection methods.
+Name Ref Augmentation Strategy
+Random Role-based Learned
+EDA [11]
+- [43]
+TEXTFOOLER [223]
+LetterTran [225]
+AEDA [46]
+INSNET [226]
+LeCA [227]
+GBS [228]
+- [47]
+STA [224]
+Subtree Swapping [156]
+Substructure Substitution [157]
+TreeMix [158]
+SSMix [159]
+VII-B3 On Selecting Nodes and Edges
+For graph data augmentation, a key problem is selecting target nodes and edges to operate on.
+Node centrality acts as an important factor that is both meaningful and convenient, supporting selection strategies of many graph data augmentation methods.
+GCA [229] uses node centrality to calculate the masking vectors and rewiring probabilities and thus decide which node to mask and which edge to rewire.
+GrpahCrop [93] applies a node-centric strategy to preserve connectivity in subgraphs.
+SUBG-CON [94] uses the Personalized PageRank centrality to select the important nodes for subgraphing.
+Smart rewiring [103] takes the degrees of chosen nodes when swapping their edges.
+Another useful factor is node similarity.
+DHGR [106] learns a similarity matrix from the original graph and rewires the graph by adding edges between high-similarity node pairs and removing edges from low-similarity ones.
+Similarly, RobustECD [230] and M-Evolve [231] also consider node similarity for edge rewiring.
+TABLE V: Node and edge selection strategy for grpah data augmentation methods.
+Name Ref Augmentation Strategy
+Random Centrality Similarity
+Attribute masking [49]
+GRAND [50]
+BGRL [51]
+GCA [229]
+GrpahCrop [93]
+SUBG-CON [94]
+Smart rewiring [103]
+DHGR [106]
+RobustECD [230]
+M-Evolve [231]
+VII-C Leveraging Augmented Samples
+A basic use of data augmentation is to enlarge the dataset, especially in data-scarce scenarios, where the augmented samples are appended to the training dataset.
+Beyond this paradigm, data augmentation provides many other possibilities to support deep learning.
+A frequent usage of augmented samples is to construct different views for contrastive learning.
+For example, Rock-ViT [232] performs rotation, flipping, blurring, contrast enhancement, irregular noise, and microscope dark-field noise on sandstone microscopic images.
+It optimizes the vision transformer with both cross-entropy loss and contrastive loss.
+When calculating the contrastive loss, Rock-ViT uses two random augmentations to build positive pairs.
+Instead of simply calculating contrastive loss between positive and negative sample pairs, DualCL [233] augments text samples by including the label in the sentence.
+Specifically, for an original sentence “[CLS] love this movie” in a sentiment classification dataset, DualCL converts it to “[CLS] POS NEG love this movie”, where “POS” and “NEG” denote the positive and negative labels.
+Then it repels the embeddings of the “[CLS]” token and the “NEG” token, since this sample is possible and thus “[CLS] should also be positive.
+Conversely, it attracts the embeddings of the “[CLS]” and “POS” tokens.
+Contrastive learning is a common practice in graph deep learning, and graph data augmentation plays a critical role in many graph deep learning pipelines.
+GraphCL[49] uses Node Dropping and Edge Perturbation to build different views, and MVGRL [99] applies graph diffusion.
+Different views of an original graph are used as positive pairs and a contrastive loss function is used to enforce maximizing their consistency.
+VIME [56] expend the usage of data augmentation beyond contrastive learning.
+It incorporates tabular data augmentation in both self-supervised and semi-supervised learning frameworks.
+For self-supervised learning, it corrupts the table with a mask and optimizes the model with table reconstruction loss and mask estimation loss.
+For semi-supervised learning, it calculates the consistency loss between masked views.
+VII-D Harnessing Latest Techniques
+Early data augmentation methods rely on hand-craft manipulations, which are constrained by effectiveness and efficiency.
+AutoAugment [13] begins to introduce reinforcement learning to find optimal augmentation policies, enabling automatic search for desired augmentation strategies.
+Generative models, such as LLMs and diffusion models, possess outstanding capacities to capture sample semantics and produce new samples with high flexibility and high quality [234, 162, 122].
+However, data augmentation is sometimes only a preprocessing step in a deep learning framework, and introducing generative models might introduce extra computation overhead.
+How to balance the tradeoff between the reward and cost of using these advanced techniques remains an open but critical problem.
+VIII Conclusion
+This survey presents a comprehensive summary of data augmentation techniques across five data modalities and proposes a modality-independent taxonomy from a data-centric perspective, which focuses on from where the augmented data is derived.
 It further enumerates related research papers on these data augmentation methods and annotates them with descriptive information in detail.
+Finally, it discusses several critical issues related to data augmentation that can help us better understand this technique.

### v4

@@
 By leveraging data augmentation techniques, AI models can achieve significantly improved applicability in tasks involving scarce or imbalanced datasets, thereby substantially enhancing AI models’ generalization capabilities.
-Existing literature surveys only focus on a certain type of specific modality data, and categorize these methods from modality-specific and operation-centric perspectives, which lacks a consistent summary of data augmentation methods across multiple modalities and limits the comprehension of how existing data samples serve the data augmentation process.
+Existing literature surveys only focus on a certain type of specific modality data and categorize these methods from modality-specific and operation-centric perspectives, which lacks a consistent summary of data augmentation methods across multiple modalities and limits the comprehension of how existing data samples serve the data augmentation process.
 To bridge this gap, this survey proposes a more enlightening taxonomy that encompasses data augmentation techniques for different common data modalities by investigating how to take advantage of the intrinsic relationship between and within instances.
 Additionally, it categorizes data augmentation methods across five data modalities through a unified inductive approach.
-Index Terms:
-Data augmentation, Data-centric taxonomy, Multi-modality
 I Introduction
 With the rapid development of Artificial Intelligence (AI) in the past decades, AI methods have shown their superiority over human beings and other traditional approaches across most tasks, from universal tasks in our daily lives such as image recognition [1] and text translation [2] to complicated scientific research tasks such as drug-drug interaction prediction [3] and single-cell RNA sequencing analysis [4].
-Recently, powered by Latent Diffusion Models [5] and Transformer [6], Stable Diffusion [5] and ChatGPT [7] have greatly changed human beings’ working, living, and entertainment.
+Recently, Stable Diffusion [5] and ChatGPT [6] have greatly changed human beings’ working, living, and entertainment.
 The success of AI products is usually attributed to AI models’ in-depth understanding of the accumulated data, thereby intrinsically uncovering data patterns and learning data-task correlations.
-The performances of AI models will be affected by the quantity and quality of training data, e.g.
-models trained with limited data suffer from over-fitting with significantly degenerative performance on testing datasets, and models trained with imbalanced samples will lead to poor generalization ability.
+The performances of AI models will be affected by the quantity and quality of training data, e.g., models trained with limited data suffer from overfitting with significantly degenerative performance on testing datasets, and models trained with imbalanced samples will lead to poor generalization ability.
 In most situations, researchers have to overcome difficulties caused by data scarcity and distribution imbalance.
-An intuitive solution to these problems is to acquire more data, but data are limited or difficult to collect in many cases, and labeling the data is yet another labor-intensive work.
+An intuitive solution to these problems is to acquire more data, but data are limited or difficult to collect in many cases, and labeling the data is yet another labor-intensive task.
 To solve these problems, data augmentation has been extensively applied and proven to be effective and efficient [7].
@@
 It also introduces more diversity and fills the gap between training datasets and real-world applications.
+Data augmentation is widely adopted in the real world and exhibits success in both the dataset preparation stage, such as the ImageNet-C and ImageNet-P datasets [8], which use data augmentation to simulate common image corruptions and perturbations, as well as the model training stage, such as scSiameseClu [9], which uses data augmentation to generate dual views for graph representation learning.
 As technology continues to evolve, a variety of techniques are applied to augment data samples.
 Some simple approaches only randomly mask part of the data [10, 11, 12], while more sophisticated augmentation schemes involve generative adversarial networks [13] or reinforcement learning agents [14].
-An ideal survey on data augmentation should be modality-independent, because only in this way can it focus on the intrinsic mechanisms of data augmentation, regardless of the data type, and thus provide insight into the nature of data augmentation.
-Existing literature surveys summarize data augmentation methods from different perspectives [14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29].
-As summarized in Table.
-I, each of them concentrates on a certain data modality.
-Most of them categorize data augmentation methods with modality-specific (e.g.
-image modality and text modality) or operation-centric (e.g.
-feature scaling operation, data perturbation operation) taxonomies.
+An ideal survey on data augmentation should be modality-independent, because only in this way can it focus on the intrinsic mechanisms of data augmentation, regardless of the data type, and provide insight into the nature of data augmentation.
+Existing literature surveys summarize data augmentation methods from different perspectives [15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33].
+As summarized in Table I, each survey concentrates on a certain data modality.
+Most of them categorize data augmentation methods with modality-specific (e.g., image modality and text modality) or operation-centric (e.g., feature scaling operation and data perturbation operation) taxonomies.
 These surveys reflect the development of data augmentation and its application in different learning scenarios.
 However, they fail to cover data augmentation methods across data types, uncover their common patterns despite different modalities, and thus limit readers’ understanding of the essence of data augmentation.
-To fill this gap, we investigate data augmentation methods comprehensively, across the five most popular data modalities.
+For example, originating from the idea of interpolation, mixup-based augmentation has been applied across different modalities, whereas such consistency is ignored in modality-specific discussions.
+To fill this gap, we investigate data augmentation methods comprehensively across the five most popular data modalities.
 We analyze how data augmentation methods utilize different sample numbers and how they leverage different components of the data information.
 Then, we propose a unified taxonomy that focuses on the way data augmentation leverages information from samples.
-Finally, we investigate up-to-date literature and categorize them by our proposed taxonomy.
+Finally, we investigate up-to-date literature and categorize it by our proposed taxonomy.
 The main contributions of this survey are as follows:
-  • We propose a novel modality-independent taxonomy from a data-centric perspective that accommodates data augmentation techniques for all modalities consistently and inductively;
-  • To the best of our knowledge, this is the first survey that covers data augmentation techniques across five data modalities, i.e. image, text, graph, tabular, and time-series modalities;
-  • We investigate how information is consistently contained in each of these data modalities and can be utilized for data augmentation;
-  • We include and categorize up-to-date literature in data augmentation.
+$\bullet$ We propose a novel modality-independent taxonomy from a data-centric perspective that accommodates data augmentation techniques for all modalities consistently and inductively.
+$\bullet$ To the best of our knowledge, this is the first survey that covers data augmentation techniques across five modalities: image, text, graph, tabular, and time-series data.
+$\bullet$ We investigate how information is consistently contained in each of these modalities and can be utilized for augmentation.
+$\bullet$ We include and categorize up-to-date literature in data augmentation and discuss future directions.
 TABLE I: A comparison of data modalities and taxonomies discussed by existing surveys and ours.
@@
 J Big Data 2023 [27]
-Neural Comput Appl 2023 [28]
+NCA [29]
 arXiv 2024 [31]
@@
 For example, the use of random distortion can be found in LeNet [34], where ninefold the distorted images are added to the dataset to verify that increasing the size of the training set can effectively reduce test error.
-Since then, data augmentation has gradually been recognized as a best practice for training Convolutional Neural Networks (CNNs) [34, 35].
-AlexNet [36] explicitly employs several data augmentation techniques to reduce overfitting.
+Since then, data augmentation has gradually been recognized as a best practice for training Convolutional Neural Networks (CNNs) [35].
+AlexNet [36] explicitly employs data augmentation to reduce overfitting.
 It augments the datasets by extracting patches from images and altering the color intensity.
-It also applies dropout, which randomly sets the output of some neurons to zero to prevent their co-adaption.
+It also applies dropout, which randomly sets the output of some neurons to zero to prevent their co-adaptation.
 Although dropout is often regarded as an approach to regularization, it has a similar idea to data augmentation and indeed inspires a set of augmentation methods [10].
 These applications of data augmentation mainly focus on increasing the size of the training dataset and introducing diversity to it.
-SMOTE [37] is another early and representative work on data augmentation but from a different perspective.
+SMOTE [37] is designed from a different perspective.
 Focusing on addressing the class imbalance problem, it suggests that oversampling the minority class can achieve better classification performance when categories are not equally represented in the dataset.
@@
 Data augmentation on text data has not been thoroughly researched as early as image data, possibly due to the discrete and correlated nature of text components.
-Still, several augmentation methods are proposed for or transplanted to text data.
+Still, several augmentation methods are proposed for text data.
 [19] is among the first surveys on text data augmentation.
 It adopts a simple taxonomy but gives more emphasis on the applications.
-Built on previous work on image data augmentation, [19] continues to investigate how data augmentation is used to enrich text datasets, and methods are classified as symbolic augmentation and neural augmentation.
+Built on previous work on image data augmentation, [20] continues to investigate how data augmentation enriches text datasets, and methods are classified as symbolic augmentation and neural augmentation.
 [21] categorizes text data augmentation methods by identifying whether they are used on the data space or feature space.
@@
 Till now, data augmentation techniques for different data types have been exhaustively discussed; however, there has not been a comprehensive survey that summarizes data augmentation for all types of data.
-Such a survey can be profitable because data augmentation techniques for different modalities share some common methodologies when leveraging information from existing data, such as altering some features of a sample or mixing values between multiple samples.
+Such a survey is profitable because data augmentation techniques for different modalities share some common methodologies when leveraging information from existing data, such as altering some features of a sample or mixing values between multiple samples.
 Summarizing these similarities can help reveal the common pattern of data augmentation.
@@
 Given a labeled dataset $\mathcal{D}_{L}=\{\mathbf{X},\mathbf{y}\}$, where $\mathbf{X}$ stands for the data, and $\mathbf{y}$ stands for the labels, data augmentation can be represented by a certain function $f_{\theta}$, such that the augmented dataset $\mathcal{\tilde{D}}_{L}=\{\mathbf{\tilde{X}},\mathbf{\tilde{y}}\}$ is derived by:
-$f_{\theta}:\mathcal{D}_{L}=\{\mathbf{X},\mathbf{y}\}\rightarrow\mathcal{\tilde{D}}_{L}=\{\mathbf{\tilde{X}},\mathbf{\tilde{y}}\}$ (1)
+$f_{\theta}:\mathcal{D}_{L}=\{\mathbf{X},\mathbf{y}\}\rightarrow\mathcal{\tilde{D}}_{L}=\{\mathbf{\tilde{X}},\mathbf{\tilde{y}}\}\vskip-5.69054pt$ (1)
 This representation also applies to an unlabeled dataset $\mathcal{D}_{U}=\{\mathbf{X}\}$ or a partially labeled dataset $\mathcal{D}_{P}=\{\mathbf{X}_{L}\cup\mathbf{X}_{U},\mathbf{y}_{L}\}$, but for simplicity, we only use a labeled dataset to explain the taxonomy.
@@
 RQ2: Which part of the information is used to generate new data?
-The answers to these questions constitute the taxonomy hierarchy which we present below.
+The answers to these questions constitute the taxonomy hierarchy, which we present below.
 By answering RQ1, we can divide data augmentation approaches into Single-instance Level Augmentation, Multi-instance Level Augmentation, and Dataset Level Augmentation, which build the first tier of the taxonomy in our survey.
@@
 We formalize single-instance level augmentation as follows, where $\epsilon(\mathbf{x}_{i})$ stands for a certain perturbation related to the original data $\mathbf{x}_{i}$:
-$\mathbf{\tilde{x}}=\mathbf{x}_{i}+\epsilon(\mathbf{x}_{i}),\qquad\tilde{y}=y_{i}$ (2)
+$\mathbf{\tilde{x}}=\mathbf{x}_{i}+\epsilon(\mathbf{x}_{i}),\qquad\tilde{y}=y_{i}\vskip-5.69054pt$ (2)
 We further answer RQ2 to build the second tier of the taxonomy.
@@
 We formalize multi-instance level augmentation as follows, where $i$ and $j$ are two data samples from the original dataset, and $\lambda$ is a scalar factor that often lies between $[0,1]$:
-$\mathbf{\tilde{x}}=\lambda\cdot\mathbf{x}_{i}+(1-\lambda)\cdot\mathbf{x}_{j},\qquad\tilde{y}=\lambda\cdot y_{i}+(1-\lambda)\cdot y_{j}$ (3)
+$\mathbf{\tilde{x}}=\lambda\cdot\mathbf{x}_{i}+(1-\lambda)\cdot\mathbf{x}_{j},\qquad\tilde{y}=\lambda\cdot y_{i}+(1-\lambda)\cdot y_{j}\vskip-5.69054pt$ (3)
 Now we answer RQ2 for multi-instance level augmentation.
@@
 Dataset Level Augmentation.
-This type of data augmentation method does not explicitly make use of one or multiple data but rather uses the entire dataset.
-It comprehensively learns the features and abstractions of the dataset, and samples completely new data from a learned distribution of the dataset.
+This type of data augmentation method does not explicitly make use of one or multiple data points but rather uses the entire dataset.
+It comprehensively learns the features and abstractions of the dataset and samples completely new data from a learned distribution of the dataset.
 The generated data should also fall into this distribution.
 We formalize dataset level augmentation as follows, where $P(\cdot)$ stands for a probability distribution learned from the original dataset $\mathcal{D}_{L}=\{\mathbf{X},\mathbf{y}\}$:
-$\vspace{-1.5mm}\mathbf{\tilde{x}}\sim P(\mathbf{X}),\qquad\tilde{y}\sim P(\mathbf{y}|\mathbf{\tilde{x}})$ (4)
+[…]
 The answer to RQ2 for dataset level augmentation is different from the two categories mentioned above.
@@
 We elaborate on how the proposed taxonomy applies to each data modality and review some typical data augmentation methods.
+[^1] A collection of discussed papers is available at https://github.com/ZaitianWang/Awesome-Data-Augmentation.
+TABLE II: Comparative summary of all methods, including detailed categories, operations, and corresponding analyses.
+# of samples Information type Modality Category Operation Analysis Computation Info loss
+Single-instance Value Image Pixel erasing Applying a rectangular- or grid-shaped mask over the image Moderate Negligible High
+Photometric transformation Brightness adjustment, color inversion, color casting, greyscaling, histogram equalization, kernel filters Negligible Moderate Negligible
+Text Token replacement Synonym, embedding similarity, knowledge graph Negligible Moderate Negligible
+Token addition Synonym, random punctuation Negligible Negligible Negligible
+Token deletion Randomly dropping tokens Negligible Negligible High
+Graph Node attribute masking Applying a mask over the node attribute matrix to hide certain nodes or features Negligible Negligible High
+Tabular Table masking Applying a mask over the table to hide some cells Negligible Negligible High
+Feature engineering Arithmetic operations High High Negligible
+Time-series Amplitude perturbation Scaling, magnitude warping, rotation, and jittering Negligible Negligible Negligible
+Sequence decomposition Extract components from a sequence High Moderate High
+Structure Image Image cropping Selecting an area and removing other parts Moderate Negligible High
+Geometric transformation Flipping, shearing, rotation, translation, stretching, distortion Negligible Moderate Negligible
+Text Sentence cropping Dividing sentence to pieces Moderate Negligible High
+Sentence morphing Inversion, passivization Moderate Negligible Negligible
+Graph Topology perturbation Node addition, edge addition, node dropping, edge dropping Negligible Negligible Moderate
+Subgraphing Cropping or substituting subgraphs Moderate Negligible High
+Graph diffusion Transition matrices sampled from random walk High Moderate Moderate
+Graph rewiring Swapping edges High Moderate Moderate
+Tabular Table subsetting Splitting table by columns Negligible Negligible High
+Time-series Window Slicing Selecting a span from a sequence Negligible Negligible High
+Window Morphing Downsampling, upsampling, shuffling Negligible Negligible Negligible
+Value-structure Image Policy-based DA Selecting an group of augmentation operations with an RL agent High High Moderate
+Prompt-based image editing Prompting visual generative models to edit image Negligible High Negligible
+Text Hierarchical data augmentation Augmenting at both word and sentence level High Negligible Moderate
+Back-translation Translating to an intermediate language and translating back to the original language Negligible High Negligible
+Paraphrase Rewriting sentence with alternative expressions while conveying the same meanings Negligible High Negligible
+Multi-instance Value Image Image mixup Mixing color from multiple images Negligible Negligible Moderate
+Text Text mixup Mixing word or sentence embeddings Negligible Moderate Moderate
+Graph Graph propagation Spreading node features or labels through edges to adjcacent nodes Moderate Negligible Moderate
+Tabular Tabular mixup Mixing cell values across rows Negligible Moderate Moderate
+Time-series Sequence averaging Calculating averaged sequence values Negligible Moderate Moderate
+Structure Image Image Patching Putting image segments together Negligible Negligible High
+Text Text fragment merging Swapping substructures between sentences Moderate Negligible High
+Tabular Table Concatenation Unifying table columns to concatenate tables by overlapping columns Moderate Negligible Negligible
+Value-structure Image Neural blending Mixing images through a neural network Negligible High Negligible
+Prompt-based mixing Prompting Diffusion Models to combine images Negligible High Negligible
+Text Generative mixing Prompting LLMs to mix sentences Negligible High Negligible
+Graph Graph mixup Aligning the topology and mixing node features Moderate Moderate Negligible
+Dataset Value-structure Cross-modality AE Using reconstructed sample as augmentation Negligible High Negligible
+GAN Using the GAN’s generator for data augmentation Negligible High Negligible
+LLM Prompting the LLM to generate large amount of textual samples Negligible High Negligible
+Diffusion Using discrete denoising diffusion model to generate new graphs Negligible High Negligible
+Exogenous Image, text Neural Style Transfer Introducing styles from other datasets High Moderate Negligible
+Tabular Table Retrieval Retrieve external database to fill and expend existing tables High Negligible Negligible
+Image Computer Graphic Modeling Designing scenes and objects with computer softwares High High Negligible
+Tabular Relational Structure Construction Building external graph structures over the table High Moderate Negligible
+Time-series Statistical Generation Using expert knowledge to expand sequences High Moderate Negligible
 IV Single-instance Level Augmentation
@@
 A pixel carries unit information that uses a numerical value to represent the color.
-Each pixel also has spatial relationships with other pixels, so, images are well-structured.
-Single-instance level augmentation for image data revolves around the perturbation of the image pixels’ colors and spatial relationships.
-For text data, the elemental components of text data are words (or sometimes, tokens).
+It also has spatial relationships with other pixels.
+Single-instance level image augmentation perturbs the image pixels’ colors or spatial relationships.
+For text data, the elemental components are words or tokens.
 A word often expresses minimal meanings.
 Words are organized into sentences following syntactic rules.
-The sentence is a sequential representation since one word comes after another, but there is also a tree-structured grammatical dependency behind the sentence.
-Single-instance level augmentation for text data often changes the wording or reshapes the syntax of a sentence.
-A graph is represented by $\mathbf{G}=\{\mathcal{V},\mathcal{E}\}$, where $\mathcal{V}$ stands for the vertex set, and $\mathcal{E}$ stands for the edge set.
-The topology structure can be alternatively represented by an adjacency matrix $\mathbf{A}\in\mathbb{R}^{|\mathcal{V}|\times|\mathcal{V}|}$, where $\mathbf{A}_{ij}$ is $1$ if there is an edge connecting vertex $i$ and vertex $j$, and is 00 otherwise.
-In a deep graph learning scenario, the graph also contains a feature matrix $\mathbf{X}\in\mathbb{R}^{|\mathcal{V}|\times d}$ that represents the attributes of the nodes.
+Single-instance level text augmentation often changes the wording or reshapes the syntax of a sentence.
+A graph is represented by $\mathbf{G}=\{\mathcal{V},\mathcal{E}\}$, where $\mathcal{V}$ and $\mathcal{E}$ stand for the vertex set and the edge set.
+The topology structure can be alternatively represented by an adjacency matrix $\mathbf{A}\in\mathbb{R}^{|\mathcal{V}|\times|\mathcal{V}|}$.
+The feature matrix $\mathbf{X}\in\mathbb{R}^{|\mathcal{V}|\times d}$ represents the attributes of the nodes.
 Hence another representation for a graph is $\mathbf{G}=\{\mathbf{A},\mathbf{X}\}$.
-Single-instance level augmentation for graph data is to perturb the values in the feature matrix $\mathbf{X}$ of nodes or the topology structure represented by the adjacency matrix $\mathbf{A}$.
-A table is made up of cells, which store numerical or categorical values that represent attributes of the described entities.
+Single-instance level graph augmentation perturbs the values in the feature matrix $\mathbf{X}$ or the topology structure represented by the adjacency matrix $\mathbf{A}$.
+A table is made up of cells with numerical or categorical values that represent attributes of the described entities.
 Cells are organized into columns and rows, but this structural information is trivial, and switching the sequence of the columns and rows does not affect the information conveyed by the table.
-So, most data augmentation methods leverage the value information.
-Still, it is also possible to perturb the structure information of a table.
-Time-series data represent sequential values within a period, and each timestamp is considered an elemental component.
+So, most data augmentation methods leverage the value information with only a few exceptions.
+Time-series data represent sequential values within a time span, and each timestamp is considered an elemental component.
 Data augmentation methods for time-series data either perturb the value at each timestamp or perturb the sequence of timestamps.
@@
 Pixel erasing masks out all color information in an area of the data; photometric transformation imposes a soft adjustment to change the appearance of the image, but the essential information remains visible.
+Both methods are easy to implement, require none to minimal analysis of the sample, and can be applied to various tasks.
 Pixel Erasing
 As illustrated in Fig.
-2b, the idea of pixel erasing is to mask some information from the original input space or the feature space.
-Cutout [9] and random erasing [10] first selects a rectangle region and erases all pixels in that region.
-The erased region is then filled with a color or Gaussian noise.
-Random Erasing typically adds one or two big rectangular masks to the image, and some other methods deploy more flexible implementations.
+2b, pixel erasing masks some information from the original sample.
+Cutout [10] and random erasing [11] select a rectangular region, erase all pixels in that region, and then fill the region with a color or Gaussian noise.
 Hide-and-Seek [38] randomly removes multiple parts of the image so that the model has to learn from multiple parts of the image other than the most discriminating part.
-Similarly, the Grid-Mask method [39] generates a set of squared masks in the grid layout.
+Grid-Mask [39] generates a set of squared masks in the grid layout.
 It can prevent excessive deletion and reservation and achieve a reasonable balance between deleted and reserved regional information.
-In most research, random erasing is used in the spatial domain, while REF [40] extends the use of random erasing to the frequency domain.
+REF [40] extends the use of random erasing from the spatial domain to the frequency domain.
 It uses the discrete Fourier transform to process the image, applies random erasing on the image’s frequency domain, and uses the inverse discrete Fourier transform to recover the image to the spatial domain.
+Despite its easy application, pixel erasing is subject to removing critical parts from the sample, and some solutions are discussed in Section VII-B1.
 Photometric Transformation
-Augmentation that changes the overall color and style of the image is called photometric transformation, as shown in Fig 2c.
+As shown in Fig.
+2c, photometric transformation changes the overall color and style of the image.
 The simplest way is to adjust the brightness of the image.
 Color inversion replaces the pixels with their “opposite” color.
-Color casting applies different operations in different channels, for example, increasing the image’s red channel while keeping the green channel and blue channel untouched.
+Color casting applies different operations in different channels.
 Noise injection adds Gaussian or other types of noises to the images.
-Greyscaling transforms a colored picture into black-and-white by combining channels for different colors, forcing the model to focus on the shapes or edges of target objects instead of their colors.
-Histogram equalization [41, 42] enhances the image by enhancing its contrast, and white balancing simulates alternative light cast on the object.
+Greyscaling transforms a colored picture into black-and-white by combining channels for different colors.
+Histogram equalization [41] enhances the image by enhancing its contrast, and white balancing simulates alternative light cast on the object.
 Kernel filters are matrices that slide across and are convoluted with the original image to generate new images, either sharpening or blurring the image.
@@
 IV-A2 Wording Perturbation
-Text data augmentation of this type uses word-level manipulation. It applies changes to the choice of words or adds or removes words. Some examples are presented in Table II.
+Token replacement, addition, and deletion perform word-level manipulation. Some examples are presented in Table III.
 Token Replacement
-Token replacement refers to using alternative wording in a given sentence.
-EDA [11] proposes a set of text augmentations, including synonym replacement.
-It randomly selects several words from a sentence, substituting them with their synonyms, i.e.
-words that mean exactly or nearly the same.
-[43] uses WordNet [44], a lexical database to find synonyms for verbs and nouns in sentences and thus make replacements.
+Token replacement uses alternative wording in a sentence.
+EDA [12] proposes synonym replacement.
+It randomly selects words from a sentence, substituting them with their synonyms, i.e., words that mean exactly or nearly the same.
+[42] uses WordNet [43], a lexical database, to find synonyms for verbs and nouns and make replacements.
 It also derives word embeddings by Word2Vec [44] and replaces a word with others that have high cosine similarity.
-Another possibility is to use a knowledge graph.
-A knowledge graph describes the relationships of entities, and by utilizing the “is equivalent” relation, we can find a series of synonyms for a given word.
+KG-NMT [45] resorts to a knowledge graph for terminology equivalence to augment samples and improve translation performance.
 Token Addition
-Token addition refers to adding tokens to the original text, either introducing junk tokens as noise or enriching the sentences and making them more informative.
+Token addition either introduces junk tokens as noise or enriches the sentences to make them more informative.
 The random insertion, proposed in EDA [12], inserts a random synonym of a random word at a random position in the sentence.
@@
 In [47], noising techniques are used to avoid overfitting, and the randomly selected word is replaced by a blank placeholder token “_”, rather than simply being removed.
-In [48], apart from word dropout which is similar to random deletion, it also uses word blanking to replace a word with a special “¡BLANK¿” placeholder.
-IV-A3 Node Attributes Perturbation
-Graph nodes often contain rich attributes in the feature matrix. This matrix is generally similar to tabular data, so augmenting node attributes does not need to consider the specialty of graph data.
-Attribute Masking and Noise Injection
-Various typical augmentation approaches on node attributes from random masking [49, 50, 51] to perturbing attributes with random noises [52] can be applied, as we do for common tabular data.
-These methods are unified in Eq.
-5, where $\mathbf{M}\in\{0,1\}^{N\times N}$ is a binary mask, $\odot$ denotes the Hadamard product, and $\mathbf{N_{W}}\sim\mathcal{N}(0,\Sigma)$ is the Gaussian white noise.
-$\mathbf{\tilde{X}}=\mathbf{X}\odot\mathbf{M}+\mathbf{N}_{W}$ (5)
-Some other methods are more sophisticated.
-For example, [53] first learns the topology information of the graph by random walk, and uses the topology embedding together with the original features as inputs to a dual graph neural network.
-Similarly, LA-GNN [54] applies a local augmentation mechanism.
-It samples features from nodes’ neighborhoods as additional information when augmenting the attributes.
-A different approach is proposed in DropMessage [55].
-Instead of directly modifying the attributes on the nodes, it applies random dropping in the message-passing phase.
+In [48], apart from word dropout, which is similar to random deletion, it also uses word blanking to replace a word with a special “¡BLANK¿” placeholder.
+IV-A3 Node Attributes Masking and Noise Injection
+[…]
 IV-A4 Cell Value Perturbation
-Tabular data is much simpler compared with other modalities, so it is hard to design complex augmentation methods that make use of spatial or sequential relationships within a data sample as we do for image or text data, and most data augmentation for tabular data perturbs the value stored in the table cells.
-Table Masking is a simple technique often applied for tabular data augmentation.
-These corrupted data are often used with the original data under a self-supervised learning framework to learn stronger representations of the dataset.
-Another approach to tabular data augmentation borrows ideas from Feature Engineering.
-Some work on Feature Engineering automates feature selection and transformation to generate optimal feature space, and the operations can also be used to augment tabular datasets.
+Table masking is a simple technique often applied for tabular data augmentation that corrupts cell values.
+Another approach borrows ideas from feature engineering, which automates feature selection and transformation to generate optimal feature space, and the operations can also be used to augment tabular datasets.
 Table Masking
 VIME [56] uses a mask generator to corrupt features.
-In a self-supervised scenario, two pretext tasks respectively reconstruct the original data and estimate the mask vector.
-The model updates its weights based on the reconstruction loss and the mask vector estimation loss of the two pretext tasks.
-VIME also considers a semi-supervised scenario.
-In that case, multiple masks are generated and applied to the feature of the unlabeled samples.
-Despite not being labeled, the model predictions on these corrupted samples should be similar, and thus a consistency loss is calculated.
-For labeled samples, a supervised loss is calculated based on the prediction.
-The two losses jointly contribute to the back-propagation phase.
+In a self-supervised scenario, two pretext tasks reconstruct the original data and estimate the mask vector, respectively.
 SCARF [57] applies data augmentation in a contrastive learning framework.
-Its corruption process is similar to that of VIME but it uses InfoNCE loss to evaluate the similarity between the embeddings of the corrupted and original views and then update the model accordingly.
-MET [58] works in a pattern similar to VIME and SCARF, which calculates the reconstruction loss of a partially masked sample.
+Its corruption process is similar to that of VIME, but it uses InfoNCE loss to evaluate the similarity between the embeddings of the corrupted and original views and then update the model accordingly.
+MET [58] works in a similar pattern, which calculates the reconstruction loss of a partially masked sample.
 Feature Engineering
-Feature transformation [59, 60] and feature generation techniques can effectively enlarge tabular datasets by generating tables with different feature sets [61].
-Powered by reinforcement learning, GRFG [62] transforms existing features with cascading agents.
-It uses arithmetic operations to provide better traceability and explainability.
-MOAT [63] extends feature transformation to continuous space.
-It trains an encoder to project the feature transformation process to continuous embeddings and decodes the optimized embeddings for the best feature transformation strategies.
+Feature transformation and generation techniques can effectively enlarge tabular datasets by generating tables with different feature sets [59, 60, 61].
+GRFG [62] transforms existing features with cascading agents that search for the best arithmetic operation sets.
+MOAT [63] extends feature transformation to continuous space by decoding the optimal transformation strategies from continuous embeddings.
 LFG [64] and TIFG [65] leverage LLM agents to generate new features that are both explainable and effective.
 TFWT [66] computes weights for samples and features, and the weights are combined with the data for later tasks.
-Although these studies focus on feature engineering, their way of thinking and application of reinforcement learning is very enlightening for data augmentation tasks.
+Although these studies focus on feature engineering, their way of thinking and application of reinforcement learning are very enlightening for data augmentation tasks.
 IV-A5 Sequence Value Perturbation
-The value information of a sequence is carried on the Y-axis, represented by the amplitude at specific timestamps or frequency channels.
-Amplitude Perturbation augments time-series data by altering the Y-axis values, and Decomposition decomposes the value of a sequence into different parts that express the sequential values from different aspects.
+The value information of a sequence is carried on the Y-axis, represented by the amplitude at specific timestamps.
+Amplitude perturbation augments time-series data by altering the Y-axis values, and sequence decomposition extracts different parts that express the sequential values from different aspects.
 Amplitude Perturbation
-Jittering adds random noise to the value at each timestamp of a sequence. In [67], jittering is used to simulated additive noise $\epsilon$ of wearable sensors:
-$\tilde{x}(\epsilon)=\{x_{1}+\epsilon_{1},\cdots,x_{t}+\epsilon_{t},\cdots,x_{T}+\epsilon_{T}\}$ (6)
-Scaling [67] alters the data by multiplying the magnitude of the entire sequence with a scalar $\alpha$.
-$\tilde{x}(\alpha)=\{\alpha x_{1},\cdots,\alpha x_{t},\cdots,\alpha x_{T}\}$ (7)
-Magnitude warping [67] is similar to scaling, but it uses a smooth curve to warp the sequence so that magnitude at different timestamps is scaled with different parameters.
-$\tilde{x}(\alpha)=\{\alpha_{1}x_{1},\cdots,\alpha_{t}x_{t},\cdots,\alpha_{T}x_{T}\}$ (8)
-Rotation [67], or flipping, assumes the data is symmetric with respect to the X-axis, and makes the data upside down by applying a rotation matrix $R$.
-$\tilde{x}(R)=\{Rx_{1},\cdots,Rx_{t},\cdots,Rx_{T}\}$ (9)
+Jittering adds random noise to the value at each timestamp of a sequence.
+In [67], jittering is used to simulate additive noise $\epsilon$ of wearable sensors: $\tilde{x}(\epsilon)=\{x_{1}+\epsilon_{1},\cdots,x_{t}+\epsilon_{t},\cdots,x_{T}+\epsilon_{T}\}$.
+Scaling [67] alters the data by multiplying the magnitude of the entire sequence with a scalar $\alpha$: $\tilde{x}(\alpha)=\{\alpha x_{1},\cdots,\alpha x_{t},\cdots,\alpha x_{T}\}$.
+Magnitude warping [67] resembles scaling but applies a smooth curve to the sequence so that magnitude at each timestamp is scaled with different parameters: $\tilde{x}(\alpha)=\{\alpha_{1}x_{1},\cdots,\alpha_{t}x_{t},\cdots,\alpha_{T}x_{T}\}$.
+Rotation [67], or flipping, assumes the data is symmetric with respect to the X-axis and makes the data upside down by applying a rotation matrix $R$: $\tilde{x}(R)=\{Rx_{1},\cdots,Rx_{t},\cdots,Rx_{T}\}$.
 Apart from the amplitude, the phase spectrum provides another option for data augmentation.
-RobustTAD [68] converts the time-series data from the time domain to the frequency domain, as shown in Eq.
-10, and then increases all phases $\theta(\omega_{k})$ by a small perturbation.
-$F(\omega_{k})=\frac{1}{N}\sum^{N-1}_{t=0}{x_{t}e^{-j\omega_{k}t}}=A(\omega_{k})exp[j\theta(\omega_{k})]$ (10)
+RobustTAD [68] converts the time-series data from the time domain to the frequency domain, as shown in equation: $F(\omega_{k})=\frac{1}{N}\sum^{N-1}_{t=0}{x_{t}e^{-j\omega_{k}t}}=A(\omega_{k})exp[j\theta(\omega_{k})]$, and then increases all phases $\theta(\omega_{k})$ by a small perturbation.
 Decomposition
-Empirical Mode Decomposition (EMD) [69] is used to produce a set of intrinsic mode functions (IMFs).
+Empirical Mode Decomposition (EMD) [69] is used to produce a set of intrinsic mode functions.
 This decomposition relies on extracting the energy linked to different time scales that are inherent to the data.
 EMD has been utilized as a method of data augmentation in [70] for the purpose of categorizing impact noise in vehicles.
-Another way to decompose time-series data is STL [71], which decomposes signals into seasonal, trend, and remainder:
-$y_{t}=\tau_{t}+s_{t}+r_{t},\quad t=1,2,\cdots N$ (11)
-Other works such as [72] and [68] focus on improving robustness with the help of STL components.
+Seasonal-trend decomposition (STL) [71] decomposes signals into seasonality, trend, and remainder: $y_{t}=\tau_{t}+s_{t}+r_{t},\quad t=1,2,\cdots N$ [72] and [68] focus on improving robustness with the help of STL components.
 IV-B Structure-based Transformation
 IV-B1 Shape Transformation
-This type of data augmentation manipulates the spatial relationship between pixels.
 Image cropping breaks the original image into pieces (Fig.
@@
 2e).
+Both of them manipulate the spatial relationship between pixels.
 Image Cropping
 Image cropping selects an area from the original image and removes other parts of the image.
-It includes manual selection, random cropping, attention-based, aesthetics-based, and change-based approaches [73].
-Attention-based approaches use certain attention scores to evaluate a subimage.
-Regions that contain the main topic or other types of information often have high attention scores.
-[74] uses the entropy, area ratio, and center distance of the image to evaluate the performance of the image segmentation and accordingly adjust the cropping parameters.
-[75] uses eye-tracking techniques to capture human fixations and thus infer the location of important contents on the image.
+Attention-based approaches use attention scores to evaluate a subimage.
+Regions with the main topic often have high attention scores.
+[73] uses the entropy, area ratio, and center distance of the image to evaluate the performance of the image segmentation.
+[74] uses eye-tracking techniques to capture human fixations and infer the location of important content on the image.
 Aesthetics-based approaches [75, 76, 77] are similar to quality evaluation.
-For example, [76] detects several subject regions and one background region, extracts features from the regions, estimates and combines posterior probabilities, and determines the quality score according to these probabilities.
+[75] detects several subject regions and one background region, extracts features from the regions, estimates and combines posterior probabilities, and determines the quality score according to these probabilities.
 A2-RL [78] integrates image cropping into a sequential decision-making process and an aesthetics-aware reinforcement learning model to solve the best cropping problem.
-Change-based approaches such as [80] take into account the content that is removed and extract two sets of specially designed features to model what is changed after cropping.
+Change-based approaches, such as [79], take into account the content that is removed and extract two sets of specially designed features to model what is changed after cropping.
 Geometric Transformation
-As one of the most classical image data augmentation methods, geometric transformation changes the spatial relationship of pixels within an image while keeping the value of each pixel untouched to create more data without extra sampling or labeling.
+Geometric transformation changes the spatial relationship of pixels within an image while keeping the value of each pixel untouched.
 Affine transformations like flipping, shearing, rotation, and translation manipulate the image while preserving the colinearity and ratios of distances.
-Image flipping reflects the image around its vertical axis, horizontal axis, or both axes, rotation turns the image by an angle about a fixed point.
+Image flipping reflects the image around its vertical axis, horizontal axis, or both axes; rotation turns the image by an angle about a fixed point.
 Translation means moving an image in a certain direction.
 It can introduce more variance in the position of target objects in the image.
-In non-affine transformation, the spatial relationship between pixels is further changed compared with affine transformations because the colinearity and ratio of distances are broken.
+Non-affine transformation allows breaking the colinearity and ratio of distances when reshaping the images.
 Perspective transformation [80] mimics images taken from other views of the object.
-Image stretching [82] preserves or even enlarges the prominent contents in the image and compresses the background or unimportant foreground information.
-Distortion [34, 83] also stretches the pixels but often focuses on a certain region.
+Image stretching [81] preserves or even enlarges the prominent contents and compresses the background or unimportant foreground information.
+Distortion [35, 82] focuses on stretching the pixels in a certain region.
 PatchShuffle [83] divides an image into block matrices and shuffles the pixels within each block under a certain probability.
 IV-B2 Syntax Transformation
-We categorize syntax transformation as: Sentence Cropping that splits a sentence into many pieces, and Sentence Morphing that swaps the location of words or grammatical components.
-Some examples are presented in Table II.
+We categorize syntax transformation as: sentence cropping that splits a sentence into many pieces and sentence morphing that swaps the location of words or grammatical components.
+Some examples are presented in Table III.
 Sentence Cropping
-Sentence cropping [85] divides sentences into several small pieces.
-It often uses dependency trees to extract the relationships between subjects and objects of the sentence and produces multiple smaller sentences that focus on different parts of the original sentence.
+Sentence cropping [84] divides sentences into small pieces.
+It uses dependency trees to extract the relationships between subjects and objects of the sentence and produces multiple smaller sentences that focus on different parts of the original sentence.
 This operation breaks the syntactic solidity and semantic comprehensiveness of the entire sentence but preserves local syntactic tags and shallow semantic labels.
@@
 Random swap [12] randomly chooses two words in a sentence and swaps their position.
-This perturbs the structure and adds noise to the text data across paragraphs.
 [85] studies two types of syntactic augmentations: inversion and passivization.
@@
 Passivization converts a sentence to its passive form, and the meaning of the new sentence is typically consistent with the original one.
-Sentence rotation [85] borrows the idea of image rotation, where the image is rotated around its center.
-In sentence rotation, the root of the sentence is chosen as the center of the sentence, and other sentence fragments are connected to their center via the dependency tree.
-TABLE II: Demonstration of some elementary operations for text data augmentation
+Sentence rotation [84] chooses the root of the sentence’s dependency tree as the center (as in an image) and swaps other fragments.
+TABLE III: Demonstration of some elementary operations for text data augmentation
 Operation Sentence
@@
 Graph topology structure has immense information representation capacity.
-Naive Topology Perturbation methods modify a single node or edge component or consider a set of nodes as a whole and modify the substructure.
+Naive topology perturbation methods modify a node or an edge.
 Subgraphing extracts subgraphs from the complete graph.
@@
 Due to their straightforward idea and easy implementation, node dropping and node addition are among the earliest augmentation approaches in graph learning.
-As Eq.
-12 shows, Node dropping [49] randomly discards some nodes from the graph.
-Such an operation will also remove the attributes that these nodes carry and the edges on these nodes.
-$\tilde{\mathcal{V}}=\text{Drop}(\mathcal{V}),\qquad\tilde{\mathbf{G}}=\{\tilde{\mathbf{A}},\tilde{\mathbf{X}}\}=\{\mathbf{A}[\tilde{\mathcal{V}},\tilde{\mathcal{V}}],\mathbf{X}[\tilde{\mathcal{V}},:]\}$ (12)
+Node dropping [49] randomly discards some nodes from the graph: $\tilde{\mathcal{V}}=\text{Drop}(\mathcal{V}),\tilde{\mathbf{G}}=\{\tilde{\mathbf{A}},\tilde{\mathbf{X}}\}=\{\mathbf{A}[\tilde{\mathcal{V}},\tilde{\mathcal{V}}],\mathbf{X}[\tilde{\mathcal{V}}$,:]\}.
 As for node addition, an example is MPNN [86].
@@
 An edge describes the relationship between two nodes, and by perturbing these relationships through edge dropping and edge addition, we also introduce diversity to the structural information of a graph.
-DropEdge [88] perturbs the edge set by randomly setting a portion of the adjacency matrix to zero.
-It first produces a sparse matrix from a random subset of the original edges, then derives the augmented edge set by subtracting the sparse matrix from the original adjacency matrix.
-Edge masking in [51] constructs a Bernoulli distribution, and randomly masks edges under that probability.
-FairDrop [89] computes a mask according to the network’s homophily, and uses this mask to guide the dropping process to remove unfair connections.
-AD-GCL [90] applies a GNN-augmenter to learn how to drop edges so that the mutual information of the augmented graph and that of the original graph is minimized.
+DropEdge [87] randomly sets a portion of the adjacency matrix to zero.
+Edge masking in [51] constructs a Bernoulli distribution and randomly masks edges under that probability.
+FairDrop [88] computes a mask according to the network’s homophily and uses this mask to guide the dropping process to remove unfair connections.
+AD-GCL [89] applies a GNN augmenter to learn how to drop edges so that the mutual information of the augmented graph and that of the original graph is minimized.
+Despite its popularity, graph topology perturbation is potentially risky in some cases and must be used carefully.
+The removal of some crucial edges might influence the connectivity of the graph, thereby limiting the use of some graph algorithms that require the graph to be fully connected.
+Conversely, adding some inter-class edges might escalate representation collapse caused by the inductive bias of the GNN models.
 Subgraphing
-Subgraph cropping is based on the idea that graph semantics are partially represented in local structures [49].
-Some subgraphing methods, such as those used in GCC [91] and SUGAR [92] use random walk to sample a subgraph from the original graph.
-GrpahCrop [93] applies a node-centric strategy to crop connectivity-preserved subgraphs and expand the dataset with these augmented subgraphs.
-These subgraphs are able to maintain some topology characteristics of the original graph.
-Similarly, SUBG-CON [94] uses the Personalized PageRank algorithm to select the important nodes.
+Subgraph cropping assumes graph semantics are partially represented in local structures [49].
+GCC [90] and SUGAR [91] use random walk to sample a subgraph from the original graph.
+GraphCrop [92] applies a node-centric strategy to crop connectivity-preserved subgraphs and expand the dataset with these augmented subgraphs.
+SUBG-CON [93] uses the Personalized PageRank algorithm to select the important nodes.
 NeuralSparse [94] proposes to obtain the subgraph by learning.
 It uses a sparsification network to generate sparsified subgraphs.
-Subgraph Substitution perturbs the substructure as a whole of a graph, which is especially beneficial for molecular graphs.
-A molecule graph represents atoms with nodes and chemical bonds with edges.
-Functional groups, which often decide the chemical characteristics of molecules, can be represented by substructures.
-So, by wisely sampling substructures, the global semantics of a molecule graph can be well contained.
-MoCL [96], for example, utilizes substructure information as one of its augmentation schemes.
-It performs graph augmentation by replacing functional groups, effectively introducing reasonable noise.
+Subgraph Substitution perturbs the substructure as a whole of a graph.
+It is especially beneficial for molecular graphs, which represent atoms with nodes and chemical bonds with edges.
+Functional groups, often deciding the chemical characteristics of molecules, can be represented by substructures.
+MoCL [95] utilizes substructure information as one of its augmentation schemes.
+Its replacement of functional groups effectively introduces reasonable noise.
 GEAR [96] also employs such a technique.
@@
 Graph Diffusion
-Traditional message passing in graph neural networks only leverages one-hop connections, while the edges of a real graph are often noisy or arbitrarily defined, and many relationships in the networks are indirectly contained in higher-order neighborhoods.
-To address this issue, GDC [98], as in Eq.
-$\mathbf{\tilde{A}}=\sum_{k=0}^{\infty}\theta_{k}\mathbf{T}^{k}$ , proposes to calculate a series of transition matrices $\mathbf{T}^{k}$ by means such as random walk and combines these transition matrices with weights $\theta_{k}$ to obtain a new adjacency matrix $\mathbf{\tilde{A}}$ and thus generate a new graph that directly contains some node connection hidden in the original graph.
-The new graph can also be understood as being processed by a denoising filter because graph diffusion smooths out the neighborhood over the graph [98].
-This technique is widely used by self-supervised approaches to transform the graph and generate different views for contrastive learning, such as MVGRL [99], SelfGNN [100], and MERIT [101].
+[…]
 Graph Rewiring
-Graph rewiring, as presented in Eq.
-13, is initially proposed as randomly swapping edges [102] to improve a network’s robustness and mitigate malicious attacks.
-Some later research such as smart rewiring [103] takes the degrees of chosen nodes when swapping their edges.
-Stochastic Discrete Ricci Flow [104] is a curvature-based graph rewiring method, which in each step adds an edge that can improve the most negatively curved edge, and then removes the edge with the most positive curvature.
-Its results demonstrate the ability of graph rewiring to alleviate the over-squashing problem.
-DiffWire [105] introduces the CT-LAYER and GAP-LAYER in a GNN model, which respectively learns the commute times and spectral gap of the graph.
-It rewires the adjacency graph so that the graph’s spectral gap is minimized, and by doing so tries to predict the optimal graph structure for the downstream task.
-DHGR [106] first learns a similarity matrix from the original graph.
-Then it rewires the graph by adding edges between high-similarity node pairs and removing edges from low-similarity ones.
-Such an operation can effectively improve the homophily of the graph and thus improve GNN performance.
-Following DHGR [106], HDHGR [107] computes multiple similarity matrices under different meta-paths and thus generates several rewired meta-path subgraphs.
-$\vspace{-2mm}\mathbf{\tilde{A}}=\overbrace{\mathbf{A}\odot(\mathbf{1}-\mathbbm{1}_{r})}^{\text{edge removal}}+\overbrace{(\mathbf{1}-\mathbf{A})\odot\mathbbm{1}_{r}}^{\text{edge addition}}$ (13)
-IV-B4 Table Structure Transformation
-Table Subsetting
+[…]
+IV-B4 Table Subsetting
 Inspired by image cropping [107, 79], SubTab [108] divides a table into multiple subsets by splitting its columns.
@@
 IV-B5 Sequence Transformation
-Unlike Value Perturbation, sequence perturbation is implemented on the X-axis, either on the time domain or the frequency domain.
+Unlike value perturbation, sequence perturbation is implemented on the X-axis.
 Window slicing splits the time-series data into pieces.
@@
 Window Slicing
-Inspired by image cropping, which is widely used for image data augmentation, MCNN [109] proposes window slicing.
-It extracts slices from the entire sequence as new training samples.
-Each slice is assigned with the label of the original sequence.
+Inspired by image cropping, MCNN [109] proposes window slicing. It extracts slices from the entire sequence as new training samples. Each slice is assigned the label of the original sequence.
 Sequence Morphing
@@
 This is meant to be consistent with how children speak.
-[67] propose to randomly permeate these slices and concatenate them as a long sequence.
-Similar processing is also called Random Shuffling and is used in [112].
+Random shuffling [67, 112], randomly permutates these slices and concatenates them as a long sequence.
 VTLP [113] applies window wrapping in the frequency domain.
 It adds a small distortion to the central frequency.
-SpecAugment [114] also proposes to augment the frequency domain.
-It generates masks of length sampled from a uniform distribution and uses each of them to mask consecutive frequency channels.
+SpecAugment [114] also augments the frequency domain.
+It masks consecutive frequency channels of length sampled from a uniform distribution.
 IV-C Value-structure Transformation
 IV-C1 Image Value-structure Transformation
-Policy-based methods search for the optimized augmentation policy that includes both value-based transformations and structure-based transformations, as demonstrated in Fig.
+Policy-based methods search for the optimized augmentation policy that includes both value-based and structure-based transformations, as demonstrated in Fig.
 2f.
-Prompt-based Image Editing employs a text-to-image model to change the content, background, or style of the image as humans desire.
-When implementing the augmentation, both coloring and the positions of pixels are changed (Fig.
+Prompt-based image editing employs a text-to-image model to change the content, background, or style of the image as humans desire, changing both the coloring and the position of pixels (Fig.
 2g).
 Policy-based Data Augmentation
-AutoAugment [13] represents data augmentation as a decision-making process and uses reinforcement learning to search for the best operations.
-In AutoAugment, a policy includes a group of basic data augmentation operations with probabilities and magnitudes.
-AutoAugment processes images under the policy, evaluates these augmented images on downstream networks, and rewards a policy generator for better performance.
-Fast AutoAugment [115] proposes a more efficient searching algorithm for augmentation policies, which works significantly faster than AutoAugment and thus allows a larger number of sub-policies, which should benefit the generalization performance.
-RandAugment [116] removes the need for a separate search phase on a proxy task, and makes the training process less complicated and less expensive.
+A policy includes a sequential group of basic augmentation operations with probabilities and magnitudes.
+AutoAugment [14] represents data augmentation as a decision-making process and uses reinforcement learning to search for the best policy.
+It processes images under the policy, evaluates these augmented images on downstream networks, and rewards a policy generator for better performance.
+Fast AutoAugment [115] proposes a more efficient searching algorithm, which is faster and thus allows a larger number of sub-policies, benefiting the generalization performance.
+These methods rely on independent reinforcement learning processes for each task, and the searched policies do not guarantee optimality when transferred to other tasks, making the methods cost-inefficient.
+RandAugment [116] removes the need for a separate search phase on a proxy task and makes the training process less complicated and less expensive.
 Population Based Augmentation [117] learns an augmentation schedule rather than a fixed policy as in AutoAugment to improve efficiency.
@@
 Prompt-based Image Editing
-Visual generative models have been rapidly evolving in recent years.
-These models are typically conditioned on text embeddings, allowing flexible image editing by modifying textual descriptions of images and prompting the models.
+Visual generative models are typically conditioned on text embeddings, allowing flexible image editing by modifying textual descriptions of images and prompting the models.
 DA-Fusion [118] learns fine-grained concepts, replaces the original category labels, and uses the Stable Diffusion model to generate new images.
 SaSPA [119] uses GPT-4 to generate prompts based on the sample’s meta-class, then inputs the prompts, reference images, and detected edges to the diffusion model to generate augmented images.
+Generating the final augmentation set with diffusion models can be labor-intensive because models are very sensitive to the prompts, so users must prepare detailed image descriptions carefully and might have to examine the prompts repeatedly.
 IV-C2 Text Value-structure Manipulation
 Hierarchical data augmentation explicitly augments both value and structure information of text data.
-Back-translation Augmentation and Paraphrase Augmentation use neural networks to generate sentences that are semantic-preserving but different from original sentences in both wording and syntax.
+Back-translation and paraphrase augmentation use neural networks to generate sentences that are semantic-preserving but different from original sentences in both wording and syntax.
 Hierarchical Data Augmentation
-Hierarchical Data Augmentation (HDA) [120] augments text data at both word and sentence levels.
+Hierarchical data augmentation [120] augments text data at both word and sentence levels.
 It evaluates the roles of words and sentences by training a hierarchical attention network to obtain the attention values of the text data at both levels.
 Attention values help decide the importance of components and augment the data by cropping and concatenating these components.
-Back-translation Augmentation
-Back-translation Augmentation.
-Back-translation [121] means translating text data from one language to another, or a sequence of other languages, and finally translating back to the original language.
+Back-translation
+[121] proposes to translate text data from one language to another, or a sequence of other languages, and finally translates back to the original language.
 Semantic information is preserved, but the translation process can introduce some variance in syntax and wording.
-Such transformation can help the model learn more flexible expressions that convey the same meaning.
+Such transformations can help the model learn more flexible expressions that convey the same meaning.
 Paraphrase Augmentation
-AugGPT [122], for example, prompts the state-of-the-art ChatGPT model to rewrite sentences in the dataset and preserves dataset coherence after data augmentation.
-Taboo [123] requires paraphrasing without using some frequently used words, Chain [124] paraphrase one sentence multiple times, and Hint [124] provide some examples, to encourage the LLM response otherwise.
+AugGPT [122], for example, prompts the state-of-the-art large language model (LLM) to rewrite sentences in the dataset and preserves dataset coherence after data augmentation.
+Taboo [123] requires paraphrasing without using some frequently used words, Chain [123] paraphrases one sentence multiple times, and Hint [123] provides some examples to encourage the LLM response otherwise.
 DoAug [124] constructs a preference dataset based on the diversity of paraphrased sentences and trains the LLM with the DPO to boost its generation diversity.
+Although such paraphrase methodology is capable of preserving sentence meanings, it is still possible that the paraphraser misunderstands the nuanced semantics and causes semantic drift that invalidates the sample’s label.
 V Multi-instance Level Data Augmentation
-Multi-instance augmentation combines value and structure information from different samples. Specific operation varies depending on the data modality in which it is applied.
+Multi-instance augmentation combines value and structure information from different samples. The specific operation varies depending on the data modality in which it is applied.
 Multi-instance level image augmentation combines the visual information, i.e., pixels, from different images.
@@
 V-A Value-based Mixture
-V-A1 Pixel Mixture
-Image Mixup computes an arithmetical mixture of the color from different images and Neural Blending resorts to neural networks to enhance such process.
-Image Mixup
-Mixup [126] mixes two random images on the input layer by calculating their interpolation pixel-wise. Pixels from two original images each contribute a portion to the new image:
-$\vspace{-2mm}\mathbf{\tilde{x}}=\lambda\mathbf{x}_{i}+(1-\lambda)\mathbf{x}_{j},\qquad\mathbf{\tilde{y}}=\lambda\mathbf{y}_{i}+(1-\lambda)\mathbf{y}_{j}$ (14)
-where $(\mathbf{x}_{i},\mathbf{y}_{i})$ and $(\mathbf{x}_{j},\mathbf{y}_{j})$ are two labeled images from the dataset, and $\lambda\in[0,1]$.
-As a variant, manifold mixup [127] is operated on the hidden layers of the neural network.
-Most mixup methods combine two different images, while AugMix [128] applies three operations to one image, and combines the three augmented images to generate a new sample.
-In some approaches, the mixing operation varies in strength across regions.
-For example, SmoothMix [129] applies a soft-edged mask when mixing.
-Most mixup methods create new samples with soft labels, as in Eq.14, while there are some exceptions.
-Sample Paring [130], for example, uses the label of one of the original images, and that of the other image is left unused.
-ReMix [131] addresses the class imbalance problem by assigning the minority class label to augmented data.
-Neural Blending
-AutoMix [132] integrates a Mix Block in the training process.
-It builds a bridge between the selection of mixup policy and the optimization of the model.
-In AutoMix, the augmentation module is optimized along with the classification feature map, and the mixup functions and mixing ratio will updated according to the classification loss.
-Another image blending technique is proposed as Smart Augmentation [133].
-At the input end of Smart Augmentation, two random samples are fed into the network through two channels, and a CNN sub-network yields a one-channel output with the same shape as the input, which means one new image is generated based on the two input images.
-The augmented image is then used for downstream tasks, and the augmentation sub-network adjusts to the task performance accordingly.
-V-A2 Text Representation Mixture
+V-A1 Image Mixup
+Mixup [125] mixes two random images on the input layer by calculating their interpolation pixel-wise.
+Pixels from two original images each contribute a portion to the new image: $\mathbf{\tilde{x}}=\lambda\mathbf{x}_{i}+(1-\lambda)\mathbf{x}_{j},\mathbf{\tilde{y}}=\lambda\mathbf{y}_{i}+(1-\lambda)\mathbf{y}_{j}$, where $(\mathbf{x}_{i},\mathbf{y}_{i})$ and $(\mathbf{x}_{j},\mathbf{y}_{j})$ are two labeled images from the dataset, and $\lambda\in[0,1]$.
+Manifold mixup [126] is operated on the hidden layers of the neural network.
+AugMix [127] applies three operations to one image and combines the three augmented images to generate a new sample.
+SmoothMix [128] applies a soft-edged mask to mix different regions in varied strength.
+Most mixup methods create new samples with soft labels, while sample pairing [129] uses the label of one of the original images, and that of the other image is left unused.
+Similarly, ReMix [130] addresses the class imbalance problem by assigning the minority class label to augmented data.
+V-A2 Text Mixup
 Unlike image data, whose coloring information is expressed by numerical values and can be interpolated at the input layer, words are difficult to interpolate directly, so text interpolation is usually done at the embedding layer.
-Text Mixup
 The adaptation of Mixup in sentence classification is studied in [131], which proposes wordMixup and senMixup.
-WordMixup interpolates a pair of word embeddings and then feeds the interpolation to the sentence encoder.
-SenMixup, on the other hand, interpolates a pair of sentence embeddings.
-The embedding vector produced by WordMixup or SenMixup is then passed to the softmax layer to produce the target distribution of the classification task.
-The use of mixup in text data augmentation has also been extensively studied in many research, such as nonlinear Mixup [135] and DoubleMix [136].
-GPT3Mix [137] makes use of the latest large language model to mix sentences.
-By inserting two source sentences in a prompt template that describes the instruction to the model, we can get a mixed sentence of high quality with a soft label.
-V-A3 Node Mixture
-Feature or label information from different nodes can be mixed by Graph Propagation, which is often used for tasks such as node classification.
-Graph Propagation
+Its wordMixup interpolates a pair of word embeddings and then feeds the interpolation to the sentence encoder.
+On the other hand, senMixup interpolates a pair of sentence embeddings.
+The use of mixup in text data augmentation has also been studied in other research, such as nonlinear Mixup [132] and DoubleMix [133].
+V-A3 Graph Propagation
 Propagation refers to spreading information from one node to other nodes along the graph.
@@
 MV-GCN [134] first creates three different views from the original graph, propagates the features in each, and then combines the three representations.
-Label propagation algorithm (LPA) [139] propagates labels from labeled nodes to all unlabeled nodes according to their proximity through areas with dense unlabeled nodes.
-AutoGRL [140] applies LPA in their semi-supervised framework to label more nodes utilizing a few ground truth nodes.
+The label propagation algorithm (LPA) [135] propagates labels from labeled nodes to all unlabeled nodes according to their proximity through areas with dense unlabeled nodes.
+AutoGRL [136] applies LPA in its framework to label more nodes utilizing a few ground truth nodes.
 [137] performs two types of label propagation, correcting the base predictions and smoothing the final prediction, respectively.
-Though not in a typical graph-based scenario, [142] constructs a nearest neighbor graph on the classifier’s manifolds.
-The affinity matrix, analogous to the adjacency matrix in this case, is constructed based on the $k$ nearest neighbors of each node’s embeddings, and this matrix decides how labels propagate.
-V-A4 Cell Value Mixture
+V-A4 Tabular Mixup
 Due to its well-aligned structure, tabular data is very suitable for interpolation.
-Tabular Mixup
-SMOTE [37] is among the first works to study data interpolation, even before the proposal of Mixup [126].
+SMOTE [37] is among the first works to study data interpolation.
 It addresses the problem of class imbalance by oversampling from the minority class.
-The samples are selected within the $k$ nearest neighbors in the feature space and new synthetic data are produced by a random interpolation operation.
-Since tabular features are often nominal, two variants, namely SMOTE-NC and SMOTE-N, are also proposed along with the initial version of SMOTE [37].
+The samples are selected from the $k$ nearest neighbors in the feature space, and new synthetic data are produced by a random interpolation operation.
 As argued in [138], SMOTE is subject to unrealistic samples when interpolating in a sparse feature space.
@@
 They interpolate data in the input layer and the hidden layer, respectively.
-V-A5 Sequence Value Mixing
-Methods that combine information from different sequences are referred to as sequence averaging or sequence mixing.
+V-A5 Sequence Mixing
 DTW Barycenter Averaging (DBA) [143] revolves around the averaging problem for Dynamic Time Warping.
@@
 It first uses a sequence autoencoder to learn a feature space from unlabeled data.
-The data is then encoded and augmented with additive noises, interpolation, and extrapolation.
+The data is then encoded and augmented with additive noise, interpolation, and extrapolation.
 Finally, the data in feature spaces is decoded and classified.
@@
 V-B Structure-based Combination
-V-B1 Image Structural Combination
+V-B1 Image patching
 Image patching puts together small fragments of many images or puts a part of an image over a second image.
-Image Patching
-RICAP [152] randomly crops four images, and patches them to construct a new image.
-CutMix [146] makes use of two images.
-It is developed based on Cutout [9], but the removed region is replaced by a patch from another image.
-Attentive CutMix [153] further enhances the CutMix method by adding a feature extraction module that selects an important or representative region, and pastes this attentive patch onto the other image.
-CowMask [154] introduces patching with irregular shape rather than rectangular.
-V-B2 Text Structural Combination
+RICAP [147] randomly crops four images and patches them to construct a new image.
+CutMix [141] is developed based on Cutout [10], but the removed region is replaced by a patch from another image.
+Attentive CutMix [148] further enhances the CutMix method by adding a feature extraction module that selects an important or representative region and pastes this attentive patch onto the other image.
+CowMask [149] introduces patching with an irregular shape rather than a rectangular one.
+V-B2 Text Fragment Merging
 Text fragment merging combines text data by merging text fragments from one sentence into another.
-Text Fragment Merging
-GECA [155] aims to remove a fragment from a sentence and populate it with another fragment that occurs in some common environment.
+GECA [150] removes a fragment from a sentence and populates it with another fragment that occurs in some common environment.
 Subtree Swapping [151], Substructure Substitution [152], and TreeMix [153] use tree structures to analyze the part of speech or role in the sentences and thus make reasonable swaps.
-SSMix [159] replaces a low saliency segment in a sentence with a segment with high saliency from another sentence and mixes sentence labels according to the ratio of their span length in the combined sentence.
+SSMix [154] replaces a low-saliency segment in a sentence with a segment with high saliency from another sentence and mixes sentence labels according to the ratio of their span length in the combined sentence.
 V-B3 Table Concatenation
@@
 If these restrictions are relaxed, models can learn across different tables, be more transferable, and better predict unseen tables.
-Structure Relaxation
 TransTab [155] suggests that multiple tables could share partially overlapped columns in the real world, but traditional methods often fail to fully utilize the data due to the removal of non-overlapping columns and mismatched samples when cleaning the data.
@@
 V-C1 Image Value-structure Mixture
+CNN models merge images by fusing pixels through channels and across the region within the reception fields.
 Diffusion models take textual prompts as inputs and can perform a flexible and smart mixture of existing images.
+Neural Blending
+AutoMix [156] integrates a mix block in the training process.
+It builds a bridge between the selection of the mixup policy and the optimization of the model.
+Another image blending technique is proposed as Smart Augmentation [157].
+At its input end, two random samples are fed into the network through two channels, and a CNN sub-network yields a one-channel output with the same shape as the input, which means one new image is generated based on the two input images.
+The augmented image is then used for downstream tasks, and the augmentation sub-network adjusts to the task performance accordingly.
 Prompt-based Mixing
@@
 Diff-II [159] calculates the inversion noise of two images, conducts random circle interpolation, and then denoises with different prompts.
-V-C2 Graph Value-structure Mixutre
-Graph Mixup refers to mixing graph nodes or graph instances. This often involves both node attribute values and the topology structures.
-Graph Mixup
-Due to the irregular connectivity and various topological structures of graph data, it is hard to directly combine nodes or graphs.
-For node-level mixup, one solution to this problem is temporarily ignoring the edges and mixing node features in latent space.
-GraphMix [163] interpolates the hidden states and labels of input nodes.
-A topology-involved node mixup method is proposed in [164].
-It randomly pairs nodes and includes the nodes’ local topology representations when interpolating them.
-S-Mixup (Structural Mixup) [165] considers neighboring edges when pairing nodes but works in the input space and discusses how to connect the node with the original graph with edges.
-It matches nodes based on GNN classification confidence and connects with high-gradient edges to the mixed node.
+V-C2 Text Value-structure Mixture by Generative Mixing
+Generative mixing leverages the text understanding and generation capabilities of LLMs to combine information from multiple sentences.
+GPT3Mix [160] makes use of the latest large language model to mix sentences.
+By inserting two source sentences in a prompt template that describes the instruction to the model, we can get a mixed sentence of high quality with a soft label.
+PromptMix [161] also prompts the LLM to combine information from two sentences and then uses another LLM to relabel the generated augmentation sentence.
+Combining information from two text pieces is a subtle problem, and the specific behaviors of the LLMs highly depend on the prompts, which include the instruction, task description, and few-shot examples.
+So, the users have to try hard to design an ideal prompt so that the model works well.
+V-C3 Graph Value-structure Mixture by Graph Mixup
+Graph mixup refers to mixing graph nodes or graph instances.
+This often involves both node attribute values and the topology structures.
+For node-level mixup, a topology-involved method [162], randomly pairs nodes and includes the nodes’ local topology representations when interpolating them.
+Structural Mixup [163] considers neighboring edges when pairing nodes and discusses how to connect the node with the original graph with edges.
 For graph-level mixup, ifMixup [164] adds dummy nodes so that two graphs have the same number of nodes.
-These dummy nodes have all-zero feature vectors and are not connected to any other nodes, or in other words, the weights of edges on dummy nodes are zeros.
-S-Mixup (Graph Mixup with Soft Alignments) [167] also aligns graphs first before mixing them.
-It obtains an assignment matrix by a soft assignment and uses the assignment matrix to transform a graph so that every node corresponds with a node in the other graph.
-Compared with the above approaches, G-Mixup [168] proposes a more sophisticated way to interpolate graph data.
-It first uses graphs within the same class to estimate a graphon, i.e.
-graph generator:
-$\vspace{-2mm}\mathbf{\mathcal{G}}\rightarrow W_{\mathbf{\mathcal{G}}},\mathbf{\mathcal{H}}\rightarrow W_{\mathbf{\mathcal{H}}}$ (15)
-The graphon $W:[0,1]^{2}\rightarrow[0,1]$ is a matrix, where $W(i,j)$ describes the probability of an edge between nodes $i$ and $j$, and graphs of the same class are considered to be generated from the same graphon.
-Graphons of two classes of graphs are then interpolated to produce a mixed graphon:
-$\vspace{-2mm}W_{\mathbf{\mathcal{I}}}=\lambda W_{\mathbf{\mathcal{G}}}+(1-\lambda)W_{\mathbf{\mathcal{H}}}$ (16)
-By sampling from the graphon mixup $W_{\mathbf{\mathcal{I}}}$, we acquire a series of graphs with the same soft labels:
-$\vspace{-2mm}\mathbf{y}_{\mathbf{\mathcal{I}}}=\lambda\mathbf{y}_{\mathbf{\mathcal{G}}}+(1-\lambda)\mathbf{y}_{\mathbf{\mathcal{H}}}$ (17)
-During graphon estimation, G-Mixup aligns the original node features and obtains the graphon node features average pooling, which are used as node features of the generated graphs.
+Graph Mixup with Soft Alignments [165] obtains an assignment matrix by a soft assignment and uses that matrix to transform a graph so that every node corresponds with a node in the other graph.
+G-Mixup [166] first uses graphs within the same class to estimate a graphon, i.e., graph generator: $\mathbf{\mathcal{G}}\rightarrow W_{\mathbf{\mathcal{G}}},\mathbf{\mathcal{H}}\rightarrow W_{\mathbf{\mathcal{H}}}$.
+The graphon describes the probability of an edge between nodes $i$ and $j$.
+Graphons of two classes of graphs are then interpolated to produce a mixed graphon: $W_{\mathbf{\mathcal{I}}}=\lambda W_{\mathbf{\mathcal{G}}}+(1-\lambda)W_{\mathbf{\mathcal{H}}}$.
+By sampling from the graphon mixup $W_{\mathbf{\mathcal{I}}}$, we acquire graphs with the same soft labels: $\mathbf{y}_{\mathbf{\mathcal{I}}}=\lambda\mathbf{y}_{\mathbf{\mathcal{G}}}+(1-\lambda)\mathbf{y}_{\mathbf{\mathcal{H}}}$.
 DAGAD [167] mixes graph representations by performing random permutations, demonstrating that data augmentation is helpful in scenarios such as graph anomaly detection.
@@
 A straightforward way to produce augmented data from the entire dataset is to learn the data distribution of existing datasets with neural networks and generate new samples that fall in that distribution.
-Some typical solutions are Autoencoder (AE), Generative Adversarial Networks (GAN), Large Language Models (LLMs), and Generative Diffusion Models.
+Some typical solutions are autoencoders (AE), generative adversarial networks (GAN), large language models (LLMs), and generative diffusion models.
 VI-A1 Autoencoder
-An Autoencoder (AE) learns from the dataset and extracts the essential structure, which can be used to generate a more robust dataset.
+An autoencoder learns from the dataset and extracts the essential structure, which can be used to generate a more robust dataset.
 It is widely used for tabular data augmentation.
@@
 The samples are iteratively selected by active learning, labeled by an oracle, reconstructed by the VAE, and discriminated and classified.
+The use of AEs for feature space augmentation on image, text, and time-series data is discussed in [145], [16], and [171].
+However, few follow-ups use such methods, potentially due to the inferior capacity compared with more advanced generation models.
+For instance, VAEs fail on complex datasets primarily because the limited approximation capacity of the encoder results in a poor, often overly simplistic, posterior distribution [172].
+When applied to graph data, VAE models also require a massive graph-matching process [173].
 VI-A2 Generative Adversarial Networks
-The rise of Generative Adversarial Networks (GANs) [173] makes it possible for humans to generate new images on a large scale using computers.
-This nature also leads to a new path of data augmentation: creating new data samples instead of modifying or combining old ones.
-It can produce considerably more augmented data and is especially useful for oversampling and alleviating the problem of class imbalance.
-For image data, GANs have now been widely used as a data augmentation technique [12], especially in the medical field [174, 175], where data is to some extent scarce.
-Apart from simply enlarging the dataset, GAN can be used in other ways such as removing background noise and cleaning the dataset [176, 177, 178].
+GANs enable generating new images on a large scale, leading to a new path of data augmentation.
+For image data, GANs have now been a common data augmentation technique, especially in data-scarce scenarios like the medical field [174, 175].
+Apart from simply enlarging the dataset, GAN can be used in other ways, such as removing background noise and cleaning the dataset [176, 177, 178].
 It can also generate different views or poses of the object based on the original image [179], and thus add more variation to the dataset.
+Despite the huge success of GAN’s application for image data augmentation, a benchmark on biomedical image augmentation indicates that GAN-based methods do not generally outperform conventional data augmentation, very likely due to their limited generalizability [180].
 Tabular data augmentation is another case where GANs play an important role.
@@
 It models functional dependencies with autoencoders and includes a term to characterize the set of functional dependency constraints in the generator loss function.
-Arguing that GANs are difficult to train on small datasets, ConvGeN [184] designed a new mechanism that requires the GAN’s generator and discriminator to cooperate, that the generator generates batches of a convex combination of minority neighborhood samples and the discriminator classifies these augmented minority samples against majority batches.
+Apart from application to general tabular data, [185] extends the use of GAN to transcriptomics data, where each row of the table represents a sample, each column represents a gene, and the table’s values record the genes’ expression levels.
 GANs are also widely used to generate virtual time-series data.
@@
 WaveGAN is used on the waveform, while SpecGAN is used on the spectrogram.
-SpecGAN demonstrates how GAN is applied to generate time-series data on the frequency domain.
+SpecGAN demonstrates how GAN is applied to generate time-series data in the frequency domain.
+Although it is possible to generate graph data with GANs, [189] points out that GANs on graphs easily fall into mode collapse, where the generators tend to output non-unique samples, limiting GANs’ application on graph datasets.
 VI-A3 Large Language Models
-Recent advances in large language models (LLM) and generative pre-training (GPT) [188] models have demonstrated unprecedented power in text understanding, classification, and generation tasks.
-Apart from supporting intelligent chat-bots that help people in their daily lives, the capacity of these state-of-the-art language generation models establishes a new paradigm of text data augmentation.
-LAMBADA [189] is among the first attempts to use a state-of-the-art language generator to synthesize new text data for data augmentation, and its procedure is as follows.
-LAMBADA first trains a baseline classifier from a small dataset, and then it fine-tunes the GPT-2 [190] model.
-A set of labels is thereafter fed to the fine-tuned language model, which generates a large dataset of labeled sentences.
-These sentences are evaluated by the classifier trained prior according to its confidence score, and top-ranked sentences are retained to compose the synthesized dataset.
-Similarly, [191] integrates GPT-2 in its text data augmentation process and improves the performance of active learning on tiny datasets.
-It uses the language model to generate a tree of tokens as candidates and construct a sentence from them.
-The set of new sentences is repeatedly refined by active learning, and the outcome is an augmented text dataset.
-Some other research also applies GPT-2 to effectively augment low-resource text datasets, such as [192] and [193].
-Despite their success, both of them point out the need for more sophisticated methods to generate text data.
-[192] for example, argues that human-in-the-loop for relabeling might be rewarding.
-[193] adopts human-in-the-loop for seed selection and further discusses the potential benefits of injecting expert knowledge for specialized domains.
-Notably, Data Boost [194] proposes a reinforcement learning framework to guide the process of text data augmentation with GPT-2 and succeeds in generating data of high quality and high similarity to the original data.
-The rise of GPT-3 [195] and later LLMs is now providing more possibilities for text generative augmentation. Their use and improvements are studied in the latest research such as [196] and [197].
+Recent advances in large language models have established a new paradigm of data augmentation, especially benefiting the NLP field.
+LAMBADA [190] is among the first attempts to use a state-of-the-art language generator to synthesize new text data for data augmentation.
+It generates a large dataset using a set of labels.
+Similarly, [191] uses the language model to generate a tree of tokens as candidates and constructs a sentence from them, and repeatedly refines the new sentences to build the augmented dataset.
+Data Boost [192] proposes a reinforcement learning framework to guide the generation process of GPT-2 to generate high-quality and high-similarity data.
+The rise of GPT-3 and later LLMs is now providing more possibilities for text generative augmentation, such as [193] and [194].
+The development of multi-modal LLMs also facilitates image data augmentation.
+DIAGen [195] uses the GPT-4 model to provide image descriptions with additional meaningful context and uses the enriched textual descriptions to generate diverse image datasets.
+[196] designs a variety of textual prompts and utilizes the DALL-E model to produce an exclusively LLM-generated agricultural dataset, avoiding traditional fieldwork with imaging sensors.
+LLMs can also be applied to tabular data augmentation by treating the table as text content.
+CLLM [197] uses the background description, in-context examples, and instruction guidelines to construct prompts and use the LLM to generate a large synthetic dataset, followed by a curator that evaluates and selects samples by their learning dynamics.
+Pred-LLM [198] first converts the table rows to text descriptions to fine-tune an LLM.
+Then it masks some features of sampled rows and uses the LLM to recover these rows.
+Finally, it uses the LLM to label the generated data.
+Time-LLM [199] demonstrates the predictive ability of Transformer-based LLMs for time-series data, suggesting another possibility of LLMs as a data augmentation technique.
+However, beyond simply using LLMs to predict and expand sequences to future timestamps, the use of LLMs for general time-series data augmentation is still underexplored.
+Some exceptions are in the audio, voice, and speech fields, where the language information can be utilized by LLMs, enabling more possible manipulations, such as ArzEn-LLM [200], which employs the code-switched augmentations for multilingual speech recognition performance.
 VI-A4 Generative Diffusion Models
-Despite that AEs and GANs can all be used to generate graph data, they face many problems in practice.
-VAE models require things like a massive graph-matching process, and GANs on graphs easily fall into mode collapse [198].
-So, graph data augmentation resorts to another recently proposed technique, namely, Generative Diffusion Models.
-The denoising diffusion model [199] is a generative paradigm that has become popular in recent years.
-Though not yet widely applied as an approach to graph data augmentation, the Generative Diffusion Model boasts enormous potential in graph generation and can possibly be adopted as a promising data augmentation technique.
-To implement diffusion models on graph data, some properties of the graph data must be properly treated.
+Diffusion models have recently emerged as a new solution to generative AI.
+Generating realistic and diverse samples by learning the underlying data distributions, they also show promising capabilities as a tool for data augmentation.
+The diffusion model is widely applied to image data.
+GIF [201] leverages Stable Diffusion to optimize the latent features of the seed data in the semantically meaningful space of the prior model and create informative and realistic new data.
+SIGD [202] employs diffusion models to generate augmented images with good image diversity.
+It takes image labels and captions as guidance to maintain semantic consistency.
+[203] prompts the diffusion models to “generate a clean background” to avoid additional objects in the background and designs a background mask erosion mechanism to avoid extending objects to the background.
+For text data, DiffusionCLS [204] leverages a diffusion LM to capture in-domain knowledge and generate pseudo samples by reconstructing from the corrupted dataset to balance between consistency and diversity.
+DiffLM [205] leverages diffusion models to reserve more information of the dataset’s original distribution and format structure in the learned latent distribution and decouples the learning of target distribution knowledge from the LLM’s generative objectives via a plug-and-play latent feature injection module.
+The diffusion model also fills the gap of tabular data augmentation.
+[206] leverages diffusion models to first learn an unconditional generative model and introduces lightweight controllers to guide the unconditional generative model in generating synthetic data that satisfies different conditions.
+TabDiff [207] develops a joint continuous-time diffusion process for numerical and categorical data and proposes feature-wise learnable diffusion processes to counter the high disparity of different feature distributions.
+Forest-Diffusion [208] introduces a novel approach for generating and imputing mixed-type tabular data utilizing score-based diffusion and conditional flow matching.
+Compared with other modalities, using the diffusion model for graph data and time-series data augmentation is relatively underexplored.
+To implement diffusion models on graphs, some properties of graphs must be properly treated.
 For example, diffusion models often rely on a continuous Gaussian noise process, but the graph structures are discrete [209].
-DiGress [201] designs a discrete denoising diffusion model for graph generation and achieves state-of-the-art performance on molecular and non-molecular datasets.
-The use of Generative Diffusion Models for data augmentation especially benefits studies on molecules, proteins, and materials, whose data are often in the form of graph instances.
+In response, DiGress [210] designs a discrete denoising diffusion model for graph generation and achieves state-of-the-art performance on molecular and non-molecular datasets.
+Future advancement of generative diffusion models for graph data augmentation will especially benefit studies on molecules, proteins, and materials, whose data are often in the form of graph instances.
+For tabular data, [211] proposes ASE-DDPM for unbalanced time series classification.
+[212] employs the diffusion model for time-series data but only uses it to generate the first moment data.
+Despite these attempts, literature focusing on the design and application of diffusion-based time-series data augmentation is still scarce.
+Given the remarkable capability of diffusion models and their success in other data modalities, it is promising to further delve into this field.
 VI-B Exogenous Generation
@@
 Table Retrieval
-Tabular data populates the web and by retrieving consistent and related tables, it is possible to generate an augmented table that incorporates extended information.
+Tabular data populates the web, and by retrieving consistent and related tables, it is possible to generate an augmented table that incorporates extended information.
 A key consideration of table retrieval augmentation is to design appropriate ways to search for ideal tables.
@@
 Here we introduce three examples of how humans utilize expert knowledge for image, tabular, and time-series data augmentation
-Computer Graphic Modeling is to leverage expert knowledge on how objects in the existing dataset look like and the power of computer software to generate new image samples [207, 208, 209].
+Computer Graphic Modeling is to leverage expert knowledge on how objects in the existing dataset look and the power of computer software to generate new image samples [218, 219].
 These datasets have several advantages over ordinary image datasets.
@@
 In other words, it learns a graph from the features in the table and obtains augmented data from that graph.
-PET [211] depicts multiple tabular data instances as a hyper-graph and uses the labels to help construct hyper-edges.
+PET [221] depicts multiple tabular data instances as a hypergraph and uses the labels to help construct hyperedges.
 These methods are based on external assumptions and introduce extra structural information beyond the tabular dataset.
@@
 VII Discussions
-Up to now this survey focuses on wherefrom the information in the augmented samples is derived. And there are many other perspectives to discuss data augmentation and understand how it works.
+Up to now, this survey focuses on wherefrom the information in the augmented samples is derived. And there are many other perspectives to discuss data augmentation and understand how it works.
 VII-A Augmentation Evaluation Metrics
@@
 So, there is a need for a set of evaluation metrics that can assess specific considerations of data augmentation.
-Two of the most critical metrics are affinity and diversity [161, 162, 214, 215, 216, 217, 125].
+Two of the most critical metrics are affinity and diversity [224, 225, 226].
 Affinity is also referred to as consistency, correctness, faithfulness, quality, similarity, or validity.
 It reflects if the augmented sample is semantically consistent with the original sample and preserves the original label.
-Model-centric evaluation [215] defines affinity as the ratio of model $m$’s performance on augmented validation dataset $D^{\prime}_{val}$ and that on the original validation dataset $D_{val}$, given by
-[…]
-where $\text{ACC}(m,D)$ stands for model $m$’s performance on evaluation dataset $D$.
+Model-centric evaluation [225] defines affinity as the ratio of model $m$’s performance on the augmented validation dataset $D^{\prime}_{val}$ and that on the original validation dataset $D_{val}$, given by $\mathit{affinity}=\text{ACC}(m,D^{\prime}_{val})/\text{ACC}(m,D_{val})$, where $\text{ACC}(m,D)$ stands for model $m$’s performance on the evaluation dataset $D$.
 Data-centric evaluation, on the other hand, defines affinity as how much the augmented dataset deviates from the original dataset in the feature space.
-For example, [214] proposes Fréchet BERT Distance (FBD) to evaluate the deviation of two textual datasets, defined as:
-[…]
-where $m$ is the mean vector and $C$ is the covariance matrix.
-Diversity reflects the breadth of patterns in the dataset and exposing the model to a diverse training dataset can improve its robustness.
-[218] considers a dataset as an ellipsoid and defines its diversity as the isocontour radius of the ellipsoid:
-[…]
-where $H$ is the dimension of the feature vectors and $\sigma_{i}$ is the standard deviation along the $i$-th feature. [219] proposes to evaluate dataset diversity by calculating the dispersion:
-[…]
-where $d(x_{i},x_{j})=1-\frac{f(x_{i}){\cdot}f(x_{j})}{\|f(x_{i})\|\|f(x_{j})\|}$.
-Some empirical studies including [219], [217], and [215] investigate how affinity and/or diversity influence the effectiveness of data augmentation, and their real-world experimental results demonstrate the model performance takes the most advantage of data augmentation that achieves both high affinity and diversity.
+For example, [224] proposes Fréchet BERT Distance (FBD) to evaluate the deviation of two textual datasets, defined as: $\mathit{FBD}=\sqrt{\|m_{1}-m_{2}\|^{2}_{2}+Tr(C_{1}+C_{2}-2(C_{1}C_{2})^{1/2}}$, where $m$ is the mean vector and $C$ is the covariance matrix.
+Diversity reflects the breadth of patterns in the dataset, and exposing the model to a diverse training dataset can improve its robustness.
+[227] considers a dataset as an ellipsoid and defines its diversity as the isocontour radius of the ellipsoid: $\mathit{diversity}=(r_{1}{\cdot}r_{2}{\cdot}...{\cdot}r_{H})^{\frac{1}{H}}=\sqrt[H]{\prod^{H}_{i}\sigma_{i}}$, where $H$ is the dimension of the feature vectors and $\sigma_{i}$ is the standard deviation along the $i$-th feature.
+[228] proposes to evaluate dataset diversity by calculating the dispersion: $\mathit{disersion}=\sum_{x_{i},x_{j}}d(x_{i},x_{j})$, where $d(x_{i},x_{j})=1-\frac{f(x_{i}){\cdot}f(x_{j})}{\|f(x_{i})\|\|f(x_{j})\|}$.
+Some empirical studies investigate how affinity and diversity influence the effectiveness of data augmentation, and their real-world experimental results demonstrate that model performance takes the most advantage of data augmentation that achieves both high affinity and diversity [225, 226, 228].
 VII-B Selecting the Target
 VII-B1 On Selecting Pixels
-For Pixel Erasing, selecting target pixels is a necessity because it is impossible to mask out all pixels in the image.
-If erasing is performed by randomly selecting an area in the image, it is called Image-aware Random Erasing (IRE) [10].
-In some tasks, the location of the target in the image is known, so, random erasing can be restricted in the object’s bounding box, known as Object-aware random erasing (ORE) [10].
-Further argued in [220], the erased area within the bounding box should be limited to a manageable size and a part of the less informative background should also be erased.
+For pixel erasing, selecting target pixels is a necessity before erasing them.
+Image-aware Random Erasing (IRE) [11] randomly selects an area in the image.
+Object-aware random erasing (ORE) [11] restricts random erasing to the object’s bounding box if its location is known.
 [229] proposed Region-aware Random Erasing (RRE), which introduces ORE with upper bounds for width and height in the bounding boxes, and then randomly erases the background.
-The same consideration occurs when performing Image Cropping and Mixup, and some importance-based strategies are used to preserve important contents in the images.
-For Image Cropping (IC), [74] uses entropy as the importance score to select the most importance regions, and [75] uses eye-tracking technique to locate important content.
-For Image Mixup, Co-mixup [221] and Puzzle mix [222] both use saliency to decide the mixing strength of different regions, potentially leading to better mixing performances.
-A summary of how these methods select target pixels are presented in Table III.
-TABLE III: Region selection strategy for Pixel Erasing, Image Cropping and Mixing methods.
-Name Ref Augmentation Strategy
+The same consideration also applies to image cropping (IC) and mixup, and some importance-based strategies are used to preserve important contents in the images.
+For IC, [73] uses entropy as the importance score to select the most important regions, and [74] uses eye-tracking technique to locate important content.
+For mixup, Co-mixup [230] and Puzzle Mix [231] both use saliency to decide the mixing strength of different regions.
+A summary of how these methods select target pixels is presented in Table IV.
+TABLE IV: Region selection strategy for pixel erasing, image cropping, and mixing methods.
 Random Location-based Importance-based
-Cutout [9]
-Image-aware RE [10]
-Object-aware RE [10]
-Region-aware RE [220]
-Hide-and-seek [38]
-Grid-Mask [39]
-Entropy-bases IC [74]
-Gaze-based IC [75]
-Co-mixup [221]
-Puzzle mix [222]
+Cutout [10], Image-aware RE [11], Region-aware RE [229], Hide-and-seek [38], Grid-Mask [39] Object-aware RE [11], Region-aware RE [229] Entropy-bases IC [73], Gaze-based IC  [74], Co-mixup [230], Puzzle mix [231]
 VII-B2 On Selecting Words
-Most value-based text transformation methods randomly select the word to be replaced, added, and deleted, but some methods considers the role a word plays in the sentence when deciding whether to impose perturbation.
+Most value-based text transformation methods randomly select the word to be replaced, added, and deleted, but some methods consider the role a word plays in the sentence when deciding whether to impose perturbation.
 For example, TEXTFOOLER [232] ranks words by their importance and replaces the important words with their synonyms.
-Syntax-aware data augmentation [48] argues that most previous text data augmentation approaches are only computation-motivated and cannot assure the good quality of the generated text.
-Instead of simply modifying words with the same probability, Syntax-aware data augmentation calculates the probability of being modified for each word based on its syntactic role in the sentence.
-Specifically, it utilizes a dependency tree to calculate the importance of each word, i.e.
-the depth in the tree.
-A word far from the root is considered less important syntactically and is more likely to be removed (by means of dropout and blanking) or replaced during the augmentation process.
-Instead of using a syntax tree to assess word importance, Selective Text Augmentation (STA) [224] focuses more on the statistical and semantic features of words to describe the role of words.
-It proposes two perspectives to view words in text classification tasks, namely, statistical correlation, which means how frequently a word co-occurs with a certain category, and semantic similarity, which measures the semantic overlap of the given words and the category label.
-Using the metrics as axes, STA divides the roles of words into four quadrants: gold, venture, bonus, and trivial.
-Words with different roles are treated differently when imposing augmentation operations on the sentence.
-For example, gold words will be kept in selective deletion and synonyms of gold words will be added to the sentences in selective insertion.
-A summary of how these methods select target words is presented in Table IV.
-TABLE IV: Word selection strategy of text data selection methods.
-Name Ref Augmentation Strategy
+Syntax-aware data augmentation [48] improves upon older methods by modifying words based on their grammatical role in a sentence.
+It uses a sentence’s structure (a dependency tree) to determine a word’s importance; words less central to the sentence’s core meaning are more likely to be changed or removed.
+Selective Text Augmentation (STA) [233] ignores grammatical structure and instead classifies words based on their statistical and semantic relationship to a category.
+It then groups words into four roles (like ”gold” for important words).
+During augmentation, it treats words differently based on their role; for example, it will keep ”gold” words but add synonyms for them.
+A summary of how these methods select target words is presented in Table V.
+TABLE V: Word selection strategy of text data selection methods.
 Random Role-based Learned
-EDA [11]
-- [43]
-TEXTFOOLER [223]
-LetterTran [225]
-AEDA [46]
-INSNET [226]
-LeCA [227]
-GBS [228]
-- [47]
-STA [224]
-Subtree Swapping [156]
-Substructure Substitution [157]
-TreeMix [158]
-SSMix [159]
+EDA [12], [42], AEDA [46], [47] TEXTFOOLER [232], LeCA [234], GBS [235], STA [233], Substructure Substitution [152], Subtree Swapping [151], TreeMix [153] LetterTran [236], INSNET [237], SSMix [154]
 VII-B3 On Selecting Nodes and Edges
@@
 Node centrality acts as an important factor that is both meaningful and convenient, supporting selection strategies of many graph data augmentation methods.
-GCA [229] uses node centrality to calculate the masking vectors and rewiring probabilities and thus decide which node to mask and which edge to rewire.
-GrpahCrop [93] applies a node-centric strategy to preserve connectivity in subgraphs.
+GCA [238] uses node centrality to calculate the masking vectors and rewiring probabilities and thus decides which node to mask and which edge to rewire.
+GraphCrop [92] applies a node-centric strategy to preserve connectivity in subgraphs.
 SUBG-CON [93] uses the Personalized PageRank centrality to select the important nodes for subgraphing.
@@
 Similarly, RobustECD [239] and M-Evolve [240] also consider node similarity for edge rewiring.
-TABLE V: Node and edge selection strategy for grpah data augmentation methods.
-Name Ref Augmentation Strategy
+TABLE VI: Node and edge selection strategy for grpah data augmentation methods.
 Random Centrality Similarity
-Attribute masking [49]
-GRAND [50]
-BGRL [51]
-GCA [229]
-GrpahCrop [93]
-SUBG-CON [94]
-Smart rewiring [103]
-DHGR [106]
-RobustECD [230]
-M-Evolve [231]
+Attribute masking [49], GRAND [50], BGRL [51] GCA [238], GrpahCrop [92], SUBG-CON [93], Smart rewiring [102] DHGR [105], RobustECD [239], M-Evolve [240]
 VII-C Leveraging Augmented Samples
@@
 A frequent usage of augmented samples is to construct different views for contrastive learning.
-For example, Rock-ViT [232] performs rotation, flipping, blurring, contrast enhancement, irregular noise, and microscope dark-field noise on sandstone microscopic images.
+Rock-ViT [241] performs rotation, flipping, blurring, contrast enhancement, irregular noise, and microscope dark-field noise on sandstone microscopic images.
 It optimizes the vision transformer with both cross-entropy loss and contrastive loss.
@@
 GraphCL [49] uses Node Dropping and Edge Perturbation to build different views, and MVGRL [98] applies graph diffusion.
-Different views of an original graph are used as positive pairs and a contrastive loss function is used to enforce maximizing their consistency.
-VIME [56] expend the usage of data augmentation beyond contrastive learning.
+Different views of an original graph are used as positive pairs, and a contrastive loss function is used to enforce maximizing their consistency.
+VIME [56] extends the usage of data augmentation beyond contrastive learning.
 It incorporates tabular data augmentation in both self-supervised and semi-supervised learning frameworks.
@@
 For semi-supervised learning, it calculates the consistency loss between masked views.
-VII-D Harnessing Latest Techniques
-Early data augmentation methods rely on hand-craft manipulations, which are constrained by effectiveness and efficiency.
+VII-D How to Choose the Appropriate Augmentation Method?
+As summarized in many surveys and benchmarks on data augmentation, although different data augmentation methods claim to be effective and boast certain advantages, their practical performance largely depends on specific task scenarios and even nuanced experimental settings.
+So, it is hard to say which method is the best one.
+Still, we summarize some characteristics, either as advantages or limitations, of each method in Table II.
+Specifically, we highlight the methods that (1) perform sample-wise analysis (e.g., saliency map or syntax parsing) before augmentation, (2) require dedicated computation to generate the augmentation output, and (3) are subject to indistinctive or non-recoverable information loss during the augmentation processes.
+We also address some practical considerations when choosing augmentation methods under different circumstances in detail in this section.
+VII-D1 Tradeoff between the Gain and Cost
+Many advanced data augmentation methods claim better performance compared with simple and random methods.
+However, these advanced methods often involve complicated techniques, such as reinforcement learning and generative models, which require even more time and resources to augment the dataset than training on the datasets.
+To balance the performance and cost, a key issue is what we expect from data augmentation.
+Suppose the current dataset is already very large; perhaps we just want to introduce some corruption to improve model robustness, then cheaper methods like image flipping, synonym replacement, and noise addition would be good choices.
+In case of low-resource scenarios where data is scarce, it might be wise to resort to generative methods for their capabilities to introduce diversity and expand dataset coverage [180, 124].
+VII-D2 Be Careful with the Labels
+A crucial consideration when choosing the data augmentation method is the characteristics of the dataset’s labels.
+The labels in some datasets are subject to change under some operations.
+For example, image rotation is not safe for the MNIST dataset, which contains “6” and “9”.
+Similarly, random token insertion might be label-preserving for the SST-2 sentiment classification dataset but is harmful for the CoLA grammar acceptability dataset.
+Multi-instance augmentation, such as promptMix [161], is a powerful tool to refine the decision boundary between categories for a classification dataset.
+However, it is hard to apply to unlabeled datasets in semi-supervised settings, where single-instance augmentation plays a dominant role.
+VII-D3 Compatibility with the Training Framework
+Multi-instance augmentation is also challenged by some model training frameworks.
+Some models only accept hard-labeled data, while mixup-based methods often produce soft labels when mixing samples.
+If data augmentation is performed as an intermediate step in the model’s framework, both input-space and hidden-space augmentation are available, while if augmentation is used before training the model, it would be hard to apply hidden-space augmentation because the embeddings have not been produced.
+VII-E Transferring between modalities
+We have proposed a cross-modality taxonomy that unifies data augmentation for five common modalities.
+From this cross-modality perspective, we notice that many data augmentation methods designed for different modalities, though dubbed with different names, share the same nature with respect to their operation on samples’ information.
+For example, pixel erasing, token deletion, node attribute masking, table masking, and sequence jittering all apply masks and introduce noise to the samples, and image cropping, sentence cropping, subgraph cropping, table subsetting, and sequence window slicing all break the structure of a sample.
+Conversely, such consistency implies that data augmentation techniques well developed for one modality can be transferred to other under-explored modalities, inspiring novel advancements in this field.
+Policy-based augmentation image combines multiple basic operations for image data [14, 115], and since abundant atomic operations are also available for text and graph data, such as token replacement or edge dropping, it could be a very good idea to design agents that search for optimal text or graph augmentation strategies.
+Image patching and text fragment merging combine pieces from multiple images or sentences together as a new sample [141, 151].
+To transfer this idea to tabular data, we could think about if it is possible to combine half columns from a row with the remaining columns from another row.
+Advanced neural networks have enabled a series of augmentation approaches that can generate high-quality and high-fidelity image and text data [118, 124].
+Though yet underexplored, we could expect more similar advancements for graph, tabular, and time-series data.
+VII-F Harnessing the Latest Techniques
+Early data augmentation methods rely on hand-crafted manipulations, which are constrained by effectiveness and efficiency.
 AutoAugment [14] begins to introduce reinforcement learning to find optimal augmentation policies, enabling automatic search for desired augmentation strategies.
 Generative models, such as LLMs and diffusion models, possess outstanding capacities to capture sample semantics and produce new samples with high flexibility and high quality [243, 159, 122].
-However, data augmentation is sometimes only a preprocessing step in a deep learning framework, and introducing generative models might introduce extra computation overhead.
-How to balance the tradeoff between the reward and cost of using these advanced techniques remains an open but critical problem.
+However, most generative data augmentation methods are constrained to simple deployment in their by-design modalities, for example, modifying images with diffusion models or paraphrasing sentences with LLMs.
+In response, we suggest two concrete directions: (1) Designing informed data augmentation methods that take into account the sample distribution and supervision power of the datasets.
+For example, DoAug [124] bridges the gap between LLM generation and datasets’ sample diversity.
+(2) Embracing the cross-modality nature of many generative models and utilizing the advantage of their representative and generative power in different modalities.
+For example, DALDA [244] uses the GPT-4 model to generate image captions and consequently employs the diffusion model to generate corresponding images.
 VIII Conclusion
@@
 Finally, it discusses several critical issues related to data augmentation that can help us better understand this technique.
+Acknowledgments
+This work was supported by the National Natural Science Foundation of China (Grant Nos.
+62406306 and 92470204) and the National Key Research and Development Program of China Grant (No.
+2024YFF0729201).